### v3

# Character Theory for Semilinear Representations

Abstract.
Let $G$ be a group acting on a field $L$, and suppose that $L/L^{G}$ is a finite extension. We show that the category of semilinear representations of $G$ over $L$ can be described in terms of the category of linear representations of $H$, the kernel of the map $G\rightarrow\operatorname{Aut}(L)$. When $G$ is finite and $L$ has characteristic 0 this provides a character theory for semilinear representations of $G$ over $L$, which recovers ordinary character theory when the action of $G$ on $L$ is trivial.
Key words and phrases:
Semilinear representation, Galois extension, Schur index, character theory
2020 Mathematics Subject Classification:
20C15, 20C99, 11R52, 15A04, 16G99
1. Introduction
Suppose that $L$ is a field, and $G$ is a group which acts on $L$ through a group homomorphism $\sigma\colon G\rightarrow\operatorname{Aut}(L)$, where $\operatorname{Aut}(L)$ is the group of field automorphisms of $L$. Write $K\coloneqq L^{G}$, so we can consider the action as a group homomorphism $\sigma\colon G\rightarrow\operatorname{Aut}(L/K)$.
In this context we can consider the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ of finite-dimensional $L$-vector spaces $V$ equipped with a semilinear action of $G$: a group homomorphism $\rho\colon G\rightarrow\operatorname{Aut}_{K}(V)$ such that each $\rho(g)$ is $\sigma_{g}$-semilinear, meaning that
$\rho(g)(\lambda\cdot v)=\sigma_{g}(\lambda)\cdot\rho(g)(v)\ \mbox{ for all }\ g\in G,\lambda\in L,v\in V.$
In this category the morphisms are those $L$-linear maps which commute with the action of $G$.
Such semilinear representations arise naturally, this category being identified through taking global sections with $\operatorname{\mathbf{Vect}}^{G}(X)$, the category of $G$-equivariant vector bundles on $X\coloneqq\operatorname{Spec}(L)$.
When the action of $G$ on $L$ is trivial, this is simply the category $\operatorname{\mathbf{Rep}}_{L}(G)$ of $L$-linear representations of $G$. When the action of $G$ on $L$ is non-trivial however, $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is not very well understood.
1.1. Matrix Interpretation
One can interpret the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ more concretely in terms of matrices. For any $n\geq 1$, to give a semilinear representation of dimension $n$ over $L$ is the same as giving a family of matrices $(A_{g})_{g\in G}$ in $\operatorname{GL}_{n}(L)$, which satisfy
$A_{g_{1}g_{2}}=A_{g_{1}}\cdot{}^{g_{1}\!}A_{g_{2}}\ \mbox{ for all }\ g_{1},g_{2}\in G,$
where $^{g_{1}\!}A_{g_{2}}$ is the matrix obtained by applying $\sigma_{g_{1}}$ to all entries of $A_{g_{2}}$. A homomorphism from this to another representation given by $(B_{g})_{g\in G}$ in $\operatorname{GL}_{m}(L)$ is a matrix $M\in M_{m\times n}(L)$ such that
(1) $B_{g}\cdot{}^{g}M=M\cdot A_{g}\ \mbox{ for all }\ g\in G.$
In particular, the matrix representations equivalent to $(A_{g})_{g\in G}$ are all given by $(B_{g})_{g\in G}$, where
[…]
for $P\in\operatorname{GL}_{n}(L)$. This point of view naturally identifies the set of isomorphism classes of $n$-dimensional semilinear representations of $G$ over $L$ with the pointed set $H^{1}(G,\operatorname{GL}_{n}(L))$.
Unlike linear representations, one cannot easily use techniques from linear algebra to describe such families of matrices when the action of $G$ on $L$ is non-trivial. For example, even a description of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ when $G$ is finite and cyclic is not apparent, as one can no longer appeal to standard techniques such as diagonalisation.
1.2. The Extension Problem
Let $H\coloneqq\ker(G\rightarrow\operatorname{Aut}(L/K))$, a normal subgroup of $G$. As $H$ acts trivially on $L$, any $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ natural restricts to an object $V|_{H}\in\operatorname{\mathbf{Rep}}_{L}(H)$.
Hand in hand with understanding the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is the extension problem for $\operatorname{\mathbf{Rep}}_{L}(H)$:
  1. (1)
    Which $W\in\operatorname{\mathbf{Rep}}_{L}(H)$ extend to a semilinear representation of $G$ over $L$?
  2. (2)
    How many different extensions are possible?
The category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ blends the representation theory of $G$ with the arithmetic of $L$ and $K$. Therefore these questions often contain arithmetic information about the extension $L/K$.
Example 1.1.
Suppose that $d\in{\mathbb{Q}}^{\times}\setminus({\mathbb{Q}}^{\times})^{2}$, $L/K$ is the extension ${\mathbb{Q}}(\sqrt{d})/{\mathbb{Q}}$, and $G=C_{4}$, acting on ${\mathbb{Q}}(\sqrt{d})$ through the natural surjection $C_{4}\twoheadrightarrow\operatorname{Gal}({\mathbb{Q}}(\sqrt{d})/{\mathbb{Q}})$ with kernel $H=C_{2}$. Write $\overline{\cdot}$ for the non-trivial automorphism of $\operatorname{Gal}({\mathbb{Q}}(\sqrt{d})/{\mathbb{Q}})$.
From the matrix description of semilinear representations, to give a one-dimensional semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$ is the same as giving an element $a\in{\mathbb{Q}}(\sqrt{d})$ with
$a\overline{a}a\overline{a}=1.$
Similarly, if $W$ is the non-trivial one-dimensional representation of $C_{2}$ over ${\mathbb{Q}}(\sqrt{d})$, then to give a semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$ extending $W$ is the same as giving an $a\in{\mathbb{Q}}(\sqrt{d})$ with
$a\overline{a}=-1.$
Writing $a=x+y\sqrt{d}$, for $x,y\in{\mathbb{Q}}$, we see that $W$ extends to a semilinear representation if and only if
$x^{2}-dy^{2}=-1$
for some $x,y\in{\mathbb{Q}}$, i.e. that $(x,y)$ is a rational solution to the negative Pell equation.
For such an $a\in{\mathbb{Q}}(\sqrt{d})$, let $V_{a}$ denote the extension of $W$ to a semilinear representation defined by $a$. From the matrix description of semilinear representations, the semilinear extensions isomorphic to $V_{a}$ are exactly $V_{ba\overline{b}^{-1}}$ for $b\in{\mathbb{Q}}(\sqrt{d})^{\times}$. If we write $b=u+v\sqrt{d}$, then we can consider the isomorphic extensions as being parametrised by $u:v]\in{\mathbb{P}}^{1}({\mathbb{Q}})$, and one can compute that
$ba\overline{b}^{-1}=\frac{(u^{2}+dv^{2})x+(2uv)dy}{u^{2}-dv^{2}}+\sqrt{d}\frac{(u^{2}+dv^{2})y+(2uv)x}{u^{2}-dv^{2}}.$
In particular, the statement that there is at most one extension of $W$ to a semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$ if equivalent to the claim that if $X^{2}-dY^{2}=-1$ has a solution $(x,y)\in{\mathbb{Q}}^{2}$, then the set of all solutions in ${\mathbb{Q}}^{2}$ is parametrised by $u:v]\in{\mathbb{P}}^{1}({\mathbb{Q}})$ via the formula
$\left(\frac{(u^{2}+dv^{2})x+(2uv)dy}{u^{2}-dv^{2}},\frac{(u^{2}+dv^{2})y+(2uv)x}{u^{2}-dv^{2}}\right).$
Repeating this with $W$ replaced by the trivial representation $\mathbf{1}$ of $C_{2}$ over ${\mathbb{Q}}(\sqrt{d})$, one considers instead the (positive) Pell equation
$X^{2}-dY^{2}=1.$
This always has the trivial solution $(1,0)$, which corresponds to the fact that $\mathbf{1}$ always extends to a semilinear representation of $C_{4}$ over ${\mathbb{Q}}(\sqrt{d})$. Similarly, taking $(1,0)$ as an initial solution, this has at most one extension to a semilinear representation if and only if all solutions are given by
$\left(\frac{u^{2}+dv^{2}}{u^{2}-dv^{2}},\frac{2uv}{u^{2}-dv^{2}}\right).$
In particular, taking d=1d=-1, this is the standard parametrisation
$\left(\frac{u^{2}-v^{2}}{u^{2}+v^{2}},\frac{2uv}{u^{2}+v^{2}}\right).$
of the $\mathbb{Q}$}-rational points of $X^{2}+Y^{2}=1$ by $u:v]\in{\mathbb{P}}^{1}({\mathbb{Q}})$.
1.3. Contribution
In this paper we provide what we believe to be the first systematic description of the semilinear representation categories $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ beyond the linear case of $\operatorname{\mathbf{Rep}}_{L}(G)$. We work under the following mild assumption, which we impose from now until the end of the introduction:
  • $L/K$ is a finite extension.
For example, this is automatically satisfied if either $G$ is finite or $L$ is finite over its prime subfield. We also note this further implies that $L/K$ is Galois and $G\rightarrow\operatorname{Gal}(L/K)$ is surjective (Remark 3.5).
1.4. Unique Extension
Our first result concerns the second half of the extension problem, and shows that any $W\in\operatorname{\mathbf{Rep}}_{L}(H)$ has at most one extension to a semilinear representation of $G$ over $L$.
Theorem A.
Suppose that $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Then:
  • $V\cong W$ if and only if $V|_{H}\cong W|_{H}$ in $\operatorname{\mathbf{Rep}}_{L}(H)$,
  • The natural map
    $L\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\rightarrow\operatorname{Hom}_{L[H]}(V|_{H},W|_{H})$
    is an isomorphism.
  • In particular, the natural map
    $L\otimes_{K}\operatorname{End}_{L\rtimes G}(V)\rightarrow\operatorname{End}_{L[H]}(V|_{H})$
    is a $K$-algebra isomorphism.
In particular, this tells us that to understand $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, we are reduced to classifying which $W\in\operatorname{\mathbf{Rep}}_{L}(H)$ extend to a semilinear representation of $G$ over $L$: the first half of the extension problem.
1.5. Group Action
One necessary condition for $W\in\operatorname{\mathbf{Rep}}_{L}(H)$ to extend can be described in terms of a group action of $\operatorname{Gal}(L/K)$ on $\operatorname{\mathbf{Rep}}_{L}(H)/\!\cong$, the set of isomorphism classes of objects of $\operatorname{\mathbf{Rep}}_{L}(H)$.
Definition 1.2.
For $g\in G$ and $W\in\operatorname{\mathbf{Rep}}_{L}(H)$ we set $g*W\in\operatorname{\mathbf{Rep}}_{L}(H)$ to be $W$ as a $K$-vector space, with action of $L$ and $H$ defined by
  • $h*w=(g^{-1}hg)(w)$,
  • $\lambda*w=\sigma_{g^{-1}}(\lambda)w$,
for $\lambda\in L$, $h\in H$ and $w\in W$.
This defines an left action of $G$ on $\operatorname{\mathbf{Rep}}_{L}(H)/\!\cong$ which is trivial on $H$, and therefore induces an action of the quotient $G/H$, and so also $\operatorname{Gal}(L/K)$ using the isomorphism $G/H\xrightarrow{\sim}\operatorname{Gal}(L/K)$.
If $W=V|_{H}$ for some $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, then for any $g\in G$, the action of $g$ on $V$ defines an isomorphism $\rho(g)\colon g*W\xrightarrow{\sim}W$, and so $W$ is fixed by $\operatorname{Gal}(L/K)$. Therefore a necessary condition for $W$ to extend to a semilinear representation of $G$ over $L$ is that $W$ is fixed by $\operatorname{Gal}(L/K)$.
The converse is not true however: returning to Example 1.1, the representation $W$ is always fixed by $\operatorname{Gal}(L/K)$, but when $d$ is square-free the negative Pell equation has a solution if and only if all odd prime divisors $p$ of $d$ satisfy $p\equiv 1\!\mod 4$.
1.6. Indecomposable Representations
The category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ satisfies the Krull-Remak-Schmidt Theorem (Lemma 2.14), and therefore one is reduced to understanding $\operatorname{Ind}_{L}^{\rtimes}(G)$, the set of isomorphism classes of indecomposable semilinear representations of $G$ over $L$. We similarly write $\operatorname{Ind}_{L}(H)$ for the set of isomorphism classes of indecomposable representations of $H$ over $L$.
Our next result describes $\operatorname{Ind}_{L}^{\rtimes}(G)$ in terms of $\operatorname{Ind}_{L}(H)$ and its action of $\operatorname{Gal}(L/K)$. To ease notation, we set $\Gamma\coloneqq\operatorname{Gal}(L/K)$, and for $W\in\operatorname{Ind}_{L}(H)$ write $\Gamma_{W}$ for the stabiliser of $W$ in $\Gamma$.
Along with the restriction functor
$(-)|_{H}\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{L}(H),$
we can, denoting by $L\rtimes G$ the skew group ring (Definition 2.9), consider the induction functor
$\operatorname{Ind}_{H}^{G}(-)\coloneqq(L\rtimes G)\otimes_{L[H]}-\colon\operatorname{\mathbf{Rep}}_{L}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G).$
Theorem B.
Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ and $W\in\operatorname{Ind}_{L}(H)$ is a direct summand of $V|_{H}$. Then:
  • $V|_{H}$ and $\operatorname{Ind}_{H}^{G}(W)$ are described by
    $V|_{H}\cong\bigoplus_{\gamma\in\Gamma/\Gamma_{W}}(\gamma*W)^{m(V)},\qquad\operatorname{Ind}_{H}^{G}(W)\cong V^{\oplus|\Gamma_{W}|/m(V)}$
    for some integer $m(V)\geq 1$ with $m(V)\mid|\Gamma_{W}|$.
  • $V$ is irreducible if and only if $W$ is irreducible.
  • $\operatorname{Ind}_{H}^{G}(W)\cong\operatorname{Ind}_{H}^{G}(\gamma*W)$ for any $\gamma\in\Gamma$.
  • Any $W\in\operatorname{Ind}_{L}(H)$ is a direct summand of $V|_{H}$ for a unique $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$.
  • This defines bijective correspondences
    $\operatorname{Ind}_{L}^{\rtimes}(G)\leftrightarrow\operatorname{Ind}_{L}(H)/\hskip 1.00006pt\Gamma,\qquad\operatorname{Irr}_{L}^{\rtimes}(G)\leftrightarrow\operatorname{Irr}_{L}(H)/\hskip 1.00006pt\Gamma.$
In particular, this gives a complete description of $\operatorname{Ind}_{L}^{\rtimes}(G)$, up to a knowledge of the numbers $m(V)$: for any orbit of $\operatorname{Gal}(L/K)$ on $\operatorname{Ind}_{L}(H)$, $m(V)$ copies of the orbit sum extends uniquely to an indecomposable semilinear representation of $G$ over $L$, and all elements of $\operatorname{Ind}_{L}^{\rtimes}(G)$ arise this way.
1.7. Semilinear Schur Index
Theorem B allows us to define the following.
Definition 1.3.
For $W\in\operatorname{Ind}_{L}(H)$, set $m_{K}^{L}(W)\coloneqq m(V)$ for the unique corresponding $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$.
We refer to $m_{K}^{L}(W)$ as the semilinear Schur index of $W$, as these generalise the classical Schur indices in an appropriate sense (see Section 10). In light of Theorem B, these measure the failure for a $\Gamma$-fixed $W\in\operatorname{Ind}_{L}(H)$ to extend to a semilinear representation.
Corollary 1.4.
$W\in\operatorname{Ind}_{L}(H)$ extends to a semilinear representation of $G$ over $L$ if and only if $\Gamma$ fixes $W$ and $m_{K}^{L}(W)=1$.
In Section 9 we establish the basic properties of the $m_{K}^{L}(W)$, including the fact that all $m_{K}^{L}(W)=1$ whenever $W\in\operatorname{Irr}_{L}(H)$ and $L$ is a finite field. In general, like their classical counterparts, the numbers $m_{K}^{L}(W)$ are hard to determine and often contain arithmetic information. For example, for $W$ as in Example 1.1, $m_{K}^{L}(W)\in\{1,2\}$, and $m_{K}^{L}(W)=1$ if and only if $x^{2}-dy^{2}=-1$ has a solution over $\mathbb{Q}$}.
1.8. Character Theory
When $G$ is finite and $L$ has characteristic 0, character theory provides a powerful and computationally efficient tool for completely describing the category $\operatorname{\mathbf{Rep}}_{L}(G)$.
Theorem A allows us to give a good definition of the character for $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ in this setting, which recovers the usual character theory for $\operatorname{\mathbf{Rep}}_{L}(G)$ when $G$ acts trivially on $L$.
Definition 1.5.
For $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, we set $\chi_{V}\coloneqq\chi_{V|_{H}}$, the character of $V|_{H}$.
As an immediate consequence of Theorem A we have the following properties.
Corollary 1.6.
Suppose that $G$ is finite and $L$ has characteristic 0. If $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, then:
  • $\chi_{V}=\chi_{W}$ if and only if $V\cong W$,
  • $\langle\chi_{V},\chi_{W}\rangle=\dim_{K}\operatorname{Hom}_{L\rtimes G}(V,W)$.
Here the inner product is the usual inner product of characters of $H$ over $L$:
$\langle\chi_{V},\chi_{W}\rangle=\frac{1}{|H|}\sum_{h\in H}\chi_{V}(h)\chi_{W}(h^{-1}).$
More generally, when $G$ is finite, $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple if and only if $|H|\in L^{\times}$ (Corollary 8.21), and in this context one can still use character theory to understand $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ (Section 12).
For example, this gives a complete description of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ using only characters when $L$ is a finite field, as in this case all $m_{K}^{L}(W)=1$.
As an example of how this character theory works, we give a complete classification of the semilinear representations of $S_{3}$ acting on a general degree two Galois extension $L/K$ via the natural projection $S_{3}\twoheadrightarrow\operatorname{Gal}(L/K)$. This is first done by hand in Section 5, in the case of characteristic 0, with explicit semilinear matrix computations. We then return to this example later in Section 13 to recover this classification more conceptually using characters and Theorems A and B.
1.9. Realisation of Division Algebras
For $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, the $K$-algebra $D\coloneqq\operatorname{End}_{L\rtimes G}(V)$ is a division algebra over $K$, which may or may not be central over $K$. When $\operatorname{End}_{L[H]}(W)=L$, for any corresponding $W\in\operatorname{Irr}_{L}(H)$, then $m(V)$ is the degree of $D$ in $\operatorname{Br}(Z(D))$, and $Z(D)=K$ exactly when $W$ is fixed by $\Gamma$.
In Section 16 we prove that this process exhausts $\operatorname{Br}(K)$: for any central division algebra $D$ over a characteristic 0 field $K$, we show that $D$ is realised as an endomorphism ring of some $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, for some finite group $G$ and finite Galois extension $L/K$. We further have that $m(V)=\operatorname{Deg}(D)$, which shows that there are no restrictions on the possible values of the $m(V)$ in general.
This is in contrast to what happens for for linear representations: the classes of $\operatorname{Br}({\mathbb{Q}})$ realised as endomorphism rings of linear representations of finite groups over $\mathbb{Q}$} are exactly those classes defined by cyclotomic algebras, which form a proper subgroup of $\operatorname{Br}({\mathbb{Q}})$.
1.10. Determination of Semilinear Schur Indices
By Theorem B, to understand $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ one is reduced to understanding the semilinear Schur indices of $W\in\operatorname{Ind}_{L}(H)$. This is hard in general.
For classical Schur indices, one has Unger’s algorithm [UNG] for computing Schur indices over number fields, and it would be interesting to see if this can be generalised to semilinear Schur indices. There is still a local-global principle here (Corollary 11.5), and so the problem can similarly be reduced to a problem for local fields. However many results upon which the algorithm rests require deep facts about classical Schur indices, which at present are not established for semilinear Schur indices.
1.10.1. Other Associated Linear Categories
One natural way one might try to glean information about the semilinear Schur indices $m_{K}^{L}(W)$ is to use other linear representation categories associated with $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Theorem B and the theory developed above can be viewed as a partial description of $\operatorname{Ind}_{L}^{\rtimes}(G)$ in terms of $\operatorname{Ind}_{L}(H)$, but there is another linear representation category that one might naturally consider, namely $\operatorname{\mathbf{Rep}}_{K}(G)$.
We can produce semilinear representations using the induction functor
$L\otimes_{K}-\colon\operatorname{\mathbf{Rep}}_{K}(G)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G),$
which is most simply described in terms of matrices: any matrix representation of $G$ over $K$ will define a semilinear matrix representation of $G$ over $L$, as all matrices are valued in $K$. Using this way to construct objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, this gives us more information about the numbers $m_{K}^{L}(W)$, but does not completely determine them (see Section 13).
1.10.2. Linear Characters
Unlike the classical Schur indices $\overline{m}_{K}^{L}(W)$, the semilinear Schur indices $m_{K}^{L}(W)$ of $1$-dimensional characters can be non-trivial. This is already illustrated in Example 1.1, for any $d$ for which $x^{2}-dy^{2}=-1$ has no solution over $\mathbb{Q}$}. In Section 15 we prove some non-trivial properties of $m_{K}^{L}(\chi)$ for $1$-dimensional representations $\chi$ (see Corollary 15.4).
1.10.3. The Case of $\mathbb{C}$$\mathbb{R}$}
Classically, the biggest success story for determining Schur indices is when $L/K$ is the extension $\mathbb{C}$$\mathbb{R}$} and $H$ is a finite group. One can associate to each $W\in\operatorname{Irr}_{{\mathbb{C}}}(H)$ a number
${\mathcal{F}}(W)\coloneqq\frac{1}{|H|}\sum_{h\in H}\chi_{W}(h^{2})\in\{-1,0,1\},$
known as the Frobenius-Schur indicator. It turns out that this number tells you everything: $W$ is fixed by $\operatorname{Gal}({\mathbb{C}}/{\mathbb{R}})$ if and only if ${\mathcal{F}}(W)\neq 0$, and such a $W$ is defined by a representation over $\mathbb{R}$} if and only if ${\mathcal{F}}(W)=1$. In particular, $\overline{m}_{{\mathbb{R}}}^{{\mathbb{C}}}(W)=2$ if ${\mathcal{F}}(W)=-1$, and $\overline{m}_{{\mathbb{R}}}^{{\mathbb{C}}}(W)=1$ otherwise.
Whilst the existence of such a simple way to determine the numbers $\overline{m}_{{\mathbb{R}}}^{{\mathbb{C}}}(W)$ is already remarkable, what is even more remarkable is that this theory extends to semilinear representations.
Given a finite group $G$, and a surjection $G\twoheadrightarrow\operatorname{Gal}({\mathbb{C}}/{\mathbb{R}})$ with kernel $H$, then for any $W\in\operatorname{Irr}_{{\mathbb{C}}}(H)$ one can define the generalised Frobenius-Schur indicator:
$\tilde{{\mathcal{F}}}(W)\coloneqq\frac{1}{|H|}\sum_{h\in G\setminus H}\chi_{W}(h^{2}).$
It turns out that, like before,
  • $\tilde{{\mathcal{F}}}(W)\in\{-1,0,1\}$,
  • $W$ is fixed by the (now twisted) action of $\operatorname{Gal}({\mathbb{C}}/{\mathbb{R}})$ if and only if $\tilde{{\mathcal{F}}}(W)\neq 0$,
  • $W$ extends to $\operatorname{\mathbf{Rep}}_{{\mathbb{C}}}^{\rtimes}(G)$ if and only if $\tilde{{\mathcal{F}}}(W)=1$ [RUMTAY, Thm. 4.2].
In particular, $m_{{\mathbb{R}}}^{{\mathbb{C}}}(W)=2$ if $\tilde{{\mathcal{F}}}(W)=-1$, and $m_{{\mathbb{R}}}^{{\mathbb{C}}}(W)=1$ otherwise.
An in-depth study of the numbers $m_{K}^{L}(W)$ in the smallest non-trivial case, when $L/K$ has degree two, is the subject of the forthcoming work [RUMTAY4], where alternatives to the generalised Frobenius-Schur indicator for determining the numbers $m_{K}^{L}(W)$ are explored.
1.11. Structure of the Paper
Let us now describe how the paper is organised, and give an overview of the strategy of the proof of Theorem A and Theorem B.
In Section 2, we recall some basic notions regarding semilinear representations, and how these can be interpreted in terms of matrices.
In Section 3 we fix our assumptions for the remainder of the paper. Crucially, even though we are ultimately interested in the case when $L/K$ is an extension of fields, we work with finite Galois extensions $L$ of a field $K$ that are potentially disconnected (so $L$ is a product of fields).
In Section 4 we study those semilinear representations arising from matrix representations defined over $K$, and give an extended example of how these can be used in some cases to describe all semilinear representations of $G$ over $L$ in Section 5, which also serves to demonstrate how one works with semilinear representations explicitly.
In Section 6, we consider one case in which the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is completely understood, which is when $H\hookrightarrow G$ is split. This recovers the theory of base change for linear representations.
In Section 7 we show that the semilinear representation category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, for $L$ a potentially disconnected finite Galois extension of $K$, is equivalent to $\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})$, where $F$ is the field defined by a connected component $\operatorname{Spec}(L)$, and $G_{e}$ is the stabiliser. This reduces general semilinear representation categories to the case when $L$ is a field, and is used in Section 8 to prove Theorems A and B.
In Section 8, we prove Theorems A and B. The key idea is to interpret the functors $(-)|_{H}$ and $\operatorname{Ind}_{H}^{G}$ between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$ as certain honest base change functors for the field extension $L/K$. This allows us to port many properties of base change to our setting, such as the preservation of semisimplicity, and how the functors $(-)|_{H}$ and $\operatorname{Ind}_{H}^{G}$ compose. Theorem A follows from Corollary 8.6, Remark 2.13, Remark 3.5, and Remark 8.9, and Theorem B follows from Remark 3.5 and Theorem 8.20. We also relate the integers $m(V)$ to the $K$-algebras $\operatorname{End}_{L\rtimes G}(V)$ in Proposition 8.29.
In Section 9 we define the semilinear Schur index of an object of $\operatorname{Ind}_{L}(H)$, and establish its basic properties, and in Section 10 we explain how the semilinear Schur indices specialise to the (classical) Schur indices in the case when $H\hookrightarrow G$ is split. We also highlight which properties the classical Schur indices satisfy which are not satisfied by all semilinear Schur indices.
In Section 11 we study the splitting behaviour of the $K$-division algebras $\operatorname{End}_{L\rtimes G}(V)$ for $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and consider base change for semilinear representations. When $K$ is a number field we use the local-global principle for $\operatorname{Br}(K)$ to prove a local-global principle for semilinear representations.
In Section 12 we reinterpret Theorems A and B in terms of characters, and in Section 13 we revisit the extended example of Section 5 and rederive the classification much more efficiently using this character theory. In Section 14 we specialise the framework slightly, and show that when the extension $L/K$ is cyclotomic there is also an action of $\operatorname{Gal}(L/K)$ on $G$, which allows one to compute the order $|\operatorname{Irr}_{L}^{\rtimes}(G)|$ purely in terms of the group theory of $G$.
In Section 15 we give a cohomological interpretation of our results in terms of Galois cohomology. In particular, we give a new description of the transgression map in our setting, which has consequences for the semilinear Schur indices of $1$-dimensional characters of $H$.
In Section 16 we show that all central division algebras over $K$ are realised as endomorphism rings of irreducible semilinear representations of $G$ over $L$, for some group $G$ and finite Galois extension $L$.
Finally, in Section 17 we indicate how the results of this paper extend to infinite Galois extensions.
1.12. Further Directions
Whilst we have defined the semilinear Schur indices $m_{K}^{L}(W)$ here and established their basic properties, it remains to better understand their structure. There is a wealth of literature regarding classical Schur indices, and it is natural to ask how much of this generalises.
For example, one obtains strong properties of the $\overline{m}_{K}^{L}(W)$ from Brauer’s induction theorem, and it would be natural to try and understand how semilinear induction theory works in this context.
It would also be interesting to work out an integral version of the theory, which in light of Example 1.1 should encode information about integral solutions of certain diophantine equations.
Acknowledgements
I would like to thank Tom Adams, Håvard Damm-Johnsen, Lorenzo La Porta and Dmitriy Rumynin for interesting conversations regarding the contents of this paper. This research was supported by an LMS Early Career Research Fellowship at the University of Cambridge and the departmental grant Progetto Sviluppo Dipartimentale - UNIPD PSDIP23O88 at the University of Padova.
Notation
[…]
2. Semilinear Representations
Suppose that $G$ is a group, $R$ is a commutative ring, and $\sigma\colon G\rightarrow\operatorname{Aut}(R)$ is group homomorphism, where $\operatorname{Aut}(R)$ denote the group of ring automorphisms of $R$.
The central objects we are interested in are semilinear representations of $G$.
Definition 2.1.
A semilinear representation of $G$ is a finite-rank free $R$-module $V$ with a group homomorphism $\rho\colon G\rightarrow\operatorname{Aut}_{{\mathbb{Z}}}(V)$ such that
$\rho(g)(\lambda\cdot v)=\sigma_{g}(\lambda)\cdot\rho(g)(v)\ \mbox{ for all }\ g\in G,\lambda\in R,v\in V.$
A morphism of semilinear representations from $V$ to $W$ is an $R$-linear map $\phi\colon V\rightarrow W$ such that
$\rho_{W}(g)\circ\phi=\phi\circ\rho_{V}(g)\ \mbox{ for all }\ g\in G.$
We denote the category of such representations by $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$.
Example 2.2.
When $\sigma$ is trivial, so $\sigma_{g}$ is the identity of $R$ for all $g\in G$, $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ is simply the category of finite-dimensional linear representations of $G$ over $R$. In this case we write omit $\sigma$ from the notation and write $\operatorname{\mathbf{Rep}}_{R}(G)$ for $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$.
We can perform certain constructions with semilinear representations, which generalise those of linear representations.
Example 2.3.
Suppose that $V,W\in\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$. Then
  • $V\oplus W$ is a semilinear representation of $G$, where $g\in G$ acts by
    $(v,w)\mapsto(\rho_{V}(g)(v),\rho_{W}(g)(w)).$
  • $V\otimes_{R}W$ is a semilinear representation of $G$, where $g\in G$ acts by
    $v\otimes w\mapsto\rho_{V}(g)(v)\otimes\rho_{W}(g)(w).$
  • $\operatorname{Hom}_{R}(V,W)$ is a semilinear representation of $G$, where $g\in G$ acts by
    $f\mapsto\rho_{W}(g)\circ f\circ\rho_{V}(g^{-1}).$
Example 2.4.
The trivial representation is $R$ with its natural action of $G$. This is irreducible if and only if $R$ has no proper non-trivial $G$-stable left ideals, and has endomorphism ring $R^{G}$.
Definition 2.5.
An $R$-submodule $W$ of a semilinear representation $V$ is a subrepresentation if
$\rho(g)(w)\in W\ \mbox{ for all }\ g\in G,w\in W.$
We say that $V$ is irreducible if $V$ is non-zero and 0 and $V$ are the only subrepresentations of $V$, and write $\operatorname{Irr}_{R}^{\hskip 0.81949pt\sigma}(G)$ for the set of isomorphism classes of irreducible objects of $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$.
We say that $V$ is indecomposable if $V$ is non-zero and $V$ cannot be written as the direct sum of two non-zero subrepresentations, and write $\operatorname{Ind}_{R}^{\hskip 0.81949pt\sigma}(G)$ for the set of isomorphism classes of indecomposable objects of $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$.
2.1. Matrix Form of Semilinear Representations
We can describe the category $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ more concretely in terms of matrices. We write $\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$ for the category of finite-rank free $R$-modules.
2.1.1. Semilinear Algebra
For any $\gamma\in\operatorname{Aut}(R)$ and $V\in\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$, we write $^{\gamma}V$ for $V$ as an abelian group but with $R$-module structure $\lambda*v=\gamma^{-1}(\lambda)v$. This defines a functor
$^{\gamma}(-)\colon\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}\rightarrow\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}},$
which is the identity on morphisms, and satisfies $^{\gamma}({}^{\mu}V)={}^{\gamma\mu}V$.
Let $V,W\in\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$ and fix bases $\mathcal{A}$} and $\mathcal{B}$} for $V$ and $W$ respectively, which induce an $R$-module identification
$\operatorname{End}_{R}(V,W)\rightarrow M_{m\times n}(R),\qquad[T\colon V\rightarrow W]\mapsto{}_{{\mathcal{B}}}[T]_{{\mathcal{A}}}.$
For any $\gamma\in\operatorname{Aut}(R)$, $\mathcal{A}$} and $\mathcal{B}$} also form bases for $^{\gamma}V$ and $^{\gamma}W$ respectively, and there is similarly an $R$-module identification
$\operatorname{End}_{R}({}^{\gamma}V,{}^{\gamma}W)\rightarrow M_{m\times n}(R),\qquad[T\colon V\rightarrow W]\mapsto{}_{{\mathcal{B}}}[T]_{{\mathcal{A}}},$
In fact, there is an equality of sets
$\operatorname{End}_{R}(V,W)=\operatorname{End}_{R}({}^{\gamma}V,{}^{\gamma}W),$
and under this equality the two identifications above are related by the commutativity of the diagram
$\operatorname{End}_{R}(V,W)$$M_{m\times n}(R)$$\operatorname{End}_{R}({}^{\gamma}V,{}^{\gamma}W)$$M_{m\times n}(R)$$\scriptstyle{\sim}$$\scriptstyle{=}$$\scriptstyle{{}^{\gamma}(-)}\scriptstyle{\sim}$
where $^{\gamma}(-)\colon M_{m\times n}(R)\rightarrow M_{m\times n}(R)$ is the ring automorphism defined by applying $\gamma$ to each entry.
Concretely, if $A\in M_{m\times n}(R)$ represents an $R$-linear map $T\colon V\rightarrow W$, then $^{\gamma\!}A$ represents the $R$-linear map $^{\gamma}T\colon{}^{\gamma}V\rightarrow{}^{\gamma}W$ (which is set-theoretically simply the map $T$).
Suppose now that $V,W\in\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$, and $S\colon V\rightarrow W$ is a $\gamma$-semilinear map, meaning that
$S(\lambda v)=\gamma(\lambda)S(v)\ \mbox{ for all }\ \lambda\in R,v\in V,$
or equivalently, that the map $S\colon{}^{\gamma}V\rightarrow W$ is $R$-linear. Let $B$ be the matrix associated to the $R$-linear map $S\colon{}^{\gamma}V\rightarrow W$ with respect to the bases $\mathcal{A}$} and $\mathcal{B}$}, and write [][-] for the identifications $V\xrightarrow{\sim}R^{n}$, $W\xrightarrow{\sim}R^{m}$ induced by these choices of basis. Then
$S(v)]=B\cdot{}^{\gamma}[v$],
where $^{\gamma}[v$] is $\gamma$ applied to all the entries of $[v]$.
2.1.2. Semilinear Matrix Representations
Suppose now that $V\in\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$. Then each $\rho(g)\colon V\rightarrow V$ is $\sigma_{g}$-semilinear and $\rho(g_{1}g_{2})=\rho(g_{1})\circ\rho(g_{2})$, so $V$ determines $R$-linear maps
$(\rho(g)\colon{}^{\sigma_{g}}V\rightarrow V)_{g\in G},$
which satisfy $\rho(g_{1}g_{2})=\rho(g_{1})\circ{}^{\sigma_{g_{1}}}\rho(g_{2})$. Conversely, given $V\in\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$, any such a family of $R$-linear maps defines a semilinear representation structure on $V$.
Let $n\coloneqq\text{rank}_{R}(V)$, and fix a basis of the $R$-module $V$. For any $g\in G$, write $A_{g}\in M_{n\times n}(R)$ for the matrix of the linear map $\rho(g)\colon{}^{\sigma_{g}}V\rightarrow V$. The compatibility relations above in matrix form become
(2) $A_{g_{1}g_{2}}=A_{g_{1}}\cdot{}^{\sigma_{g_{1}}\!}A_{g_{2}}\ \mbox{ for all }\ g_{1},g_{2}\in G.$
In particular, as $I_{n}=A_{gg^{-1}}=A_{g}\cdot{}^{\sigma_{g}\!}A_{g^{-1}}$, each $A_{g}\in\operatorname{GL}_{n}(R)$. The following is direct to verify.
Proposition 2.6.
Suppose that $V\in\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$ has rank $n\geq 1$, and fix a basis of $V$.
Then taking the matrix of a semilinear map defines a canonical bijection between semilinear representation structures $\rho$ on $V$ and families $(A_{g})_{g\in G}$ in $\operatorname{GL}_{n}(R)$ satisfying (2). Under this bijection,
(3) $\rho(g)v]=A_{g}\cdot{}^{\sigma_{g}}[v]\ \mbox{ for all }\ g\in G,v\in V.$
In particular, an $R$-submodule $W$ of $V$ is a subobject of $V$ in $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ exactly when
$A_{g}\cdot{}^{\sigma_{g}}[w]\in W\ \mbox{ for all }\ g\in G,w\in W.$
Suppose that $W\in\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$ has rank $m\geq 1$, and fix a basis of $W$. Let $V$ has semilinear representation structure corresponding to $(A_{g})_{g\in G}$ in $\operatorname{GL}_{n}(R)$ and let $W$ have semilinear representation structure corresponding to $(B_{g})_{g\in G}$ in $\operatorname{GL}_{m}(R)$.
Then an $R$-linear morphism $V\rightarrow W$ corresponding to $M\in M_{m\times n}(R)$ is a morphism in $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ if and only if $M$ satisfies
(4) $B_{g}\cdot{}^{\sigma_{g}}M=M\cdot A_{g}\ \mbox{ for all }\ g\in G.$
Example 2.7.
Suppose that $V\in\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ has $R$-rank $n\geq 1$, and $\mathcal{A}$} and $\mathcal{B}$} are $R$-bases of $V$. Suppose that $(A_{g})_{g\in G}$ and $(B_{g})_{g\in G}$ represent $V$ with respect to $\mathcal{A}$} and $\mathcal{B}$} respectively, and that $P\in\operatorname{GL}_{n}(R)$ is the change of basis matrix from $\mathcal{A}$} to $\mathcal{B}$}, meaning that $P\circ[-]_{{\mathcal{A}}}=[-]_{{\mathcal{B}}}$. Then
[…]
Example 2.8.
We can give matrix descriptions of the constructions of Examples 2.3 and 2.4. The trivial representation has matrix representation by $(I_{1})_{g\in G}$, for $I_{1}$ the identity of $M_{1}(R)$, and if $V$ and $W$ in $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ have matrix representations $(A_{g})_{g\in G}$ and $(B_{g})_{g\in G}$ respectively, then
  • $V\oplus W$ is represented by the block sum $(A_{g}\oplus B_{g})_{g\in G}$,
  • $V\otimes_{R}W$ is represented by the Kronecker product $(A_{g}\otimes B_{g})_{g\in G}$,
  • $\operatorname{Hom}_{R}(V,R)$ is represented by $((A_{g}^{-1})^{T})_{g\in G}$.
2.2. Semilinear Representations and Twisted Group Rings
Semilinear representations can also be viewed as modules over twisted group algebras. Suppose as above that $\sigma\colon G\rightarrow\operatorname{Aut}(R)$ is a group homomorphism.
Definition 2.9.
The ring $R\rtimes_{\sigma}\!G$ is $R[G]$ as an abelian group, but with multiplication
$(r_{1}\cdot g_{1})*(r_{2}\cdot g_{2})\coloneqq r_{1}\sigma_{g_{1}}(r_{2})\cdot g_{1}g_{2}\ \mbox{ for all }\ r_{1},r_{2}\in R,\ g_{1},g_{2}\in G,$
extended to all of $R[G]$ by linearly.
Note that $R$ is not central in $R\rtimes_{\sigma}\!G$, but $R^{G}$ is. We therefore naturally view $R\rtimes_{\sigma}\!G$ as an $R^{G}$-algebra. The following is well-known, and direct the verify.
Proposition 2.10.
Suppose that $V\in\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$. Then there is a canonical bijection between $R\rtimes_{\sigma}\!G$-module structures on $V$ and semilinear structures on $V$, under which a morphism in $\operatorname{\mathbf{Mod}}_{R}^{\operatorname{ff}}$ is a morphism in $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ if and only if it is a morphism in $\operatorname{\mathbf{Mod}}_{R\rtimes_{\sigma}\hskip-0.81949ptG}$.
In particular, we see that $\operatorname{\mathbf{Rep}}_{R}^{\hskip 0.81949pt\sigma}(G)$ is canonically identified with the full subcategory of $\operatorname{\mathbf{Mod}}_{R\rtimes_{\sigma}\hskip-0.81949ptG}$ where the underlying $R$-module is free of finite rank.
2.3. Finite Products of Fields
In this paper we are ultimately interested in the case when $R$ is a Galois extension of a field and $G$ acts through the Galois action. We first consider the more general situation when $R$ is any finite product of fields.
Therefore, suppose until the end of this section that $L=\prod_{i}L_{i}$ is a product of fields with an action $\sigma\colon G\rightarrow\operatorname{Aut}(L)$ of a group $G$ such that the induced action of $G$ on the set $\{e_{i}\}_{i}$ of principal idempotents is transitive. We note that this implies $K:=L^{G}$ is a field by [TAY4, Lem. 3.2], taking the triple $(F,G,H)$ of [TAY4, §3.1] to be $(L,G,G)$.
Proposition 2.11.
Suppose that the index set $I$ of the product $L=\prod_{i}L_{i}$ is finite. Then any $L\rtimes G$-module is free as an $L$-module. If additionally $G$ is finite and $|G|\in L^{\times}$, then $\operatorname{\mathbf{Rep}}_{L}^{\hskip 0.81949pt\sigma}(G)$ is semisimple.
Proof.
The first claim follows from [TAY4, Rem. 2.5], as $|I|<\infty$ and the natural homomorphism $V\rightarrow\prod_{i}V_{i}$ is an isomorphism for any $L\rtimes G$-module $V$ (because $|I|<\infty$). Then the second claim follows from [MONT2, Cor. 0.2], because $|G|\in L^{\times}$ and any $L\rtimes G$-module $V$ has an $L$-module complement. ∎
Lemma 2.12.
If $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\hskip 0.81949pt\sigma}(G)$, then the natural map
(5) $L\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\rightarrow\operatorname{Hom}_{L[H]}(V,W)$
is injective, where $K=L^{G}$ and $H=\ker(G\rightarrow\operatorname{Aut}(L/K))$. Furthermore,
[…]
and (5) is an isomorphism if and only if this is an equality.
Remark 2.13.
This injection is compatible with composition of morphisms in the obvious sense. In particular, when $V=W$ this injection is a ring homomorphism
$L\otimes_{K}\operatorname{End}_{L\rtimes G}(V)\hookrightarrow\operatorname{End}_{L[H]}(V).$
Proof.
Firstly, because $V$ and $W$ are free of finite rank over $L$, $N\coloneqq\operatorname{Hom}_{L}(V,W)$ is also free of finite rank over $L$. Furthermore, because $H$ acts trivially on $L$, the isomorphism
$N\xrightarrow{\sim}\prod_{i}e_{i}\cdot N,$
restricts to an isomorphism
$N^{H}\xrightarrow{\sim}\prod_{i}e_{i}\cdot N^{H}.$
In particular, $V\coloneqq N^{H}$ if free of finite rank over $L$ by [TAY4, Lem. 2.4]. Therefore we may apply [TAY4, Cor. 3.8] with the $(F,G,H)$ of loc. cit. taken to be $(L,G,G)$ to deduce the required result, noting that
$V^{G}=\operatorname{Hom}_{L[H]}(V,W)^{G}=\operatorname{Hom}_{L\rtimes G}(V,W).\qed$
Lemma 2.14.
Suppose that the index set $I$ of the product $L=\prod_{i}L_{i}$ is finite. Then the Krull-Remak-Schmidt Theorem holds in $\operatorname{\mathbf{Rep}}_{L}^{\hskip 0.81949pt\sigma}(G)$: any non-zero $V\in\operatorname{\mathbf{Rep}}_{L}^{\hskip 0.81949pt\sigma}(G)$ can be written as a direct sum of indecomposable objects
$V\cong V_{1}\oplus\cdots\oplus V_{n},$
and any such description is unique up to isomorphism and permutation.
Proof.
This follows from [ATI, Thm. 1], using [ATI, §3 Cor.] and the fact that $\operatorname{Hom}(V,W)$ is finite dimensional over $K$ for any $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\hskip 0.81949pt\sigma}(G)$ because $|I|<\infty$. ∎
3. Semilinear Representations for Galois Extensions
3.1. Galois Extensions of a Field
Definition 3.1.
If $K$ is a field, and $L$ is a finite-dimensional $K$-algebra with an action of a finite group $\Gamma$, we say that $L/K$ is a Galois extension with Galois group $\Gamma$ if $K=L^{\Gamma}$ and the natural map
$L\otimes_{K}L\rightarrow\prod_{\gamma\in\Gamma}L,\qquad a\otimes b\mapsto(a\gamma(b))_{\gamma\in\Gamma}$
is an isomorphism [TAY3, Def. 2.47].
Note that unlike the definition of a Galois field extension, the action of $\Gamma$ is part of the data of a Galois extension. However, the following shows that when $L$ is a field this essentially coincides with the notion of a Galois extension of fields.
Example 3.2.
Any Galois field extension $L/K$ with its natural action of $\operatorname{Gal}(L/K)$ is a Galois extension with Galois group $\operatorname{Gal}(L/K)$. Conversely, if $L/K$ is a Galois extension with Galois group $\Gamma$ and $L$ is a field, then $L/K$ is Galois and the action map $\Gamma\rightarrow\operatorname{Gal}(L/K)$ is an isomorphism.
Example 3.3.
For a Galois extension $L/K$, $\operatorname{Spec}(L)$ need not be connected, or equivalently, $L$ need not be a field. For example, for any finite group $\Gamma$ one has the split Galois extension defined by setting $K_{\gamma}=K$ for any $\gamma\in\Gamma$, with diagonal inclusion
$K\hookrightarrow\prod_{\gamma\in\Gamma}K_{\gamma}$
and action of $\Gamma$ by permutation $\sigma\cdot(\lambda_{\gamma})_{\gamma}\coloneqq(\lambda_{\sigma\gamma})_{\gamma}$. This is a Galois extension with Galois group $\Gamma$.
Remark 3.4.
Generally, if $L/K$ is a Galois extension with Galois group $\Gamma$, then for any connected component $\operatorname{Spec}(L_{i})$ of $\operatorname{Spec}(L)$, $L_{i}/K$ is a Galois field extension with Galois group $G_{i}\coloneqq\operatorname{Stab}_{\Gamma}(\operatorname{Spec}(L_{i}))$, and the action of $\Gamma$ on $\pi_{0}(\operatorname{Spec}(L))$ is transitive [TAY3, Lem. 2.52].
3.2. Assumptions and Notation
From now on until the end of the paper we suppose that:
  1. (1)
    $K$ is a field, and $K\rightarrow L$ is a Galois extension with Galois group $\Gamma$,
  2. (2)
    $G$ is a group which acts on $L$ through a surjective homomorphism $\sigma\colon G\rightarrow\Gamma$ with kernel $H$.
As $\sigma\colon G\rightarrow\Gamma$ is fixed, we suppress it from the notation and write:
  • $g(\lambda)\coloneqq\sigma_{g}(\lambda)$ for $\lambda\in L,g\in G$,
  • $L\rtimes G\coloneqq L\rtimes_{\sigma}\!G$,
  • $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ for $\operatorname{\mathbf{Rep}}_{L}^{\hskip 0.81949pt\sigma}(G)$,
  • $\operatorname{Irr}_{L}^{\rtimes}(G)$ for $\operatorname{Irr}_{L}^{\hskip 0.81949pt\sigma}(G)$.
We also fix the following notation:
  • $e$ is a primitive idempotent of $L$,
  • $F\coloneqq e\cdot L$,
  • $\Gamma_{e}\coloneqq\operatorname{Stab}_{\Gamma}(e)$, and
  • $G_{e}\coloneqq\operatorname{Stab}_{G}(e)$.
From Remark 3.4, the induced action of $\Gamma_{e}$ on $F$ makes $F/K$ a Galois extension with Galois group $\Gamma_{e}$, and the action of $\Gamma$ on the connected components of $\operatorname{Spec}(L)$ is transitive. In particular, the tuple
$(G_{e},H,F/K,\sigma\colon G_{e}\rightarrow\Gamma_{e})$
also satisfies the above assumptions, with the additional property that $F$ is a field.
Remark 3.5.
Suppose momentarily that $L$ is a field and $G$ is a group that acts on $L$ through a group homomorphism $G\rightarrow\operatorname{Aut}(L)$ (as in the introduction). Setting $K\coloneqq L^{G}$, then the assumption that $L/K$ is a finite extension automatically implies that $L/K$ is Galois and that the induced map $G\rightarrow\operatorname{Gal}(L/K)$ is surjective (and therefore satisfies the assumptions of this section). Indeed, $K$ is the fixed field of a subgroup of the finite group $\operatorname{Aut}(L/K)$, so $L/K$ is Galois, and if $S\subset\operatorname{Gal}(L/K)$ denotes the image of $G$ then $L^{S}=L^{G}=K$, hence $S=\operatorname{Gal}(L/K)$ by the Galois correspondence.
We also note that this setup is stable under base change, which is an important tool for our analysis, and why we don’t consider only Galois extensions of fields.
More precisely, if $K^{\prime}/K$ is a field extension, then the base change $L^{\prime}=L\otimes_{K}K^{\prime}$ with is a Galois extension of $K^{\prime}$ with Galois group $\Gamma$, where $\Gamma$ acts through the left factor of $L\otimes_{K}K^{\prime}$. With the induced action of $G$ on $L^{\prime}$ the tuple
$(G,H,L^{\prime}/K^{\prime},\sigma\colon G\rightarrow\Gamma)$
also satisfies the assumptions of this section.
Lemma 3.6.
If $K^{\prime}/K$ is a field extension and $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, then the natural map
$K^{\prime}\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\rightarrow\operatorname{Hom}_{L^{\prime}\rtimes G}(V_{L^{\prime}},W_{L^{\prime}})$
is an isomorphism.
Proof.
As an $L$-module $V$ is free of finite rank, hence by [BOUR, Ch. 2, §5.3, Prop. 7(ii)] the natural map
$K^{\prime}\otimes_{K}\operatorname{Hom}_{L}(V,W)\equiv L^{\prime}\otimes_{L}\operatorname{Hom}_{L}(V,W)\rightarrow\operatorname{Hom}_{L^{\prime}}(V_{L^{\prime}},W_{L^{\prime}})$
is an isomorphism, and we are done after taking $G$-invariants. ∎
We are interested in describing the semilinear representation category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ in terms of associated linear representation categories, in particular $\operatorname{\mathbf{Rep}}_{F}(H)$. Let us now examine these categories.
4. Associated Linear Representation Categories
There is a diagram of $K$-algebra inclusions
[…]
Recall that an homomorphism of (unital associative) rings $R\rightarrow S$ is called Frobenius if the induction
$S\otimes_{R}-\colon\operatorname{\mathbf{Mod}}_{R}\rightarrow\operatorname{\mathbf{Mod}}_{S}$
and co-induction
$\operatorname{Hom}_{R}(S,-)\colon\operatorname{\mathbf{Mod}}_{R}\rightarrow\operatorname{\mathbf{Mod}}_{S}$
functors are naturally isomorphic.
Proposition 4.1.
All inclusions in the diagram (6) are Frobenius.
Proof.
We use the criterion of [KAD, Thm. 1.2], that $R\rightarrow S$ is Frobenius if and only if there is an $(R,R)$-bimodule homomorphism $E\colon S\rightarrow R$ and elements $x_{1},...,x_{n},y_{1},...,y_{n}\in S$ with
(7) $\sum_{i=1}^{n}E(sx_{i})y_{i}=s=\sum_{i=1}^{n}x_{i}E(y_{i}s)$
for all $s\in S$.
To define these maps and elements in our context, let $g_{1},...,g_{n}$ be a set of left coset representatives for $H$ in $G$. For $K[H]\hookrightarrow K[G]$ or $L[H]\hookrightarrow L\rtimes G$, one can take $E\colon S\rightarrow R$ to be the $K$-linear (resp. $L$-linear) extension of the map which has $E(h)=h$ for $h\in H$ and $E(g)=0$ if $g\not\in H$, with $x_{i}=g_{i}$ and $y_{i}=g_{i}^{-1}$. For $K[H]\hookrightarrow L[H]$ or $K[G]\hookrightarrow L\rtimes G$, let $E\colon S\rightarrow R$ be the linear extension of $E(\lambda g)=\sum_{\gamma\in\Gamma}\gamma(\lambda)g$. The $\{x_{i}\}_{i}$ and $\{y_{i}\}_{i}$ are those given by [CHR, Thm. 1.3(b)]. It is direct to verify that in each case $E$ is $(R,R)$-bilinear and with the $\{x_{i}\}_{i}$ and $\{y_{i}\}$ satisfying the identity (7). ∎
The induction and restriction functors between the rings of diagram (6) restrict to functors between the categories
$\operatorname{\mathbf{Rep}}_{K}(G)$$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{K}(H)$$\operatorname{\mathbf{Rep}}_{L}(H)$}}
and therefore one can naturally use induction from the linear representation categories $\operatorname{\mathbf{Rep}}_{K}(G)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$ to construct objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$.
4.1. Semilinear Representations from $\operatorname{\mathbf{Rep}}_{K}(G)$
In this section we consider the most natural way to construct semilinear representations, and consider semilinear representations arising from $\operatorname{\mathbf{Rep}}_{K}(G)$. The base change functor $(L\rtimes G)\otimes_{K[G]}$- is naturally isomorphic to the functor
$L\otimes_{K}-\colon\operatorname{\mathbf{Rep}}_{K}(G)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$
defined for $V\in\operatorname{\mathbf{Rep}}_{K}(G)$ by letting $g\in G$ act diagonally on $L\otimes_{K}V$. The isomorphism is the composition of identifications
$L\otimes_{K}V\equiv L\otimes_{K}K[G]\otimes_{K[G]}V\equiv(L\rtimes G)\otimes_{K[G]}V,$
the composition of which is given by $\lambda\otimes v\mapsto\lambda\otimes v$. This is $L\rtimes G$-linear, as in $(L\rtimes G)\otimes_{K[G]}V$,
$g\cdot(\lambda\otimes v)=g(\lambda)g\otimes v=g(\lambda)\otimes g(v).$
Concretely, in terms of matrices (cf. Section 2.1), if $\rho\colon G\rightarrow\operatorname{GL}_{n}(K)$ corresponds to $V\in\operatorname{\mathbf{Rep}}_{K}(G)$, then $L\otimes_{K}V$ has matrix representation $\rho\colon G\rightarrow\operatorname{GL}_{n}(K)\hookrightarrow\operatorname{GL}_{n}(L)$. The relations of (2) are satisfied because all matrices are valued in $K$.
It turns out that the functor $L\otimes_{K}$- is neither fully faithful nor essentially surjective in general. However we can at least say the following.
Corollary 4.2.
For any irreducible $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$,
$V\hookrightarrow L\otimes_{K}(V|_{K}),$
hence any irreducible $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is a sub-object of an object in the essential image of $L\otimes_{K}$-.
Proof.
Because $K[G]\hookrightarrow L\rtimes G$ is Frobenius by Proposition 4.1, $L\otimes_{K}$- is right-adjoint to restriction, hence there is a $K$-linear isomorphism
$\operatorname{Hom}_{L\rtimes G}(V,L\otimes_{K}(V|_{K}))\equiv\operatorname{Hom}_{L\rtimes G}(V|_{K},V|_{K})\neq 0,$
and therefore, as $V$ is irreducible, $V\hookrightarrow L\otimes_{K}(V|_{K})$. ∎
5. Example: $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ for $L/K$ of Degree 22
In this section we give an example of how one can use the functor
$L\otimes_{K}-\colon\operatorname{\mathbf{Rep}}_{K}(G)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$
to construct semilinear representations. This example also illustrates how one practically works with semilinear matrices in terms of matrices, as described in Section 2.1. This example will also help to motivate the general theory we will develop in Section 8, using which we will rederive the results of this section more conceptually in Section 13.
Let $L/K$ be a degree two Galois field extension, and let $S_{3}$ act on $L$ via the non-trivial homomorphism $\sigma\colon S_{3}\twoheadrightarrow\operatorname{Gal}(L/K)$, which has kernel $H=\langle(123)\rangle$. Write $\alpha\colon L\rightarrow L$ for the non-trivial element of $\operatorname{Gal}(L/K)$. We restrict to characteristic 0 for simplicity.
Because $K$ has characteristic 0, the irreducible $K$-representations of $S_{3}$ are the trivial representation $\rho_{1}$, the sign representation $\rho_{2}$, and the unique two-dimensional irreducible representation $\rho_{3}$.
5.1. Representations from $\rho_{1}$ and $\rho_{2}$
Let us first understand the semilinear representations $L\otimes_{K}\rho_{1}$ and $L\otimes_{K}\rho_{2}$. $L\otimes_{K}\rho_{1}$ has matrix representation $A_{g}=1$ for all $g\in S_{3}$, and $L\otimes_{K}\rho_{2}$ has matrix representation $B_{g}=1$ if $g\in\langle(123)\rangle$, and $B_{g}=-1$ if $g\not\in\langle(123)\rangle$. By Proposition 2.6, an isomorphism between these representations would be a scalar $\lambda\in L^{\times}$ with the property that
$B_{g}\cdot{}^{\sigma_{g}}\lambda=\lambda A_{g}$
for all $g\in G$, or in other words, with $\alpha(\lambda)=-\lambda$. This can be found, as the $K$-linear operator $\alpha\colon L\rightarrow L$ has order $2$, and $L$ splits into eigenspaces
$L=\{x\in L\mid\alpha(x)=x\}\oplus\{x\in L\mid\alpha(x)=-x\}=K\oplus\{x\in L\mid\alpha(x)=-x\}.$
Therefore, we see that the two non-isomorphic representations of $S_{3}$ $\rho_{1}$ and $\rho_{2}$ become isomorphic as $L$-linear representations of $S_{3}$. Using Proposition 2.6 we can also compute
$\operatorname{End}_{L\rtimes S_{3}}(L\otimes_{K}\rho_{1})\subset L$
as those $\lambda\in L$ with $A_{g}\cdot{}^{\sigma_{g}}\lambda=\lambda A_{g}$ for all $g\in G$, or in other words, with $^{\sigma_{g}}\lambda=\lambda$ for all $g\in G$, which, because $\sigma$ is surjective, is simply $K$.
5.2. Decomposing $L\otimes_{K}\rho_{3}$
Now let us consider the irreducible two dimensional representation $\rho_{3}$. From the natural permutation representation with basis $e_{1},e_{2},e_{3}$ we may obtain $\rho_{3}$ as the quotient by the $L$-subspace spanned by $e_{1}+e_{2}+e_{3}$, which has matrix representation with respect to the basis $\overline{e_{1}},\overline{e_{2}}$ given by
$A_{(12)}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\qquad A_{(123)}=\begin{pmatrix}0&-1\\ 1&-1\end{pmatrix}$.
Because $\rho_{3}$ has dimension $2$, if $V\coloneqq L\otimes_{K}\rho_{3}$ is not irreducible then there is some $1$-dimensional $L$-subspace preserved by the action of $G$, and so some vector $\mathbf{v}\in L^{2}$ with
$A_{g}\cdot{}^{g}\mathbf{v}=\lambda_{g}\mathbf{v}$
for some $\lambda_{g}\in L^{\times}$, by Proposition 2.6. In particular, there are $\lambda,\mu\in L^{\times}$ with
$\begin{pmatrix}\lambda v_{1}\\ \lambda v_{2}\end{pmatrix}=A_{(12)}\cdot{}^{\alpha}\mathbf{v}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\cdot{}^{\alpha}\mathbf{v}=\begin{pmatrix}\alpha(v_{2})\\ \alpha(v_{1})\end{pmatrix}$,
and
$\begin{pmatrix}\mu v_{1}\\ \mu v_{2}\end{pmatrix}=A_{(123)}\mathbf{v}=\begin{pmatrix}0&-1\\ 1&-1\end{pmatrix}\mathbf{v}=\begin{pmatrix}-v_{2}\\ v_{1}-v_{2}\end{pmatrix}$.
From the second pair of equations, we have that $v_{2}(\mu^{2}+\mu+1)=0$, and from the first pair that $v_{2}\neq 0$ (otherwise $\mathbf{v}=\mathbf{0}$). Therefore $\mu^{2}+\mu+1=0$ and $L$ contains a primitive third root of $1$.
In particular, if $L$ contains no primitive third root of $1$, then $V$ is irreducible.
5.2.1. The case when $\omega\in K$
Suppose then, that $L$ contains $\omega$, a primitive third root of $1$. Then there are two cases. In the first case, where $\omega\in K$, we may diagonalise over $K$ the matrix of $(123)$ (which has characteristic polynomial $x^{2}+x+1$) to obtain a matrix representation of $\rho_{3}$ over $K$ defined by
$B_{(12)}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\qquad B_{(123)}=\begin{pmatrix}\omega&0\\ 0&\omega^{-1}\end{pmatrix}$.
In particular, similarly to above, if $V$ was reducible then one could obtain equations
$\begin{pmatrix}\alpha(v_{2})\\ \alpha(v_{1})\end{pmatrix}=\begin{pmatrix}\lambda v_{1}\\ \lambda v_{2}\end{pmatrix},\qquad\begin{pmatrix}\omega v_{1}\\ \omega^{-1}v_{2}\end{pmatrix}=\begin{pmatrix}\mu v_{1}\\ \mu v_{2}\end{pmatrix}$
for some $\lambda,\mu\in L^{\times}$. From the first pair of equations we must have that $v_{1},v_{2}$ are both non-zero, and therefore from the second equations we must have that $\omega=\mu=\omega^{-1}$, a contradiction. So also in this case $V$ is also irreducible.
5.2.2. The case when $\omega\in L$ and $\omega\not\in K$
The most interesting case is the third, when $\omega\in L$ but $\omega\not\in K$. In this case, $L=K(\omega)$, and the automorphism $\alpha\colon L\rightarrow L$ swaps $\omega$ and $\omega^{-1}$. We can then define $\mathbf{v}=(-1,\omega)$, $\mathbf{w}=(-1,\omega^{-1})$. The $L$-span of each defines a $G$-stable subspace because
$A_{(12)}\cdot{}^{\alpha}\mathbf{v}=\begin{pmatrix}\alpha(\omega)\\ \alpha(-1)\end{pmatrix}=\begin{pmatrix}\omega^{-1}\\ -1\end{pmatrix}=-\omega^{-1}\begin{pmatrix}-1\\ \omega\end{pmatrix}=\omega^{-1}\mathbf{v}$
and
$A_{(123)}\cdot\mathbf{v}=\begin{pmatrix}-\omega\\ -1-\omega\end{pmatrix}=\begin{pmatrix}-\omega\\ \omega^{2}\end{pmatrix}=\omega\begin{pmatrix}-1\\ \omega\end{pmatrix}=\omega\mathbf{v},$
and similarly for $\mathbf{w}$. We also see that these two subrepresentations restrict to the two non-trivial (linear) representations of $H$ over $L$, and therefore must be non-isomorphic in $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$.
5.3. Summary
To summarise:
  • $L\otimes_{K}\rho_{1}\cong L\otimes_{K}\rho_{2}$, with endomorphism ring $K$,
  • If there is some $\omega\in L$ which is a primitive third root of $1$ and $\omega\not\in K$, then $L\otimes_{K}\rho_{3}$ is the direct sum of two non-isomorphic $1$-dimensional irreducible representations,
  • Otherwise $V=L\otimes_{K}\rho_{3}$ is irreducible.
By Corollary 4.2 we know that these are all the irreducible representations, and that any object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ is a direct sum of these irreducible representations by Proposition 2.11.
5.4. Endomorphism Rings
We can also compute endomorphism rings for these representations.
First note that each irreducible semilinear representation $\sigma_{i}$ of $S_{3}$ over $L$ corresponds to a matrix factor in the $K$-algebra $L\rtimes S_{3}$, which is semisimple by Proposition 2.11. This is of the form $M_{n_{i}}(D_{i})$ where $D_{i}\coloneqq\operatorname{End}_{L\rtimes S_{3}}(\sigma_{i})$. The simple module $\sigma_{i}$ is recovered from the action of $L\rtimes S_{3}$ on the first column, and so by comparing $K$-dimensions we have that
$n_{i}\cdot\dim_{K}D_{i}=[L:K]\cdot\dim_{L}\sigma_{i}.$
5.4.1. The case when $\omega\in L\setminus K$.
In the case where there is a primitive third root $\omega\in L$ but $\omega\not\in K$, we have seen that there are two other non-isomorphic representations $\sigma_{2}$ and $\sigma_{3}$ with $\sigma_{2}\oplus\sigma_{3}=V$. There is an injection $L\otimes_{K}D_{i}\hookrightarrow\operatorname{End}_{L[H]}(\sigma_{i})$ from Lemma 2.12, and each $\operatorname{End}_{L[H]}(\sigma_{i})=L$ as each $\sigma_{i}$ restricted to $L[H]$-module is absolutely irreducible. In particular, $D_{2}=K=D_{3}$, each $n_{i}=n_{i}\dim_{K}D_{i}=[L:K]\dim_{L}\sigma_{i}=2$, and we have a decomposition
$L\rtimes S_{3}\cong M_{2}(K)\times M_{2}(K)\times M_{2}(K).$
5.4.2. The case when $V$ is irreducible
In the other case, where $V$ is irreducible, the above expression reduces to
$4=n_{V}\dim_{K}D_{V},$
so $n_{V}=2$, and $\dim_{K}D_{V}=2$. In particular, $D_{V}$ is a degree $2$ field extension of $K$ and
$L\rtimes S_{3}\cong M_{2}(K)\times M_{2}(D_{V}).$
We can be more precise about $D_{V}$, using the injection $L\otimes_{K}D_{V}\hookrightarrow\operatorname{End}_{L[H]}(V)$ from Lemma 2.12. If $\omega\in L$ (and hence $\omega\in K$), then $\operatorname{End}_{L[H]}(V)=L\times L$, so $L\otimes_{K}D_{V}\xrightarrow{\sim}L\times L$ and thus $D_{V}\cong L$ by [COHN, Thm. 2.2]. Conversely, if $\omega\not\in L$, then $\operatorname{End}_{L[H]}(V)=L(\omega)$, and all we can say in this level of generality is that $L\otimes_{K}D_{V}\xrightarrow{\sim}L(\omega)$.
6. Split Extensions
In Section 8 we will describe the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ in terms of $\operatorname{\mathbf{Rep}}_{F}(H)$. Before this, we describe one situation in which we can already give a complete description of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, namely when the sequence
$1\rightarrow H\rightarrow G\rightarrow\Gamma\rightarrow 1$
is left-split, so that $G\cong H\times\Gamma$ and $H\rightarrow G$ and $G\rightarrow\Gamma$ are the natural inclusion and projection.
Proposition 6.1.
Suppose that $H$ is a group and $L/K$ is a Galois extension with Galois group $\Gamma$. Then the functors
$L\otimes_{K}$- $\colon\operatorname{\mathbf{Rep}}_{K}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\Gamma),$
$(-)^{\Gamma}$ $\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\Gamma)\rightarrow\operatorname{\mathbf{Rep}}_{K}(H)$
define mutually inverse equivalences of categories.
Proof.
This is just Galois descent, see for example [TAY3, Prop. 2.53]. ∎
The following lemma shows that under these identifications, restriction from $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\Gamma)$ to $\operatorname{\mathbf{Rep}}_{L}(H)$ corresponds to induction from $\operatorname{\mathbf{Rep}}_{K}(H)$ to $\operatorname{\mathbf{Rep}}_{L}(H)$.
Lemma 6.2.
Suppose that $H$ is a group and $L/K$ is a Galois extension with Galois group $\Gamma$. Then the diagram
$\operatorname{\mathbf{Rep}}_{K}(H)$$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\Gamma)$$\operatorname{\mathbf{Rep}}_{L}(H)$$\scriptstyle{L\otimes_{K}-}$$\scriptstyle{(-)_{L}}$$\scriptstyle{(-)^{\Gamma}}$$\scriptstyle{(-)|_{H}}$
commutes up to natural isomorphism, where $H\times\Gamma$ acts on $L$ through the projection to $\Gamma$.
Proof.
For $V\in\operatorname{\mathbf{Rep}}_{K}(H)$, the two actions of $L[H]$ on $L\otimes_{K}V$ coincide. ∎
Similarly, induction from $\operatorname{\mathbf{Rep}}_{L}(H)$ to $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ corresponds to restriction from $\operatorname{\mathbf{Rep}}_{L}(H)$ to $\operatorname{\mathbf{Rep}}_{K}(H)$.
Lemma 6.3.
Suppose that $H$ is a group and $L/K$ is a Galois extension with Galois group $\Gamma$. Then the diagram
[…]
commutes up to natural isomorphism, where $H\times\Gamma$ acts on $L$ through the projection to $\Gamma$.
Proof.
Suppose that $V\in\operatorname{\mathbf{Rep}}_{L}(H)$. We have $K[H]$-linear isomorphisms
[…]
defined by $\lambda\otimes v\mapsto(\gamma(\lambda)v)_{\gamma}$, $(v_{\gamma})_{\gamma}\mapsto\gamma^{-1}\otimes v_{\gamma}$ and $x\otimes v\mapsto x\otimes v$ respectively, where the first map comes from applying $\otimes_{L}V$ to the defining isomorphism of the Galois extension $L/K$. It is direct to check that this composition commutes with the action of $L$ and $\Gamma$, and thus is $L\rtimes(H\times\Gamma)$-linear. ∎
7. Connected Components
In this section, we show that the study of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ reduces to the case that $L$ is connected. The results of this section are used crucially in Section 8 when we describe the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ in terms of $\operatorname{\mathbf{Rep}}_{F}(H)$ and prove Theorems A and B.
Recall that $e$ denotes a primitive idempotent of $L$, $F=e\cdot L$, and $G_{e}=\operatorname{Stab}_{G}(e)$. From Remark 3.4, the induced action of $G_{e}$ on $F$ makes $F/K$ a Galois extension with Galois group $G_{e}$, and the action of $G$ on the connected components of $\operatorname{Spec}(L)$ is transitive.
There is a functor
$e\cdot(-)\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e}),$
and a functor in the other direction defined by
$(L\rtimes G)\otimes_{L\rtimes G_{e}}-\colon\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G),$
where we view any $F\rtimes G_{e}$-module as a $L\rtimes G_{e}$-module via the natural projection $L\rtimes G_{e}\rightarrow F\rtimes G_{e}$.
Proposition 7.1.
The functors
$e\cdot(-)$ $\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e}),$
$(L\rtimes G)\otimes_{L\rtimes G_{e}}$- $\colon\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$
define mutually inverse equivalences of categories.
Proof.
Suppose first that $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and define
$\phi\colon(L\rtimes G)\otimes_{L\rtimes G_{e}}e\cdot V\rightarrow V,\qquad\phi(\lambda g\otimes v)=\lambda g(v).$
To see that this gives a well-defined map out of the tensor product, let $\mu h\in L\rtimes G_{e}$ and note that
$\phi((\lambda g)(\mu h)\otimes v)=\phi(\lambda g(\mu)gh\otimes v)=\lambda g(\mu)(gh)(v),$
which, noting that $e\cdot v=v$ and hence
$(gh)(v)=(gh)(e\cdot v)=g(h(e)\cdot h(v))=g(e\cdot h(v))=g(e)\cdot(gh)(v),$
is equal to
$\phi(\lambda g\otimes(e\cdot\mu)h(v))=\lambda g(e\cdot\mu)(gh)(v)=\lambda g(e)g(\mu)(gh)(v).$
It is direct to see that $\phi$ is $L\rtimes G$-linear. Furthermore, $e\cdot V$ has $K$-dimension $\dim_{K}(V)/[L:F]$, so the left-hand side has $K$-dimension $G:G_{e}]\dim_{K}(V)/[L:F]=\dim_{K}(V)$ and it is sufficient to show that $\phi$ is surjective. For this, let $g_{i}$ be a set of left coset representatives of $G_{e}$ in $G$. Because the action of $G$ on $\pi_{0}(\operatorname{Spec}(L))$ is transitive, the primitive idempotents of $L$ are $g_{i}(e)$. Any $v\in V$ can be written as $\sum_{i}g_{i}(e)v$, which is in the image of $\phi$ as $\phi(g_{i}\otimes e\cdot g_{i}^{-1}(v))=g_{i}(e)v$.
Suppose now that $W\in\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})$, and define
$\psi\colon W\rightarrow e\cdot(L\rtimes G\otimes_{L\rtimes G_{e}}W),\qquad\psi(w)=e\otimes w,$
which easily seen to be $F\times G_{e}$-linear. Similarly to above, to show this is an isomorphism it is sufficient to show that $\psi$ is surjective. To see this, note that any element of $L\rtimes G\otimes_{L\rtimes G_{e}}W$ is a sum of elements of the form $\lambda g\otimes w$ for $\lambda\in L$, $g\in G$, $w\in W$, and on these
$e\cdot(\lambda g\otimes w)=e\lambda g\otimes w=gg^{-1}(e)g^{-1}(\lambda)\otimes w=g\otimes eg^{-1}(e)g^{-1}(\lambda)w,$
which is zero whenever $g\not\in G_{e}$, as in this case $eg^{-1}(e)=0$. ∎
For example, this allows us to give a complete description of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ in certain cases.
Definition 7.2.
The extension $L/K$ is called split if there is a $K$-algebra homomorphism $L\rightarrow K$.
This can be characterised in other ways. The following are equivalent:
  • $L/K$ is split,
  • $H=G_{e}$,
  • The inclusion $K\rightarrow F$ is an isomorphism,
  • $L$ is the product of copies of $K$,
  • $\dim_{K}(L)=|\pi_{0}(\operatorname{Spec}(L))|$.
Split extensions are therefore essentially all of the form as described in Example 3.3. For such extensions one obtains a complete description of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, as in this case $\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})=\operatorname{\mathbf{Rep}}_{K}(H)$ is a linear representation category.
Therefore, in summary, together with the results of Section 6 we see that when either the group extension $1\rightarrow H\rightarrow G\rightarrow\Gamma\rightarrow 1$ or the Galois extension $L/K$ is split we have an equivalence from $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ to $\operatorname{\mathbf{Rep}}_{K}(H)$. The following shows that if both are true, then these are essentially the same equivalence.
Proposition 7.3.
Suppose that $L/K$ is split Galois extension with Galois group $\Gamma$, $H$ is a group, and $H\times\Gamma$ acts on $L$ through the projection to $\Gamma$. Then the diagram
$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\Gamma)$$\operatorname{\mathbf{Rep}}_{K}(H)$$\operatorname{\mathbf{Rep}}_{F}(H)$$\scriptstyle{(-)^{\Gamma}}$$\scriptstyle{e\cdot(-)}$$\scriptstyle{F\otimes_{K}-}\scriptstyle{\sim}$
commutes up to natural isomorphism.
Proof.
For $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\Gamma)$, this is defined by
$F\otimes_{K}V^{\Gamma}\rightarrow e\cdot V,\qquad\lambda\otimes v\mapsto\lambda v.$
which is directly seen to be $K[H]$-linear. Both sides have the same $F$-dimension, and therefore this map is an isomorphism as it is surjective: for $v\in e\cdot V$,
$e\otimes\left(\sum_{\gamma\in\Gamma}\gamma(v)\right)\mapsto\sum_{\gamma\in\Gamma}e\gamma(v)=\sum_{\gamma\in\Gamma}e\gamma(e)v=v,$
using that $e\cdot v=v$ and that $e\gamma(e)=0$ for $\gamma\neq 1$. ∎
8. Restriction and Induction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{F}(H)$
In this section we come to the heart of the paper and study restriction and induction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{F}(H)$, which we use to give a description of $\operatorname{Irr}_{L}^{\rtimes}(G)$ in terms of $\operatorname{Irr}_{F}(H)$. We continue with our running notation, where $e$ denote a primitive idempotent of $L$, $F=e\cdot L$, and $\Gamma_{e}=\operatorname{Stab}_{\Gamma}(e)$.
The restriction and induction functors we are interested in are:
$e\cdot(-)$ $\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{F}(H),$
$(L\rtimes G)\otimes_{L[H]}$- $\colon\operatorname{\mathbf{Rep}}_{F}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G).$
Here any $V\in\operatorname{\mathbf{Rep}}_{F}(H)$ is viewed as an $L[H]$-module via the projection $L[H]\rightarrow F[H]$ and then $L\rtimes G$ acts on the left of $(L\rtimes G)\otimes_{L[H]}V$.
8.1. Base Change
The main idea is to interpret these functors in terms of base change.
Definition 8.1.
To ease notation, we set
$R\coloneqq\prod_{\gamma\in\Gamma}F.$
We consider $R$ with an action of of $\Gamma$ by
$\sigma*(\lambda_{\gamma})_{\gamma}\coloneqq(\lambda_{\sigma^{-1}\gamma})_{\gamma}\ \mbox{ for all }\ \sigma\in\Gamma,\ (\lambda_{\gamma})_{\gamma}\in R$
and action of $\Gamma_{e}$ by
$\sigma*(\lambda_{\gamma})_{\gamma}\coloneqq(\sigma(\lambda_{\gamma\sigma}))_{\gamma}\ \mbox{ for all }\ \sigma\in\Gamma_{e},\ (\lambda_{\gamma})_{\gamma}\in R.$
These actions commute.
We also consider $F\otimes_{K}L$ with its natural action of $\Gamma_{e}\times\Gamma$, where $(\gamma_{e},\gamma)$ acts as $\gamma_{e}\otimes\gamma$.
Lemma 8.2.
The map of $F$-algebras
$F\otimes_{K}L\rightarrow R,\qquad\lambda\otimes\mu\mapsto(\lambda\gamma^{-1}(\mu))_{\gamma\in\Gamma},$
is a $\Gamma_{e}\times\Gamma$-equivariant isomorphism.
Proof.
This is an isomorphism, being the composition of the chain of isomorphisms
$F\otimes_{K}L\rightarrow F\otimes_{K}\left(\prod_{\alpha\in\Gamma/\Gamma_{e}}e_{\alpha}\cdot L\right)\rightarrow\prod_{\alpha\in\Gamma/\Gamma_{e}}F\otimes_{K}F\rightarrow\prod_{\alpha\in\Gamma/\Gamma_{e}}\prod_{\beta\in\Gamma_{e}}F\equiv\prod_{\gamma\in\Gamma}F.$
making sure to choose a consistent choice of representatives $\alpha\in\Gamma/\Gamma_{e}$. Here $e_{\alpha}=\alpha(e)$, and we use the isomorphisms $\alpha^{-1}\colon e_{\alpha}\cdot L\rightarrow e\cdot L=F$. It is direct to check that it is $\Gamma_{e}\times\Gamma$-equivariant with respect to the action of $\Gamma_{e}\times\Gamma$ on $R$ described above. ∎
Letting $G$ act on $F\otimes_{K}L$ and $R$ through the action of $\Gamma$, this induces an isomorphism
$F\otimes_{K}(L\rtimes G)\xrightarrow{\sim}(F\otimes_{K}L)\rtimes G\xrightarrow{\sim}R\rtimes G.$
In this setting, the tuple $(G,H,R/F,\sigma\colon G\rightarrow\Gamma)$ satisfies the hypothesis of Section 3.2. A primitive idempotent of $R$ is given by $f\coloneqq(\delta_{\gamma,1})_{\gamma}$. This is a split Galois extension, $f\cdot R=F$, and $H=\operatorname{Stab}_{G}(f)$. In particular, from Proposition 7.1 we have an equivalence of categories
$f\cdot(-)$ $\colon\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{F}(H),$
$(R\rtimes G)\otimes_{R[H]}$- $\colon\operatorname{\mathbf{Rep}}_{F}(H)\rightarrow\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G).$
We also have base change and restriction functors
$F\otimes_{K}$- $\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{F\otimes_{K}L}^{\rtimes}(G),$
[…]
Our next results allow us to interpret our restriction functor $e\cdot(-)\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{F}(H)$ as this base change functor $F\otimes_{K}$-, and our induction functor $(L\rtimes G)\otimes_{F[H]}-\colon\operatorname{\mathbf{Rep}}_{F}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ as this restriction functor.
Proposition 8.3.
The diagram
[…]
commutes up to natural transformation.
Proof.
Suppose that $W\in\operatorname{\mathbf{Rep}}_{F}(H)$. The induced $L\rtimes G$-module is defined by first viewing $W$ as a $L[H]$-module via the projection $L[H]\rightarrow F[H]$, and then letting $L\rtimes G$ act on the left of $(L\rtimes G)\otimes_{L[H]}W$. On the other hand, we obtain a $R\rtimes G$-module by letting $R[H]$ act on $W$ through the projection $R[H]\rightarrow F[H]$ and letting $R\rtimes G$ act on the left of $(R\rtimes G)\otimes_{R[H]}W$. There is a natural map
$(L\rtimes G)\otimes_{L[H]}W\rightarrow(R\rtimes G)\otimes_{R[H]}W$
induced by the inclusions $L[H]\hookrightarrow R[H]$ and $L\rtimes G\hookrightarrow R\rtimes G$, which is $L\rtimes G$-linear. To see that it is an isomorphism, first note that it is surjective, as any pure tensor of $rg\otimes w\in(R\rtimes G)\otimes_{R[H]}W$ can be expressed as
$rg\otimes w=gg^{-1}(r)\otimes w=g\otimes g^{-1}(r)w,$
and thus lies in the image. There it is an isomorphism, being $K$-linear and both modules having $K$-dimension $|\Gamma_{e}|\dim_{K}W$. ∎
Proposition 8.4.
The diagram
$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{F\otimes_{K}L}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{F}(H)$$\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)$$\scriptstyle{F\otimes_{K}-}$$\scriptstyle{e\cdot(-)}\scriptstyle{\sim}$$\scriptstyle{f\cdot(-)}\scriptstyle{\sim}$
commutes up to natural transformation.
Proof.
Suppose that $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. There is a well-defined action of $(F\otimes_{K}L)\rtimes G$ on $(F\otimes_{K}L)\otimes_{L}V$ defined by $xg\cdot(y\otimes v)\coloneqq xg(y)\otimes g(v)$, for which is natural identification
$F\otimes_{K}V\xrightarrow{\sim}(F\otimes_{K}L)\otimes_{L}V$
is $F\otimes_{K}(L\rtimes G)$-equivariant, where $F\otimes_{K}(L\rtimes G)$ acts on $(F\otimes_{K}L)\otimes_{L}V$ via the isomorphism $F\otimes_{K}(L\rtimes G)\xrightarrow{\sim}(F\otimes_{K}L)\rtimes G$. Then under the above functor diagram, $V$ maps to the $R\rtimes G$-module $R\otimes_{L}V$, where similarly $R\rtimes G$ acts by $xg\cdot(y\otimes v)\coloneqq xg(y)\otimes g(v)$. Then
$f\cdot(R\otimes_{L}V)=(f\cdot R)\otimes_{L}V=F\otimes_{L}V\xrightarrow{\sim}e\cdot V,$
as $F[H]$-modules, where the last isomorphism is defined by $\lambda\otimes v\mapsto\lambda v$. The map $L\rightarrow R$ is defined by $\mu\mapsto(e\gamma^{-1}(\mu))_{\gamma}$, and so the map to $L\rightarrow f\cdot R=F$ is $\mu\mapsto e\cdot\mu$, so this map is well-defined, and an isomorphism, having inverse $v\mapsto e\otimes v$. ∎
8.2. Consequences
This allows us to transport results about base change functors to our setting.
Corollary 8.5.
For any $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$,
$(L\rtimes G)\otimes_{L[H]}e\cdot V\cong V^{\oplus|\Gamma_{e}|}.$
Proof.
This follows from the commutativity of the diagrams of Propositions 8.3 and 8.4, and the fact that for a $K$-algebra AA, and $V\in\operatorname{\mathbf{Mod}}_{A}$, $F\otimes_{K}V\cong V^{\oplus|\Gamma_{e}|}$ as AA-modules. ∎
Corollary 8.6.
$V\cong W$ in $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ if and only if $e\cdot V\cong e\cdot W$ in $\operatorname{\mathbf{Rep}}_{F}(H)$.
Proof.
This follows directly from Corollary 8.5 and Lemma 2.14. ∎
Corollary 8.7.
For $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, the natural map
$F\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\rightarrow\operatorname{Hom}_{F[H]}(e\cdot V,e\cdot W)$
is an isomorphism.
Remark 8.8.
This isomorphism is compatible with composition of morphisms in the obvious sense. In particular, when $V=W$ this map is a ring isomorphism
$F\otimes_{K}\operatorname{End}_{L\rtimes G}(V)\xrightarrow{\sim}\operatorname{End}_{F[H]}(e\cdot V).$
Proof.
We have that
$F\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\rightarrow\operatorname{Hom}_{F[H]}(e\cdot V,e\cdot W),$
corresponds under the $K$-linear equivalences of Proposition 8.4 to the natural map
$F\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\xrightarrow{\sim}\operatorname{Hom}_{(L\rtimes G)_{F}}(F\otimes_{K}V,F\otimes_{K}W),$
which is an isomorphism by [CR, Lem. 29.5]. ∎
Remark 8.9.
Using this we can also show that the natural map
$L\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\rightarrow\operatorname{Hom}_{L[H]}(V,W),\qquad\lambda\otimes f\mapsto\lambda f$
of Lemma 2.12 is also an isomorphism, as this factors as
$L\otimes_{K}\operatorname{Hom}_{L\rtimes G}(V,W)$ $\xrightarrow{\sim}L\otimes_{F}\operatorname{Hom}_{F[H]}(e\cdot V,e\cdot W),$
$\xrightarrow{\sim}\operatorname{Hom}_{L[H]}(L\otimes_{F}e\cdot V,L\otimes_{F}e\cdot W),$
$\xrightarrow{\sim}\operatorname{Hom}_{L[H]}(V,W),$
using the $L[H]$-module isomorphisms $L\otimes_{F}e\cdot V\xrightarrow{\sim}V$, $L\otimes_{F}e\cdot W\xrightarrow{\sim}W$.
Example 8.10.
Suppose that $L$ is connected, so $L=F$. Then by Lemma 6.2 we see that Corollary 8.7 generalises the fact that for any group $H$ the map
$L\otimes_{K}\operatorname{Hom}_{K[H]}(V,W)\rightarrow\operatorname{Hom}_{L[H]}(L\otimes_{K}V,L\otimes_{K}W).$
is an isomorphism for $V,W\in\operatorname{\mathbf{Rep}}_{K}(H)$.
Remark 8.11.
Corollary 8.5 and Corollary 8.7 generalise [RUMTAY, Prop. 2.9] and [RUMTAY, Thm. 5.1] respectively, which consider the special case where $L/K={\mathbb{C}}/{\mathbb{R}}$ and $G$ is finite.
Corollary 8.7 also has implications for the base change functor $L\otimes_{K}$- considered in Section 4.1.
Corollary 8.12.
For $V,W\in\operatorname{\mathbf{Rep}}_{K}(G)$, we have an equality
$\dim_{K}\operatorname{Hom}_{L\rtimes G}(L\otimes_{K}V,L\otimes_{K}W)=\dim_{K}\operatorname{Hom}_{K[H]}(V,W).$
Remark 8.13.
We stress that this is only an equality of dimensions - there is only a canonical map between these spaces after we tensor with $L$ over $K$, and not before.
Proof.
From Remark 8.9 and Example 8.10 we have isomorphisms
$L\otimes_{K}\operatorname{Hom}_{L\rtimes G}(L\otimes_{K}V,L\otimes_{K}W)\xrightarrow{\sim}\operatorname{Hom}_{L[H]}(L\otimes_{K}V,L\otimes_{K}W)\xleftarrow{\sim}L\otimes_{K}\operatorname{Hom}_{K[H]}(V,W).\qed$
8.3. Preservation of Semisimplicity
Suppose now that $A$ is a $K$-algebra, and $M$ is a simple $A$-module which is finite dimensional over $K$. Then because $M$ is finite dimensional, the $K$-algebra homomorphism $A\rightarrow\operatorname{End}_{K}(M)$ factors as
$A\twoheadrightarrow B\hookrightarrow\operatorname{End}_{K}(M),$
where $B$ is the image of $A$, a finite dimensional $K$-algebra. Note that because $M$ is a simple $B$-module, $B$ is a simple $K$-algebra by [CR1, Prop. 3.31]. Furthermore, because $F/K$ is separable, $F\otimes_{K}M$ is a semisimple $F\otimes_{K}B$-module by [CR2, Cor. 7.8(ii)] and hence a semisimple $F\otimes_{K}A$-module.
Conversely, if $N$ is a simple $F\otimes_{K}A$-module, which is finite dimensional over $K$, then writing $B$ for the image of $A$ in $\operatorname{End}_{K}(N)$, $B$ is a finite dimensional $K$-algebra with $J(B)$ acting trivially on $N$ (because $J(F\otimes_{K}B)=F\otimes_{K}J(B)$ [CR1, Thm. 7.9] and $N$ is a simple $F\otimes_{K}B$-module), hence $N$ is a semisimple $B$-module, and thus a semisimple $A$-module.
Corollary 8.14.
The functors
$e\cdot(-)$ $\colon\operatorname{\mathbf{Rep}}_{L}(G)\rightarrow\operatorname{\mathbf{Rep}}_{F}(H),$
$(L\rtimes G)\otimes_{L[H]}$- $\colon\operatorname{\mathbf{Rep}}_{F}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}(G)$
preserve semisimplicity.
Proof.
Taking $A=L\rtimes G$, this follows from the above discussion and the commutativity of the diagrams of Propositions 8.3 and 8.4, the equivalences of which preserve semisimplicity. ∎
8.4. Action of $\Gamma_{e}$
In this section we consider the two natural actions of the group $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{F}(H)$, and show that these agree under the equivalences Propositions 8.3 and 8.4.
8.4.1. Action of $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)$
For any $\sigma\in\Gamma_{e}$, we have a $K$-algebra isomorphism
$\sigma\otimes 1\colon F\otimes_{K}(L\rtimes G)\rightarrow F\otimes_{K}(L\rtimes G).$
Under the isomorphism $F\otimes_{K}(L\rtimes G)\xrightarrow{\sim}R\rtimes G$ induced from Lemma 8.2, this corresponds to
$\sigma\colon R\rtimes G\rightarrow R\rtimes G,\qquad(\lambda_{\gamma})_{\gamma}\cdot g\mapsto(\sigma(\lambda_{\gamma\sigma}))_{\gamma}\cdot g.$
In particular, we obtain a left action on $\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)/\!\cong$, the set of isomorphism classes of objects:
Definition 8.15.
$M\in\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)$, $^{\sigma}M$ is $M$ but with $R\rtimes G$ action through $\sigma^{-1}\colon R\rtimes G\rightarrow R\rtimes G$.
8.4.2. Action of $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{F}(H)$
Recall $\Gamma_{e}=\operatorname{Stab}_{\Gamma}(e)$ and $G_{e}=\sigma^{-1}(\Gamma_{e})$, where $\sigma$ denotes the surjection $\sigma\colon G\rightarrow\operatorname{Gal}(L/K)$. We have a left action of $G_{e}$ on $\operatorname{\mathbf{Rep}}_{F}(H)/\!\cong$ defined as follows.
Definition 8.16.
For $g\in G_{e}$ and $W\in\operatorname{\mathbf{Rep}}_{F}(H)$, $g*W\in\operatorname{\mathbf{Rep}}_{F}(H)$ is $W$ as an abelian group, with
  • $h*w=(g^{-1}hg)(w)$,
  • $\lambda*w=\sigma_{g^{-1}}(\lambda)w$,
for $\lambda\in F$, $h\in H$ and $w\in W$.
Note that the induced action on isomorphism classes of objects is trivial when restricted to $H$, and thus becomes an action of the finite group $G_{e}/H$ and therefore of $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{F}(H)/\!\cong$ using the isomorphism $G_{e}/H\xrightarrow{\sim}\Gamma_{e}$.
Using this action, we can describe the other composition of $L\rtimes G\otimes_{F[H]}$- and $e\cdot(-)$. We could do this using Propositions 8.3 and 8.4, but it is easier to see directly.
Proposition 8.17.
For any $W\in\operatorname{\mathbf{Rep}}_{F}(H)$ there is an isomorphism
$e\cdot((L\rtimes G)\otimes_{L[H]}W)\xrightarrow{\sim}\bigoplus_{g\in G_{e}/H}\>g*W.$
Proof.
Pick coset representatives $g_{i}$ for $G/H$. First, $L\rtimes G=\oplus_{i}\>g_{i}L[H]$ as an $L$-module, hence
$(L\rtimes G)\otimes_{L[H]}W=\bigoplus_{i}g_{i}L[H]\otimes_{L[H]}W.$
If $g_{i}\not\in G_{e}$, then $\sigma_{g_{i}^{-1}}(e)$ is sent to zero under $L\rightarrow F$, and so $e\cdot(g_{i}L[H]\otimes_{L[H]}W)=0$ as
$e\cdot g_{i}x\otimes w=g_{i}\sigma_{g_{i}^{-1}}(e)x\otimes w=g_{i}x\otimes 0=0.$
Therefore
$e\cdot((L\rtimes G)\otimes_{L[H]}W)=\bigoplus_{g_{i}\in G_{e}}g_{i}L[H]\otimes_{L[H]}W.$
Each summand $g_{i}L[H]\otimes_{L[H]}W$ is stable under the action of $F[H]$, and for $h\in H$, $\lambda\in F$ and $w\in W$,
$\lambda(g_{i}\otimes w)$ $=\lambda g_{i}\otimes w=g_{i}\sigma_{g_{i}^{-1}}(\lambda)\otimes w=g_{i}\otimes\sigma_{g_{i}^{-1}}(\lambda)w,$
$h(g_{i}\otimes w)$ $=hg_{i}\otimes w=g_{i}(g_{i}^{-1}hg_{i})\otimes w=g_{i}\otimes(g_{i}^{-1}hg_{i})w,$
which is exactly the action of $F[H]$ on $g_{i}*W$. ∎
The next lemma shows that the natural action of of $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)/\cong$ agrees under the equivalences of Propositions 8.3 and 8.4 with the action of $\Gamma_{e}$ on $\operatorname{\mathbf{Rep}}_{F}(H)/\cong$ described above. We will use this later in Proposition 8.29, to describe the endomorphism ring of irreducible objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$.
Lemma 8.18.
For any $g\in G_{e}$ and $W\in\operatorname{\mathbf{Rep}}_{F}(H)$,
$\begin{matrix}\phi:&{}^{\sigma_{g}}((R\rtimes G)\otimes_{R[H]}W)&\rightarrow&(R\rtimes G)\otimes_{R[H]}g*W\\ &x\otimes w&\mapsto&\sigma_{g}(x)\cdot g^{-1}\otimes w\end{matrix}$
is a natural isomorphism in $\operatorname{\mathbf{Rep}}_{R}^{\rtimes}(G)$.
Proof.
To see that this induces a well-defined map on the tensor product, let $\mu h\in R[H]$, which acts on $W$ by $w\mapsto\mu_{1}h(w)$. Then
$x\otimes\mu_{1}h(w)\mapsto\sigma_{g}(x)\cdot g^{-1}\otimes\mu_{1}h(w).$
On the other hand,
$x\cdot\mu h\otimes w$ $\mapsto\sigma_{g}(x)\cdot\sigma_{g}(\mu)h\cdot g^{-1}\otimes w,$
$=\sigma_{g}(x)\cdot\sigma_{g}(\mu)g^{-1}\cdot ghg^{-1}\otimes w,$
$=\sigma_{g}(x)\cdot g^{-1}\cdot g(\sigma_{g}(\mu))\cdot ghg^{-1}\otimes w,$
$=\sigma_{g}(x)\cdot g^{-1}\cdot g(\sigma_{g}(\mu_{\gamma\sigma_{g}}))_{\gamma}\cdot ghg^{-1}\otimes w,$
$=\sigma_{g}(x)\cdot g^{-1}\cdot(\sigma_{g}(\mu_{\sigma_{g}^{-1}\gamma\sigma_{g}}))_{\gamma}\cdot ghg^{-1}\otimes w,$
$=\sigma_{g}(x)\cdot g^{-1}\otimes\sigma_{g}(\mu_{1})(ghg^{-1})*w,$
$=\sigma_{g}(x)\cdot g^{-1}\otimes\mu_{1}h(w).$
Where here * denote the action of $F[H]$ on gWg*W. For $y\in R$,
$y\cdot\phi(x\otimes w)=y\cdot\sigma_{g}(x)\cdot g^{-1}\otimes w,$
whereas
$\phi(y*(x\otimes w))=\phi(\sigma_{g^{-1}}(y)\cdot x\otimes w)=y\cdot\sigma_{g}(x)\cdot g^{-1}\otimes w$
and thus the map is $R$-linear. It also directly seen to be an isomorphism, natural in $W$. ∎
8.5. Description of Indecomposables
We can use Corollary 8.14 to obtain a precise description of the restriction $e\cdot V$ for indecomposable $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Recall that $\operatorname{Ind}_{L}^{\rtimes}(G)$ and $\operatorname{Ind}_{F}(H)$ denote the sets of isomorphism classes of indecomposable objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{F}(H)$ respectively.
Definition 8.19.
For $W\in\operatorname{Ind}_{F}(H)$, we write $\Gamma_{e,W}\coloneqq\operatorname{Stab}_{\Gamma_{e}}(W)$.
Theorem 8.20.
Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$, and $W$ is any indecomposable direct summand of $e\cdot V$. Then
$e\cdot V=\bigoplus_{\gamma\in\Gamma_{e}/\Gamma_{e,W}}(\gamma*W)^{m(V)}$
for some $m(V)\geq 1$. Furthermore, if $W\in\operatorname{Ind}_{F}(H)$ then
  • $W$ appears in $e\cdot V$ for a unique $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$,
  • This $V$ is irreducible if and only if $W$ is irreducible,
  • For this $V$, $(L\rtimes G)\otimes_{L[H]}(g*W)=V^{\oplus r}$ for any $g\in G_{e}$, where $r=|\Gamma_{e,W}|/m(V)$.
In particular, $m(V)\mid|\Gamma_{e,W}|$, and sending $W$ to this unique $V$ defines bijective correspondences
$\operatorname{Ind}_{L}^{\rtimes}(G)\xleftrightarrow{\sim}\operatorname{Ind}_{F}(H)/\Gamma_{e},\qquad\operatorname{Irr}_{L}^{\rtimes}(G)\xleftrightarrow{\sim}\operatorname{Irr}_{F}(H)/\Gamma_{e}$
where conversely $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ is sent to any indecomposable direct summand of $e\cdot V$.
Proof.
For the first claim, write $e\cdot V$ as a direct sum of indecomposable objects of $\operatorname{\mathbf{Rep}}_{L}(H)$ using Lemma 2.14, and let $W$ be an indecomposable direct summand of $e\cdot V$. For any $g\in G_{e}$, the number of copies of each gWg*W in $e\cdot V$ must be equal to the number of copies of $W$, using Lemma 2.14 and the fact that the action $\rho(g)$ of $g$ on $V$ defines an isomorphism $\rho(g)\colon g*(e\cdot V)\xrightarrow{\sim}e\cdot V$.
It therefore remains remains to see that only one orbit of the action of $\Gamma_{e}$ on $\operatorname{Ind}_{F}(H)$ appears in $e\cdot V$. Indeed, suppose that $W_{1},W_{2}\in\operatorname{Ind}_{L}(H)$ both appear in $e\cdot V$, so $e\cdot V=W_{1}\oplus W_{2}\oplus U$ for some $U\in\operatorname{\mathbf{Rep}}_{L}(H)$. Then $(L\rtimes G)\otimes_{L[H]}(e\cdot V)\cong V^{\oplus|\Gamma_{e}|}$ by Corollary 8.5, and so by Lemma 2.14 both $(L\rtimes G)\otimes_{L[H]}W_{1}$ and $(L\rtimes G)\otimes_{L[H]}W_{2}$ are isomorphic to a direct sum of a number of copies of $V$. In particular, there are $a,b\geq 1$ with
$\left((L\rtimes G)\otimes_{L[H]}W_{1}\right)^{\oplus a}\cong\left((L\rtimes G)\otimes_{L[H]}W_{2}\right)^{\oplus b}$
and applying $e\cdot V$ we see that
$\bigoplus_{x\in\Gamma_{e}}(x\cdot W_{1})^{\oplus a}\cong\bigoplus_{y\in\Gamma_{e}}(y\cdot W_{2})^{\oplus b}$
by Proposition 8.17 and therefore $W_{2}\cong\gamma\cdot W_{1}$ for some $\gamma\in\Gamma_{e}$, by Lemma 2.14.
For the remaining claims, suppose that $W\in\operatorname{Ind}_{F}(H)$. From Proposition 8.17, $W$ is a direct summand of $e\cdot((L\rtimes G)\otimes_{L[H]}W)$, and thus of $e\cdot V$ for some $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ after writing $(L\rtimes G)\otimes_{L[H]}W$ as a direct sum of indecomposable objects and using Lemma 2.14. For this $V$, from the above we have that
$V^{|\Gamma_{e}|}\cong(L\rtimes G)\otimes_{L[H]}(e\cdot V)\cong\bigoplus_{\gamma\in\Gamma_{e}/\Gamma_{e,W}}((L\rtimes G)\otimes_{L[H]}(\gamma*W))^{m(V)}$
and therefore $(L\rtimes G)\otimes_{L[H]}(\gamma*W)=V^{\oplus r}$ for any $\gamma\in\Gamma_{e}$ by Lemma 2.14, where $r=|\Gamma_{e,W}|/m(V)$.
In particular, the $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ for which $W$ is a direct summand of $e\cdot V$ is unique. Finally, from Corollary 8.14 we have that $V$ is irreducible if and only if $W$ is irreducible. ∎
We can use this to characterise when $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple, improving on Proposition 2.11. It also possible to show this directly, using a generalisation of Maschke’s Theorem [KUN, Lem. 1.3].
Corollary 8.21.
Suppose that $G$ is finite. Then $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple if and only if $|H|\in L^{\times}$.
Proof.
By Lemma 2.14, the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple if and only if $\operatorname{Irr}_{L}^{\rtimes}(G)=\operatorname{Ind}_{L}^{\rtimes}(G)$, which by Theorem 8.20 is equivalent to the statement that $\operatorname{Irr}_{F}(H)=\operatorname{Ind}_{F}(H)$, or in other words that $\operatorname{\mathbf{Rep}}_{F}(H)$ is semisimple. By Maschke’s Theorem this is equivalent to $|H|$ being invertible in $F$, which is the same as $|H|$ being invertible in $L$, as $L$ is the product of copies of $F$. ∎
In general, we will see that any positive integer can occur as some $m(V)$ (Section 16). One case when we know that the numbers $m(V)$ are equal to $1$ is the following. This applies in particular when $K$ is a finite field, by Wedderburn’s Little Theorem [milneCFT, Thm. IV.4.1].
Proposition 8.22.
Suppose that $K$ is a field such that $\operatorname{Br}(K^{\prime})=0$ for any finite extension $K^{\prime}/K$, or equivalently, such that any finite dimensional division algebra over $K$ is a field. Then for any $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, $m(V)=1$.
Proof.
By Corollary 8.7, there is a $K$-algebra isomorphism
$L\otimes_{K}\operatorname{End}_{L\rtimes G}(V)\xrightarrow{\sim}\operatorname{End}_{F[H]}(e\cdot V),$
and thus $\operatorname{End}_{F[H]}(e\cdot V)$ is commutative because $\operatorname{End}_{L\rtimes G}(V)$ is a field. In particular, $e\cdot V$ cannot contain any factors with multiplicity greater than $1$, so $m(V)=1$. ∎
We also record here that the induction and restriction functors of Theorem 8.20 are compatible with the equivalence of Proposition 7.1, the proof of which is direct. We use the alternative notation for the functors between $\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})$ and $\operatorname{\mathbf{Rep}}_{F}(H)$, described at the start of Section 8.6 below.
Proposition 8.23.
The diagrams
$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})$$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{F}^{\rtimes}(G_{e})$$\operatorname{\mathbf{Rep}}_{F}(H)$$\operatorname{\mathbf{Rep}}_{F}(H)$$\scriptstyle{e\cdot(-)}\scriptstyle{\sim}$$\scriptstyle{e\cdot(-)}$$\scriptstyle{(-)|_{H}}$$\scriptstyle{(L\rtimes G)\otimes_{L\rtimes G_{e}}-}\scriptstyle{\sim}$$\scriptstyle{(L\rtimes G)\otimes_{L[H]}-}$$\scriptstyle{\operatorname{Ind}_{H}^{G}(-)}$
both commute up to natural isomorphism.
8.6. The Case When $L$ is a Field
From now on in this paper on we assume that $L$ is a field.
Note that one is always reduced to this case in practice by Proposition 7.1 and Proposition 8.23. As $L$ is a field, the action map $\Gamma\rightarrow\operatorname{Gal}(L/K)$ is an isomorphism (cf. Example 3.2), and so without loss of generality one could assume that $\Gamma=\operatorname{Gal}(L/K)$. Furthermore, $G_{e}=G$, $\Gamma_{e}=\Gamma$, $L=F$, and the functors of the previous section become
$(-)|_{H}$ $\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{L}(H),$
$\operatorname{Ind}_{H}^{G}(-)$ $\colon\operatorname{\mathbf{Rep}}_{L}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G),$
where $\operatorname{Ind}_{H}^{G}$ is as defined as described at the end of the introduction.
We first give a strengthening of Theorem 8.20 to intermediate extensions, which we will use in Section 9 when we consider the semilinear Schur index.
Let $S$ be any intermediate subgroup $H\leq S\leq G$, which corresponds to a unique subgroup $\Gamma_{S}\leq\Gamma$ with $S=\sigma^{-1}(\Gamma_{S})$. If $F$ is the fixed field of $\Gamma_{S}$, then $K\leq F\leq L$ and $L/F$ is Galois. In particular, the tuple $(S,H,L/F,\sigma\colon S\rightarrow\Gamma_{S})$ satisfies the assumptions of Section 3.2, and so by Theorem 8.20 there are canonical bijections
$\begin{array}[]{cc}\begin{aligned} \operatorname{Ind}_{L}^{\rtimes}(G)&\xleftrightarrow{\sim}\operatorname{Ind}_{L}(H)/\Gamma,\\ \operatorname{Ind}_{L}^{\rtimes}(S)&\xleftrightarrow{\sim}\operatorname{Ind}_{L}(H)/\Gamma_{S}.\end{aligned}&\begin{aligned} \operatorname{Irr}_{L}^{\rtimes}(G)&\xleftrightarrow{\sim}\operatorname{Irr}_{L}(H)/\Gamma,\\ \operatorname{Irr}_{L}^{\rtimes}(S)&\xleftrightarrow{\sim}\operatorname{Irr}_{L}(H)/\Gamma_{S}.\end{aligned}\end{array}$
Theorem 8.20 further provides a description of induction and restriction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$, and induction and restriction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$. We now wish to understand induction and restriction between $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S)$.
Proposition 8.24.
Let $\mathcal{A}$} be an orbit of $G$ on $\operatorname{Ind}_{L}(H)$, corresponding to $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$, and let
${\mathcal{A}}_{1}\sqcup\cdots\sqcup{\mathcal{A}}_{r}$
be a decomposition of $\mathcal{A}$} into $S$-orbits. Let $U_{i}\in\operatorname{Ind}_{L}^{\rtimes}(S)$ correspond to ${\mathcal{A}}_{i}$, so
$U_{i}|_{H}\cong\bigoplus_{X\in{\mathcal{A}}_{i}}X^{\oplus m(U_{i})}$
for some $m(U_{i})\geq 1$ (by Theorem 8.20). Then $m(U_{i})\mid m(V)$, $m(V)\mid m(U_{i})\cdot[\Gamma_{X}:\Gamma_{S,X}]$,
$V|_{S}\cong\bigoplus_{i=1}^{r}U_{i}^{\oplus\frac{m(V)}{m(U_{i})}},$
and
$\operatorname{Ind}_{S}^{G}(U_{i})\cong V^{\oplus\frac{[\Gamma_{X}:\Gamma_{S,X}]\cdot m(U_{i})}{m(V)}}$
for any $X\in{\mathcal{A}}_{i}$. Furthermore, $V$ is irreducible if and only if each $U_{i}$ is irreducible.
Proof.
For the first claim, note that using Corollary 8.6 for $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S)$,
$(V|_{S})^{m(U_{1})\cdots m(U_{r})}\cong\bigoplus_{i=1}^{r}U_{i}^{\oplus m(V)m(U_{1})\cdots m(U_{i-1})m(U_{i+1})\cdots m(U_{r})}$
because their restrictions to $H$ are isomorphic. Then the description of $V|_{S}$ follows from Lemma 2.14.
For the second claim, take any $X\in{\mathcal{A}}_{i}$, and note that by Theorem 8.20
$V^{\oplus\frac{|\Gamma_{X}|}{m(V)}}\cong\operatorname{Ind}_{H}^{G}(X)\cong\operatorname{Ind}_{S}^{G}(\operatorname{Ind}_{H}^{S}(X))\cong\operatorname{Ind}_{S}^{G}(U_{i})^{\oplus\frac{|\Gamma_{S,X}|}{m(U_{i})}}.$
The description of $\operatorname{Ind}_{S}^{G}(U_{i})$ then simiarly follows directly from Lemma 2.14.
For the final claim, note that $V$ is irreducible if and only if each element of $\mathcal{A}$} is irreducible, and $U_{i}$ is irreducible if and only if each element of ${\mathcal{A}}_{i}$ is irreducible, in each case by Theorem 8.20. ∎
Example 8.25.
If $S=G$, then $r=1$, $m(V)=m(U_{1})$, and $\Gamma_{X}:\Gamma_{S,X}]=1$. If $S=H$, then $r=[\Gamma:\Gamma_{X}]$, $m(U_{i})=1$, and $\Gamma_{X}:\Gamma_{S,X}]=[\Gamma_{X}:1]$.
The main case of interest is the following. To state it, note that if $S$ is a normal subgroup of $G$, then there is an action of $G/S$ on $\operatorname{Ind}_{L}^{\rtimes}(S)$, where for $g\in G$, and $U\in\operatorname{Ind}_{L}^{\rtimes}(S)$, gUg*U is the abelian group $U$ but with $L\rtimes S$-module structure
  • $s*u=(g^{-1}sg)(u)$,
  • $\lambda*u=\sigma_{g}^{-1}(\lambda)u$,
for $\lambda\in L$, $s\in S$ and $u\in U$. This action preserves the subset $\operatorname{Irr}_{L}^{\rtimes}(S)$ of $\operatorname{Ind}_{L}^{\rtimes}(S)$.
Remark 8.26.
The action of $G/H$ on $\operatorname{Ind}_{L}(H)$ is the restriction to the diagonal of an action of the product $G/H\times G/H$ on $\operatorname{Ind}_{L}(H)$ (where one copy acts by conjugation in $G$ and the other acts on scalars - see Section 14). When $S>H$, however, this action of $G/S$ on $\operatorname{Ind}_{L}^{\rtimes}(S)$ is not the restriction of an action of the product $G/S\times G/S$ in the same way: one only has an action of the diagonal.
Corollary 8.27.
Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$, and that for some corresponding $W\in\operatorname{Ind}_{L}(H)$, $G_{W}$ is normal in $G$. Set $G_{V}\coloneqq G_{W}$, which is independent of the choice of $W$. Let $U\in\operatorname{Ind}_{L}^{\rtimes}(G_{V})$ correspond to $W$. Then
$V|_{G_{W}}$ $\cong\bigoplus_{g\in G/G_{V}}g*U,$
$V$ $\cong\operatorname{Ind}_{G_{V}}^{G}(g*U),$
$V$ is irreducible if and only if $U$ is irreducible, and for any $g\in G$, gUg*U corresponds to gWg*W, so
$(g*U)|_{H}$ $\cong(g*W)^{\oplus m(V)},$
$\operatorname{Ind}_{H}^{G_{V}}(g*W)$ $\cong(g*U)^{\oplus\frac{[G_{V}:H]}{m(V)}}.$
Proof.
We have that for any $g\in G$, $G_{g*W}=gG_{W}g^{-1}=G_{W}$. In particular, the $G$-orbit $\{g*W\}_{g\in G/G_{V}}$ decomposes into singleton $G_{V}$-orbits, which give the $\{U_{1},...,U_{r}\}$ from Proposition 8.24. We have that m(gU)|m(V)m(g*U)|m(V) and $m(V)\mid m(g*U)\cdot[G_{V}:G_{V}]=m(g*U)$, and therefore m(V)=m(gV)m(V)=m(g*V). Then the statements follow from Theorem 8.20 and Proposition 8.24. ∎
Now we are interested in the condition that $L$ is furthermore a splitting field for $W\in\operatorname{Ind}_{L}(H)$ (meaning that $\operatorname{End}_{L[H]}(W)=L$, or equivalently that $W_{L^{\prime}}$ is irreducible for any field extension $L^{\prime}/L$).
Proposition 8.28.
Suppose that $W\in\operatorname{Ind}_{L}(H)$, corresponding to $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$. Then the $K$-algebra $\operatorname{End}_{L\rtimes G}(V)$ is central over $K$ if and only if $\Gamma$ fixes $W$ and $Z(\operatorname{End}_{L[H]}(W))=L$.
Proof.
From the isomorphism
$L\otimes_{K}\operatorname{End}_{L\rtimes G}(V)\rightarrow\operatorname{End}_{L[H]}(V|_{H})$
of Corollary 8.7, $Z(\operatorname{End}_{L\rtimes G}(V))=K$ if and only if $Z(\operatorname{End}_{L[H]}(V|_{H}))=L$, by [milneCFT, Prop. IV.2.3]. From Theorem 8.20, this occurs if and only if $\Gamma$ fixes $W$ and $Z(\operatorname{End}_{L[H]}(W))=L$. ∎
Proposition 8.29.
Suppose that $W\in\operatorname{Irr}_{L}(H)$, corresponding to $U\in\operatorname{Irr}_{L}^{\rtimes}(G_{W})$ and $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and that $\operatorname{End}_{L[H]}(W)=L$. Set $L_{W}\coloneqq L^{G_{W}}$, and $D\coloneqq\operatorname{End}_{L\rtimes G}(V)$. Then
  1. (1)
    $Z(D)\cong L_{W}$,
  2. (2)
    $\dim_{Z(D)}D=m(V)^{2}$,
  3. (3)
    $\operatorname{End}_{L\rtimes G_{W}}(U)$ is central over $L_{W}$,
  4. (4)
    $\dim_{L_{W}}(\operatorname{End}_{L\rtimes G_{W}}(U))=m(V)^{2}$.
Proof.
For the first point, let $A\coloneqq(L\rtimes G)/\text{Ann}_{L\rtimes G}(V)$, which, as $V$ is finite-dimensional over $L$, is a finite dimensional $K$-algebra. Note that because $V$ is a simple $A$-module, AA is furthermore simple by [CR1, Prop. 3.31]. As $A$ and $A_{L}$ are both quotients, they define full-subcategories
[…]
which are closed under sub-objects, and the base change and restriction functors restrict to the base change and restriction functors between $A$ and $A_{L}$. Let $M\in\operatorname{\mathbf{Mod}}^{\operatorname{fd}}_{A_{L}}$ correspond to $W$. The $K$-algebra $\operatorname{End}_{A_{L}}(M)$ contains $L$, and has the same $K$-dimension as $\operatorname{End}_{L[H]}(W)=L$, and therefore $\operatorname{End}_{A_{L}}(M)=L$ and $A_{L}$ is split over $L$. We may therefore apply [CR2, Thm. 74.4 (ii)] to obtain that $Z(\operatorname{End}_{L\rtimes G}(V))\cong L^{\Gamma_{M}}$, where $\Gamma_{M}$ is the stabiliser of $M$ in $\Gamma$. The action of $\Gamma$ on $\operatorname{Irr}_{L}(H)$ corresponds to that of $\Gamma$ on $\operatorname{Irr}(A_{L})$ by Lemma 8.18, and therefore by Proposition 8.4 $\Gamma_{W}=\Gamma_{M}$ and
$Z(\operatorname{End}_{L\rtimes G}(V))\cong L^{\Gamma_{M}}=L^{\Gamma_{W}}=L_{W}.$
For the second point, letting $D\coloneqq\operatorname{End}_{L\rtimes G}(V)$, then applying Corollary 8.7,
$\dim_{Z(D)}(D)$ $=[Z(D):K]\dim_{K}(D),$
$=\dim_{L}\operatorname{End}_{L[H]}(V|_{H}),$
$=\dim_{L}\left(\prod_{i=1}^{[G:G_{W}]}M_{m(V)}(L)\right),$
$=[G:G_{W}]\cdot m(V)^{2},$
and hence point (2) follows from point (1), as $G:G_{W}]=[L_{W}:K]$. Similarly, by Corollary 8.7, there is an isomorphism
$L\otimes_{L_{W}}\operatorname{End}_{L\rtimes G_{W}}(U)\xrightarrow{\sim}\operatorname{End}_{L[H]}(U|_{H})\cong M_{m(U)}(L),$
because the $G_{W}$ orbit of $W$ is a singleton. Then taking the centre, by [milneCFT, Prop. IV.2.3], we have that
$L\otimes_{L_{W}}Z(\operatorname{End}_{L\rtimes G_{W}}(U))\xrightarrow{\sim}Z(\operatorname{End}_{L[H]}(U|_{H})\cong M_{m(U)}(L))=L,$
hence $\dim_{L_{W}}Z(\operatorname{End}_{L\rtimes G_{W}}(U))=1$. The final claim again follows Corollary 8.7, using the fact that
$U|_{H}\cong W^{\oplus m(V)}$
by Proposition 8.24, as $m(U)\mid m(V)$ and $m(V)\mid m(U)\cdot[G_{W}:G_{W}]=m(U)$. ∎
Remark 8.30.
Note that $Z(\operatorname{End}_{L\rtimes G}(V))\cong L_{W}$, and so the field $L_{W}$ only depends on the $\Gamma$-orbit of $W$. This can also be seen directly, as the $L_{W}$ for $W$ in a fixed $\Gamma$-orbit are all Galois conjugate in $L$.
9. Semilinear Schur Index
Definition 9.1.
For $W\in\operatorname{Ind}_{L}(H)$, set $m_{K}^{L}(W)\coloneqq m(V)$ for the unique corresponding $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$.
As a direct consequence of Theorem 8.20 we have the following.
Corollary 9.2.
Let $W\in\operatorname{Ind}_{L}(H)$. Then $W$ is extendable to a semilinear representation of $G$ if and only if $\Gamma$ fixes $W$ and $m_{K}^{L}(W)=1$.
From the results of the previous section, the numbers $m_{K}^{L}(W)$ have the following properties. Recall that $G_{W}$ is the preimage of $\Gamma_{W}$ in $G$, and we write $L_{W}\coloneqq L^{\Gamma_{W}}$. For any intermediate field $E$ with $K\leq E\leq L$, we write $\Gamma_{E}$ for the preimage of $\operatorname{Gal}(L/E)$ under the isomorphism $\Gamma\xrightarrow{\sim}\operatorname{Gal}(L/K)$.
Proposition 9.3.
Suppose that $W\in\operatorname{Ind}_{L}(H)$, and set $\mathcal{O}(W)\coloneqq\bigoplus_{\gamma\in\Gamma/\Gamma_{W}}\gamma*W$. Then:
  1. (1)
    $\mathcal{O}(W)^{\oplus m_{K}^{L}(W)}=V|_{H}$ for a unique $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$,
  2. (2)
    $m_{K}^{L}(W)$ is the unique $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ extends to an indecomposable object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$,
  3. (3)
    $m_{K}^{L}(W)$ is the smallest $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ extends to an object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$,
  4. (4)
    For any $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $V|_{H}$ is a multiple of $m_{K}^{L}(W)$.
We furthermore have that
  1. (5)
    $m_{K}^{L}(W)=m_{K}^{L}(g*W)$ for any $g\in G$.
  2. (6)
    $m_{K}^{L}(W)\mid|\Gamma_{W}|$.
  3. (7)
    $m_{L_{W}}^{L}(W)=m_{K}^{L}(W)$.
In particular, applying the above to $G_{W}\twoheadrightarrow\Gamma_{W}$,
  1. (1’)
    $W^{\oplus m_{L_{W}}^{L}(W)}=U|_{H}$ for a unique $U\in\operatorname{Ind}_{L}^{\rtimes}(G_{W})$,
  2. (2’)
    $m_{L_{W}}^{L}(W)$ is the unique $m\geq 1$ such that $W^{\oplus m}$ extends to an indecomposable object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G_{W})$,
  3. (3’)
    $m_{L_{W}}^{L}(W)$ is the smallest $m\geq 1$ such that $W^{\oplus m}$ extends to an object of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G_{W})$,
  4. (4’)
    For any $U\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G_{W})$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $U|_{H}$ is a multiple of $m_{L_{W}}^{L}(W)$.
Now suppose that $K\subset E\subset L$ is an intermediate field. Then
  1. (8)
    $m_{E}^{L}(W)\mid m_{K}^{L}(W)$,
  2. (9)
    $m_{K}^{L}(W)\mid m_{E}^{L}(W)\cdot[\Gamma_{W}:\Gamma^{E}_{W}]$, and
  3. (10)
    $m_{K}^{L}(W)\leq m_{E}^{L}(W)\cdot[E:K]$.
If additionally either $E$ or $L_{W}$ is Galois over $E\cap L_{W}$, then $\Gamma_{W}:\Gamma^{E}_{W}]=[E:E\cap L_{W}]$, so
  1. (11)
    $m_{K}^{L}(W)\mid m_{E}^{L}(W)\cdot[E:E\cap L_{W}]$, and in particular
  2. (12)
    $m_{K}^{L}(W)\mid m_{E}^{L}(W)\cdot[E:K]$.
Proof.
Points (1), (2), (3), (4), (5) and (6) follow from Theorem 8.20 and Corollary 8.6. For point (7), we may take $S$ to be $G_{W}$ in Proposition 8.24, to obtain that $m_{L_{W}}^{L}(W)\mid m_{K}^{L}(W)$ and $m_{K}^{L}(W)\mid m_{L_{W}}^{L}(W)\cdot[\Gamma_{W}:(\Gamma_{W})_{W}]=m_{L_{W}}^{L}(W)$.
Now suppose that $E$ is an intermediate field between $K$ and $L$. Taking $S\coloneqq G^{E}$, the preimage of $\Gamma^{E}$ in $G$, we have that $m_{E}^{L}(W)\mid m_{K}^{L}(W)$ and $m_{K}^{L}(W)\mid m_{E}^{L}(W)\cdot[\Gamma_{W}:\Gamma^{E}_{W}]$. Furthermore,
$\Gamma:\Gamma_{W}][\Gamma_{W}:\Gamma_{W}^{E}]=[\Gamma:\Gamma^{E}][\Gamma^{E}:\Gamma^{E}_{W}],$
and $\Gamma^{E}:\Gamma^{E}_{W}]\leq[\Gamma:\Gamma_{W}]$, hence $\Gamma_{W}:\Gamma_{W}^{E}]\leq[\Gamma:\Gamma^{E}]=[E:K]$.
Suppose now that $E$ or $L_{W}$ is Galois over $E\cap L_{W}$. Then in the square
$\Gamma^{E}_{W}$$\Gamma^{E}$$\Gamma_{W}$$\Gamma^{E\cap L_{W}}$}}
at least one of $\Gamma_{W}$ or $\Gamma^{E}$ is normal in $\Gamma^{E\cap L_{W}}$, and so by the second isomorphism theorem,
$\Gamma_{W}:\Gamma^{E}_{W}]=[\Gamma^{E\cap L_{W}}:\Gamma^{E}]=[E:E\cap L_{W}].\qed$
As a direct consequence of Proposition 8.22, we also have the following.
Corollary 9.4.
If $L$ is a finite field, $m_{K}^{L}(W)=1$ for any $W\in\operatorname{Irr}_{L}(H)$.
The following example shows how the numbers $m_{K}^{L}(W)$ capture arithmetic information.
Example 9.5.
Let $K$ be any field, and let $L=K(\sqrt{d})$, where $d\in K^{\times}\setminus(K^{\times})^{2}$. Taking $G=C_{4}=\langle y\rangle$, there is a natural homomorphism
$C_{4}\twoheadrightarrow\Gamma\coloneqq\operatorname{Gal}(K(\sqrt{d})/K),$
with kernel $C_{2}=\langle x\rangle$, where $y^{2}=x$. Let $W$ be the $1$-dimensional $L$-representation of $C_{2}$ where $x$ acts by scalar multiplication by 1-1. Then, $m_{K}^{L}(W)$ is either $1$ or $2$ by Proposition 9.3(6). Explicitly, using the matrix description of Section 2.1, $W$ extends to a semilinear representation of $G$ over $K(\sqrt{d})$ if and only if there is an $a=u+v\sqrt{d}\in K(\sqrt{D})^{\times}$ (defining the matrix of the action of $y$) with
$(u+v\sqrt{d})(u-v\sqrt{d})=-1,$
this condition arising from the relation that $y^{2}=x$. In particular, as $W$ is fixed by the action of $\Gamma$, we see from Corollary 9.2 that $m_{K}^{L}(W)=1$ if and only if $1=u^{2}-dv^{2}$ for some $u,v\in K$.
This can be extended to produce arbitrarily large values of $m_{K}^{L}(W)$ (see also Corollary 16.2).
Example 9.6.
Let $p$ be a prime, and let $L/K$ be a degree $p$ Galois field extension. For any $n\geq 1$, there is a quotient map
$C_{np}\twoheadrightarrow\operatorname{Gal}(L/K),$
with kernel $C_{n}=\langle x\rangle\subset C_{np}=\langle y\rangle$, where $y^{p}=x$. Then for any $1$-dimensional representation $W$ of $H$ valued in $K$, then similarly to above, $m_{K}^{L}(W)\in\{1,p\}$, and $m_{K}^{L}(W)=1$ if and only if $N_{K}^{L}(a)=\chi_{W}(x)$ for some $a\in L^{\times}$.
10. Recovery of the Classical Schur Index
Suppose now that $H$ is a group, and $L/K$ is a Galois extension of fields. Then, as in Section 6, we may take $G\coloneqq H\times\operatorname{Gal}(L/K)$ and $G\twoheadrightarrow\operatorname{Gal}(L/K)$ to be the natural projection, and consider the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. By Proposition 6.1 there are equivalences
$\operatorname{\mathbf{Rep}}_{K}(H)$$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\operatorname{Gal}(L/K))$$\scriptstyle{L\otimes_{K}-}$$\scriptstyle{(-)^{\operatorname{Gal}(L/K)}}$
under which induction
$(-)_{L}\colon\operatorname{\mathbf{Rep}}_{K}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}(H)$
corresponds to restriction
$(-)|_{H}\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\operatorname{Gal}(L/K))\rightarrow\operatorname{\mathbf{Rep}}_{L}(H),$
by Lemma 6.2, and restriction
$(-)|_{K[H]}\colon\operatorname{\mathbf{Rep}}_{L}(H)\rightarrow\operatorname{\mathbf{Rep}}_{K}(H)$
corresponds to induction
$\operatorname{Ind}_{H}^{H\times\operatorname{Gal}(L/K)}\colon\operatorname{\mathbf{Rep}}_{L}(H)\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(H\times\operatorname{Gal}(L/K))$
by Lemma 6.3. For example, given $\in\operatorname{\mathbf{Rep}}_{L}(H), VV$ is induced from a $K$-linear representation of $H$ if and only if $V$ extends to a semilinear representation of $H\times\operatorname{Gal}(L/K)$.
As $H$ commutes with $\operatorname{Gal}(L/K)$, the action of $\operatorname{Gal}(L/K)$ of Definition 8.16 on $\operatorname{\mathbf{Rep}}_{L}(H)/\cong$ simplifies: for $\gamma\in\operatorname{Gal}(L/K)$ and $W\in\operatorname{\mathbf{Rep}}_{L}(H)$, $\gamma*W$ is $W$ as a $K$-vector space, with same action of $H$, and $L$-linear structure $\lambda\cdot w\coloneqq\gamma^{-1}(\lambda)w$. In terms of matrices, if $W$ has matrix representation $(A_{g})_{g\in G}$, then $\gamma*W$ has matrix representation $({}^{\gamma\!}A_{g})_{g\in G}$ (in the notation of Section 2.1).
From Theorems A and B, we recover the following relationship between $\operatorname{\mathbf{Rep}}_{K}(H)$ and $\operatorname{\mathbf{Rep}}_{L}(H)$.
Corollary 10.1.
Let $V_{1},V_{2}\in\operatorname{\mathbf{Rep}}_{K}(H)$. Then:
  • If $V_{1,L}\cong V_{2,L}$ in $\operatorname{\mathbf{Rep}}_{L}(H)$ then $V_{1}\cong V_{2}$.
  • The natural map
    $L\otimes_{K}\operatorname{Hom}_{K[H]}(V_{1},V_{2})\rightarrow\operatorname{Hom}_{L[H]}(V_{1,L},V_{2,L})$
    is an isomorphism.
  • In particular, if $V\in\operatorname{\mathbf{Rep}}_{K}(H)$ then the natural map
    $L\otimes_{K}\operatorname{End}_{K[H]}(V)\rightarrow\operatorname{End}_{L[H]}(V_{L})$
    is an $L$-algebra isomorphism.
Let $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ and let $W\in\operatorname{Ind}_{L}(H)$ be an indecomposable direct summand of $V_{L}$. Then:
  • $V_{L}$ and $W|_{K[H]}$ are described by
    $V_{L}\cong\bigoplus_{\gamma\in\operatorname{Gal}(L/K)/\operatorname{Gal}(L/K)_{V}}(\gamma*W)^{\overline{m}_{K}^{L}(W)},\qquad W|_{K[H]}\cong V^{\oplus|\operatorname{Gal}(L/K)_{W}|/\overline{m}_{K}^{L}(W)}$
    for some integer $\overline{m}_{K}^{L}(W)\geq 1$. In particular, $\overline{m}_{K}^{L}(W)\mid|\operatorname{Gal}(L/K)_{W}|$.
  • $W$ is irreducible if and only if $V$ is irreducible.
  • $W|_{K[H]}\cong(\gamma*W)|_{K[H]}$ and $\overline{m}_{K}^{L}(W)=\overline{m}_{K}^{L}(\gamma*W)$ for any $\gamma\in\operatorname{Gal}(L/K)$.
  • Any $W\in\operatorname{Ind}_{L}(H)$ is a direct summand of $V_{L}$ for a unique $V\in\operatorname{Ind}_{K}(H)$. Sending $W$ to this unique $V$ defines bijections
    $\operatorname{Ind}_{K}(H)\leftrightarrow\operatorname{Ind}_{L}(H)/\operatorname{Gal}(L/K),\qquad\operatorname{Irr}_{K}(H)\leftrightarrow\operatorname{Irr}_{L}(H)/\operatorname{Gal}(L/K).$
10.1. Classical Schur Index
For $L/K$ a finite Galois extension and $W\in\operatorname{Ind}_{L}(H)$, we call the number $\overline{m}_{K}^{L}(W)$ appearing in Corollary 10.1 the (classical) Schur index of $W$ over $K$. When $W$ is irreducible (which is the setting in which Schur indices are usually considered) this agrees by definition with the usual definition of the Schur index [HUP, Thm. 38.1].
We following record how Proposition 9.3 recovers the basic properties of the Schur index, and further generalises them to indecomposable modules.
Corollary 10.2.
Let $W\in\operatorname{Ind}_{L}(H)$. Then $W$ is defined over $K$ if and only if $\operatorname{Gal}(L/K)$ fixes $W$ and $\overline{m}_{K}^{L}(W)=1$.
Corollary 10.3.
Suppose that $W\in\operatorname{Ind}_{L}(H)$, and set $\mathcal{O}(W)\coloneqq\bigoplus_{\gamma\in\operatorname{Gal}(L/K)/\operatorname{Gal}(L/K)_{W}}\gamma*W$. Then:
  1. (1)
    $\mathcal{O}(W)^{\oplus\overline{m}_{K}^{L}(W)}=V_{L}$ for a unique $V\in\operatorname{Ind}_{K}(H)$,
  2. (2)
    $\overline{m}_{K}^{L}(W)$ is the unique $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ is induced from an indecomposable object of $\operatorname{\mathbf{Rep}}_{K}(H)$,
  3. (3)
    $\overline{m}_{K}^{L}(W)$ is the smallest $m\geq 1$ such that $\mathcal{O}(W)^{\oplus m}$ is induced from an object of $\operatorname{\mathbf{Rep}}_{K}(H)$,
  4. (4)
    For any $V\in\operatorname{Irr}_{K}(H)$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $V_{L}$ is a multiple of $\overline{m}_{K}^{L}(W)$.
We furthermore have that
  1. (5)
    $\overline{m}_{K}^{L}(W)=\overline{m}_{K}^{L}(\gamma*W)$ for any $\gamma\in\operatorname{Gal}(L/K)$.
  2. (6)
    $\overline{m}_{K}^{L}(W)\mid|\operatorname{Gal}(L/K)_{W}|$.
  3. (7)
    $\overline{m}_{L_{W}}^{L}(W)=\overline{m}_{K}^{L}(W)$.
In particular, applying the above to the Galois extension $L/L_{W}$,
  1. (1’)
    $W^{\oplus\overline{m}_{L_{W}}^{L}(W)}=U_{L}$ for a unique $U\in\operatorname{Ind}_{L_{W}}(H)$,
  2. (2’)
    $\overline{m}_{L_{W}}^{L}(W)$ is the unique $m\geq 1$ such that $W^{\oplus m}$ is induced from an indecomposable object of $\operatorname{\mathbf{Rep}}_{L_{W}}(H)$,
  3. (3’)
    $\overline{m}_{L_{W}}^{L}(W)$ is the smallest $m\geq 1$ such that $W^{\oplus m}$ is induced from an object of $\operatorname{\mathbf{Rep}}_{L_{W}}(H)$,
  4. (4’)
    For any $U\in\operatorname{\mathbf{Rep}}_{L_{W}}(H)$, the number of times $W$ appears as a direct summand in the Krull-Remak-Schmidt decomposition of $U_{L}$ is a multiple of $\overline{m}_{L_{W}}^{L}(W)$.
Now suppose that $K\subset E\subset L$ is an intermediate field. Then
  1. (8)
    $\overline{m}_{E}^{L}(W)\mid\overline{m}_{K}^{L}(W)$,
  2. (9)
    $\overline{m}_{K}^{L}(W)\mid\overline{m}_{E}^{L}(W)\cdot[\operatorname{Gal}(L/L_{W}):\operatorname{Gal}(L/(E\cdot L_{W}))]$, and
  3. (10)
    $\overline{m}_{K}^{L}(W)\leq\overline{m}_{E}^{L}\cdot[E:K]$.
If additionally either $E$ or $L_{W}$ is Galois over $E\cap L_{W}$, then
$\operatorname{Gal}(L/L_{W}):\operatorname{Gal}(L/(E\cdot L_{W}))]=[E:E\cap L_{W}],$
thus
  1. (11)
    $\overline{m}_{K}^{L}(W)\mid\overline{m}_{E}^{L}\cdot[E:E\cap L_{W}]$, and in particular
  2. (12)
    $\overline{m}_{K}^{L}(W)\mid\overline{m}_{E}^{L}\cdot[E:K]$.
The above properties are all special cases of properties of all semilinear Schur indices. We now highlight some properties that the classical Schur indices $\overline{m}_{K}^{L}(W)$ satisfy, which are not true for the semilinear Schur indices $m_{K}^{L}(W)$.
10.1.1. Galois Splitting Field
When $H$ is a finite group (no assumptions on the characteristic of $L$), the field $L_{W}$ admits a nicer description. For a character $\chi$ of an irreducible representation of $H$ over $L$, we write $K(\chi)$ for the subfield of $L$ generated by the set $\{\chi(h)\mid h\in H\}$.
Lemma 10.4.
Suppose that $H$ is finite, and $W\in\operatorname{Irr}_{L}(H)$. Then $L_{W}=K(\chi_{W})$.
Proof.
If $\gamma\in\operatorname{Gal}(L/K)$ fixes $W$, then $\gamma*\chi_{W}=\chi_{W}$, where $(\gamma*\chi_{W})(h)\coloneqq\gamma(\chi_{W}(h))$ for any $h\in H$. In particular, $\gamma$ fixes $K(\chi)$, and so $K(\chi_{W})\subset L_{W}$. Conversely, if $\gamma$ fixes $K(\chi_{W})$, then $\gamma*\chi_{W}=\chi_{W}$, and so, as $\chi_{\gamma*W}=\gamma*\chi_{W}$, then $W\cong\gamma*W$ by [ISA, Cor. 9.22]. In particular, $\operatorname{Gal}(L/K(\chi))\subset\operatorname{Gal}(L/K)_{W}$, and so $L_{W}\subset K(\chi_{W})$. ∎
Remark 10.5.
In particular, when $H$ is finite and $W\in\operatorname{Irr}_{L}(H)$, $L_{W}=K(\chi_{W})$ is normal over $K$, being a subfield of a cyclotomic extension. In particular, from Corollary 10.3,
$\overline{m}_{K}^{L}(W)\mid\overline{m}_{E}^{L}(W)\cdot[E:E\cap K(\chi_{W})]$
for any intermediate field $K\subset E\subset L$.
10.1.2. Divisibility by Dimension
The following result, amongst other things, implies that any linear character of $H$ over $L$ has Schur index $1$. This makes sense, as any such representation fixed by $\operatorname{Gal}(L/K)$ is easily seen to be defined over $K$. The proof does not extend to the non-split setting, and indeed the statement for semilinear Schur indices is false. This can be seen from Example 9.5, taking $L/K$ to be ${\mathbb{Q}}(i)/{\mathbb{Q}}$, which exhibits linear characters of $H$ over $L$ with Schur index $2$.
Proposition 10.6.
Suppose that $W\in\operatorname{Irr}_{L}(H)$, and let $D_{W}\coloneqq\operatorname{End}_{L[H]}(W)$, a (potentially non-central) division algebra over $L$. Then
$\overline{m}_{K}^{L}(W)\mid\frac{\dim_{L}(W)}{\dim_{L}D_{W}}.$
Proof.
Let $V\in\operatorname{Irr}_{K}(H)$ correspond to $W$, and set $m\coloneqq\overline{m}_{K}^{L}$. Writing $D_{V}\coloneqq\operatorname{End}_{K[G]}(V)$, we have that $\dim_{K}D_{V}\mid\dim_{K}V$ as $V$ is a free module over $D_{V}$. From Corollary 10.1 we have that
$V_{L}=\oplus_{\gamma\in\Gamma/\Gamma_{W}}(\gamma*W)^{m}$
and thus
$\dim_{K}D_{V}=\dim_{L}\operatorname{End}_{L[H]}(V_{L})=\dim_{L}\prod_{\Gamma/\Gamma_{W}}M_{m}(D_{W})=[\Gamma:\Gamma_{W}]\cdot m^{2}\cdot\dim_{L}D_{W}.$
From $\dim_{K}D_{V}\mid\dim_{K}V$ and noting that $\dim_{K}V=\dim_{L}V_{L}$ we have that
$\Gamma:\Gamma_{W}]\cdot m^{2}\cdot\dim_{L}D_{W}\mid[\Gamma:\Gamma_{W}]\cdot m\cdot\dim_{L}W,$
and therefore $m\cdot\dim_{L}D_{W}\mid\dim_{L}W$ as required. ∎
11. Base Change for Semilinear Representations
We would like to better understand the division algebras $\operatorname{End}_{L\rtimes G}(V)$ for $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, and more specifically their splitting behaviour under base extension. In this section only we briefly relax the condition, in place since the start of Section 8.6, that $L$ is a field - this is only needed for the final result of this Section, Corollary 11.5. We also suppose in this section that $K^{\prime}$ is an arbitrary, potentially infinite, field extension of $K$.
In this section we consider the base change functor
$(-)_{L^{\prime}}\colon\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{L^{\prime}}^{\rtimes}(G)$
as defined in Section 3, where $L^{\prime}\coloneqq K^{\prime}\otimes_{K}L$, which has Galois action of $\Gamma$ through the action of $\Gamma$ on $L$. We let $f$ denote a primitive idempotent of $K^{\prime}\otimes_{K}F$, which unlike $F$ need no longer be connected, and set $F^{\prime}\coloneqq f\cdot(K^{\prime}\otimes_{K}F)\hookrightarrow L^{\prime}$. Note that $F^{\prime}\neq K^{\prime}\otimes_{K}F$ in general, unlike what the notion might suggest, but instead $F^{\prime}$ defines a connected component of the extension $L^{\prime}/K^{\prime}$, just as $F$ defines a connected component of the extension $L/K$.
We first show that base change is compatible with the induction and restriction functors appearing in Theorem 8.20.
Proposition 11.1.
The diagrams
$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{L^{\prime}}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{L^{\prime}}^{\rtimes}(G)$$\operatorname{\mathbf{Rep}}_{F}(H)$$\operatorname{\mathbf{Rep}}_{F^{\prime}}(H)$$\operatorname{\mathbf{Rep}}_{F}(H)$$\operatorname{\mathbf{Rep}}_{F^{\prime}}(H)$$\scriptstyle{(-)_{L^{\prime}}}$$\scriptstyle{(-)_{L^{\prime}}}$$\scriptstyle{e\cdot(-)}$$\scriptstyle{f\cdot(-)}$$\scriptstyle{(L\rtimes G)\otimes_{L[H]}-}$$\scriptstyle{(-)_{F^{\prime}}}$$\scriptstyle{(L^{\prime}\rtimes G)\otimes_{L^{\prime}[H]}-}$$\scriptstyle{(-)_{F^{\prime}}}$
both commute up to natural isomorphism.
Proof.
For the first diagram, if $W\in\operatorname{\mathbf{Rep}}_{F}(H)$, there is a natural morphism between both compositions
$L^{\prime}\otimes_{L}((L\rtimes G)\otimes_{L[H]}W)\rightarrow(L^{\prime}\rtimes G)\otimes_{L^{\prime}[H]}(F^{\prime}\otimes_{F}W),\qquad\lambda\otimes x\otimes w\mapsto\lambda x\otimes 1\otimes w.$
Every pure tensor of the right-hand term can be put in the form $x\otimes 1\otimes w$, and therefore this is surjective. If $W$ has rank $n\geq 0$ over $F$, both sides have rank $|\Gamma|n$ over $L^{\prime}$, so the map is an isomorphism.
For the second diagram, if $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, there is a natural morphism between both compositions
$F^{\prime}\otimes_{F}(e\cdot V)\rightarrow f\cdot(L^{\prime}\otimes_{L}V),\qquad\lambda\otimes v\mapsto\lambda\otimes v.$
Because $L^{\prime}/L$ is free, $f\cdot(L^{\prime}\otimes_{L}V)=(f\cdot L^{\prime})\otimes_{L}V$. Furthermore, as $fe=f$ in $F^{\prime}$, any pure tensor of $(f\cdot L^{\prime})\otimes_{L}V$ is of the form $\lambda\otimes v$, for $v\in e\cdot V$, and so the map is surjective. If $W$ has rank $n\geq 0$ over $L$, then both sides have rank $n$ over $F^{\prime}$, and so the map is an isomorphism. ∎
Corollary 11.2.
Suppose that $W\in\operatorname{Ind}_{F}(H)$, and that $W_{F^{\prime}}=W_{1}^{n_{1}}\oplus\cdots\oplus W_{r}^{n_{r}}$ is a direct sum decomposition of $W_{F^{\prime}}$, where $W_{i}\in\operatorname{Ind}_{F^{\prime}}(H)$ are pairwise non-isomorphic. Suppose that $V\in\operatorname{Ind}_{L}^{\rtimes}(G)$ corresponds to $W$, and $V_{i}\in\operatorname{Ind}_{L^{\prime}}^{\rtimes}(G)$ corresponds to $W_{i}$ for i=1,,ri=1,...,r. Then
$V_{L^{\prime}}\cong\bigoplus_{i=1}^{r}V_{i}^{\frac{|\Gamma_{f,W_{i}}|\cdot m(V)\cdot n_{i}}{|\Gamma_{e,W}|\cdot m(V_{i})}}.$
In particular, when $W$ is absolutely irreducible, if $V_{0}\in\operatorname{Irr}_{L^{\prime}}^{\rtimes}(G)$ corresponds to $W_{F^{\prime}}$, then
$V_{L^{\prime}}\cong V_{0}^{\frac{|\Gamma_{f,W}|\cdot m(V)}{|\Gamma_{e,W}|\cdot m(V_{0})}}$
and so $m(V_{0})\cdot[\Gamma_{e,W}:\Gamma_{f,W}]\mid m(V)$. If additionally, $W$ is fixed by $\Gamma$, then
$\Gamma_{e,W}:\Gamma_{f,W}]=[\Gamma_{e}:\Gamma_{f}]=\deg_{F^{\prime}}(K^{\prime}\otimes_{K}F),$
thus
$V_{L^{\prime}}\cong V_{0}^{\frac{m(V)}{\deg_{F^{\prime}}(K^{\prime}\otimes_{K}F)\cdot m(V_{0})}}$
and so $m(V_{0})\cdot\deg_{F^{\prime}}(K^{\prime}\otimes_{K}F)\mid m(V)$.
Proof.
From the first diagram of Proposition 11.1 and Theorem 8.20 we have that
$V_{L^{\prime}}^{\frac{|\Gamma_{e,W}|}{m(V)}}\cong\bigoplus_{i=1}^{r}V_{i}^{\frac{|\Gamma_{f,W_{i}}|\cdot n_{i}}{m(V_{i})}},$
from which the description of $V_{L^{\prime}}$ follows from Lemma 2.14. For the second statement, note that $\Gamma_{f,W}=\Gamma_{f,W_{F^{\prime}}}$: if $\gamma\in\Gamma$ and $\gamma*W_{F^{\prime}}\cong W_{F^{\prime}}$, then as $\gamma*W_{F^{\prime}}=(\gamma*W)_{F^{\prime}}$, $(\gamma*W)_{F^{\prime}}\cong W_{F^{\prime}}$, hence $\gamma*W\cong W$ by Corollary 10.1. ∎
Remark 11.3.
If $F^{\prime}/F$ is a finite Galois extension of fields, then from Corollary 10.1 we have that
$W_{F^{\prime}}=\bigoplus_{\gamma\in\operatorname{Gal}(F^{\prime}/F)/\operatorname{Gal}(F^{\prime}/F)_{W_{0}}}(\gamma*W_{0})^{\overline{m}_{F}^{F^{\prime}}(W)}$
where $W_{0}$ is taken to be any of the $W_{i}$. In particular,
$n_{1}=\cdots=n_{s}=\overline{m}_{F}^{F^{\prime}}(W_{0})\ \text{ and }\ s=[\operatorname{Gal}(F^{\prime}/F):\operatorname{Gal}(F^{\prime}/F)_{W_{0}}].$
Corollary 11.4.
Suppose that $W\in\operatorname{Irr}_{L}(H)$ is fixed by $\Gamma$ and $\operatorname{End}_{L[H]}(W)=L$. Let $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$ correspond to $W$, and set $D\coloneqq\operatorname{End}_{L\rtimes G}(V)$, a central division algebra over $K$. Then
$\operatorname{Deg}(D_{K^{\prime}})=m_{K^{\prime}}^{F^{\prime}}(W_{F^{\prime}}),$
and in particular $\operatorname{Deg}(D_{K^{\prime}})\cdot\deg_{F^{\prime}}(K^{\prime}\otimes_{K}F)\mid m_{K}^{F}(W)$ and $\operatorname{Deg}(D_{K^{\prime}})\mid[F^{\prime}:K^{\prime}]$.
Proof.
From Lemma 3.6, there is an isomorphism
$D_{K^{\prime}}\xrightarrow{\sim}\operatorname{End}_{L^{\prime}\rtimes G}(V_{L^{\prime}}),$
and from the equivalence of Proposition 7.1 an isomorphism
$\operatorname{End}_{L^{\prime}\rtimes G}(V_{L^{\prime}})\xrightarrow{\sim}\operatorname{End}_{F^{\prime}\rtimes G_{f}}(f\cdot V_{L^{\prime}}).$
Considering the tuple $(G_{f},H,F^{\prime}/K^{\prime})$, let $V_{0}\in\operatorname{Irr}_{F^{\prime}}^{\rtimes}(G_{f})$ correspond to $W_{F^{\prime}}$. We have by Corollary 11.2 that
$f\cdot V_{L^{\prime}}=V_{0}^{\frac{m_{K}^{F}(W)}{[F^{\prime}:F]\cdot m_{K^{\prime}}^{F^{\prime}}(W_{F^{\prime}})}},$
and therefore that
$\operatorname{Deg}(\operatorname{End}_{F^{\prime}\rtimes G_{f}}(f\cdot V_{L^{\prime}}))=\operatorname{Deg}(\operatorname{End}_{F^{\prime}\rtimes G_{f}}(V_{0}))=m_{K^{\prime}}^{F^{\prime}}(W_{F^{\prime}}),$
the final equality by Proposition 8.29, as $\operatorname{End}_{F^{\prime}[H]}(W_{F^{\prime}})=F^{\prime}$ and $W_{F^{\prime}}$ is fixed by $\Gamma_{f}$. The divisibility relations come from Proposition 9.3 and Corollary 11.2. ∎
We can use this to give a local-global principle in our context.
Corollary 11.5.
Suppose that $L/K$ is a Galois extension of number fields, $W\in\operatorname{Irr}_{L}(H)$ is fixed by $\Gamma$ and $\operatorname{End}_{L[H]}(W)=L$. For each place of $K$, let $w$ be a place of $L$ over $v$, and let $G_{w}$ be the stabiliser of of $L_{w}$ in $K_{v}\otimes_{K}L$. Then
$m_{K}^{F}(W)=\operatorname{lcm}_{v}m_{K_{v}}^{L_{w}}(W_{L_{w}}).$
In particular, $W$ extends to a semilinear representation of $G$ over $L$ if and only if $W_{L_{w}}$ extends to a semilinear representation of $G_{w}$ for any place $v$ of $K$.
Proof.
Letting $D$ denote the central simple $K$-algebra $\operatorname{End}_{L\rtimes G}(V)$, $D$ is split over $K$ if and only if $m_{K}^{L}(W)=1$, as $\operatorname{Deg}(D)=m_{K}^{L}(W)$ by Proposition 8.29. For any place $v$ of $K$, similarly $m_{K_{v}}^{L_{w}}(W_{L_{v}})=1$ if and only if $W_{L_{w}}$ extends to a semilinear representation of $G_{w}$. We have that $\operatorname{Deg}(D_{K_{v}})=m_{K_{v}}^{L_{w}}(W_{L_{v}})$ by Corollary 11.4, noting that $L_{w}$ is a connected component of $K_{v}\otimes_{K}L$, so plays the role of $F^{\prime}$. The statement then follows from the fact that $\operatorname{Deg}(D)$ is the lowest common multiple of the $\operatorname{Deg}(D_{K_{v}})$ by [milneCFT, Thm. VIII.2.6]. ∎
We can use the results of this section to compute endomorphism rings of semilinear representations explicitly.
Example 11.6.
We continue Example 9.5, and suppose additionally that $K={\mathbb{Q}}$, so $L={\mathbb{Q}}(\sqrt{d})$ for some $d\in{\mathbb{Q}}^{\times}\setminus({\mathbb{Q}}^{\times})^{2}$. The group $G=C_{4}$, with subgroup $H$, and natural quotient map
$C_{4}/C_{2}\xrightarrow{\sim}\Gamma\coloneqq\operatorname{Gal}({\mathbb{Q}}(\sqrt{d})/{\mathbb{Q}}).$
We take $W$ to be the non-trivial $1$-dimensional representation of $C_{2}$, and suppose that $m_{K}^{L}(W)=2$, so $D\coloneqq\operatorname{End}_{L\rtimes C_{4}}(V)$ for corresponding $V\in\operatorname{Irr}_{L}^{\rtimes}(C_{4})$. From Example 9.5, this is equivalent to the claim that $1\neq x^{2}-dy^{2}$ for any $x,y\in{\mathbb{Q}}^{\times}$. We wish to describe the quaternion algebra $D$ over $\mathbb{Q}$}.
For any place $v$ of $\mathbb{Q}$}, by Corollary 11.4 $D_{{\mathbb{Q}}_{v}}$ is split if and only if $m_{K_{v}}^{L_{w}}(W_{L_{w}})=1$, where $w$ is a place of $L$ above $v$. Set $L_{v}\coloneqq K_{v}\otimes_{K}L$.
When $L_{v}$ is a field, $L_{v}=K_{v}(i)$, and therefore by Example 9.5, $m_{K_{v}}^{L_{v}}(W_{L_{v}})=1$ if and only if $1\neq x^{2}-dy^{2}$ for any $x,y\in{\mathbb{Q}}_{v}^{\times}$.
When $L_{v}$ is not a field, or equivalently the place $v$ splits completely, we see that $m_{K_{v}}^{L_{w}}(W_{L_{w}})=1$ by Corollary 11.4, taking $F=K_{v}$ and $F^{\prime}=L_{w}$. We can rephrase this condition as follows. Writing $V^{v}\in\operatorname{Irr}_{L_{v}}^{\rtimes}(C_{4})$ for the irreducible semilinear representation of $C_{4}$ over $L_{v}$ corresponding to $W$, $m(V^{v})=m_{K_{v}}^{L_{w}}(W_{L_{w}})$ from the compatibility of Proposition 8.23. Similarly to Example 9.5, $m(V^{v})=1$ if and only if $W\cong f\cdot V^{v}$, if and only if there is an $a=x+\sqrt{d}y\in L_{v}$, $x,y\in K_{v}$, with
$f\cdot(x+\sqrt{d}y)(x-\sqrt{d}y)=f\cdot a\overline{a}=-1$
which, as $(x+iy)(x-iy)=x^{2}-dy^{2}\in K_{v}$, is the same as the condition that $x^{2}-dy^{2}=-1$ in $K_{v}$. Of course, as $m_{K_{v}}^{L_{w}}(W_{L_{w}})=1$ this is always satisfied, and indeed $\sqrt{d}\in L\subset L_{w}$ and $L_{w}=K_{v}$, so we can take $x=0$ and $y=1/\sqrt{d}$.
Altogether, we see that for any place $v$ of $K$, $D_{K_{v}}$ is split if and only if $1=x^{2}-dy^{2}$ for some $x,y\in K_{v}$. Therefore $D$ has the same splitting behaviour at all places as $(-1,d)_{{\mathbb{Q}}}$, so $D\cong(-1,d)_{{\mathbb{Q}}}$.
12. Character Theory for Semilinear Representations
Suppose in this section that $G$ is finite, and $|H|\in L^{\times}$, which by Corollary 8.21 is equivalent to the claim that $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple, or that $\operatorname{\mathbf{Rep}}_{L}(H)$ is semisimple. For $V,W\in\operatorname{\mathbf{Rep}}_{L}(H)$, we can consider the inner product of characters
$\langle\chi_{V},\chi_{W}\rangle=\frac{1}{|H|}\sum_{h\in H}\chi_{V}(h)\chi_{W}(h^{-1}),$
which satisfies [LOR, Lem. 3.20(b)]
$\langle\chi_{V},\chi_{W}\rangle=\dim_{L}\operatorname{Hom}_{L[H]}(V,W)\cdot 1_{L},$
where the multiplication by $1_{L}$ is to ensure this is interpreted as an equality in $L$.
Definition 12.1.
For $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, we define the character of $V$
$\chi_{V}\colon H\rightarrow L$
to be the character of the linear representation $V|_{H}\in\operatorname{\mathbf{Rep}}_{L}(H)$.
From Corollary 8.6 and Corollary 8.7 the following are immediate.
Corollary 12.2.
For $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, $\langle\chi_{V},\chi_{W}\rangle=\dim_{K}\operatorname{Hom}_{L\rtimes G}(V,W)\cdot 1_{L}$.
Corollary 12.3.
If $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ and $\operatorname{char}(L)=0$, then $V\cong W$ if and only if $\chi_{V}=\chi_{W}$.
We can also describe the characters of the irreducible objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Let $\operatorname{Fun}(H/\!/H,L)$ denote the $L$-algebra of $L$-valued functions on $H$ which are invariant under conjugation by $H$.
Definition 12.4.
We let $G$ act on $\operatorname{Fun}(H/\!/H,L)$ by
$g*f\coloneqq\sigma_{g}(f(g^{-1}-g))$
for any $g\in G$.
On $H$ this action is trivial, and therefore this induces an action of $\Gamma$ on $\operatorname{Fun}(H/\!/H,L)$ via the isomorphism $G/H\xrightarrow{\sim}\Gamma$. This is compatible with the action of $\Gamma$ on isomorphism classes of objects in $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$, in the sense that for any $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$,
$\chi_{g*V}=g*\chi_{V}.$
In particular, we see that the action of $\Gamma$ on $\operatorname{Fun}(H/\!/H,L)$ preserves the subset $\operatorname{Char}(H,L)$ of characters of objects of $\operatorname{\mathbf{Rep}}_{L}(H)$, and the smaller subset $\operatorname{Irr}(H,L)$ of characters of of irreducible representations of $H$ over $L$. Note that $\operatorname{Irr}(H,L)$ forms a basis for $\operatorname{Fun}(H/\!/H,L)$ whenever $L$ splits $H$ [LOR, Cor. 3.21], but need not span $\operatorname{Fun}(H/\!/H,L)$ in general. However, it is still true that the character of any irreducible representation of $H$ over $L$ is non-zero, irreducible representations are determined by their characters, and that the set $\operatorname{Irr}(H,L)$ is $L$-linearly independent [ISA, Cor. 9.22].
Definition 12.5.
For $\chi\in\operatorname{Irr}(H,L)$, we set $m_{K}^{L}(\chi)\coloneqq m_{K}^{L}(W)$, where $W\in\operatorname{Irr}_{L}(H)$ has $\chi_{W}=\chi$.
The numbers $m_{K}^{L}(\chi)$ inherit all the properties of $m_{K}^{L}(W)$ from Proposition 9.3. For example, they are constant on the orbits under the action of $\Gamma$, and satisfy certain divisibility relations.
Theorem 12.6.
Let ${\mathcal{C}}_{1},...,{\mathcal{C}}_{r}$ denote the orbits of the action of $\Gamma$ on $\operatorname{Irr}(H,L)$. Then there are $r$ irreducible semilinear representations $V_{1},...,V_{r}$ of $G$ over $L$, and for any k=1,,rk=1,...,r, $V_{k}$ has character
$\psi_{k}\coloneqq m_{K}^{L}(\chi_{k})\cdot\sum_{\chi\in{\mathcal{C}}_{k}}\chi,$
where $\chi_{k}$ is any representative of the orbit ${\mathcal{C}}_{k}$. Moreover, writing any $V,W\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ as
$V\cong V_{1}^{\oplus a_{1}}\oplus\cdots\oplus V_{r}^{a_{r}},\qquad W\cong V_{1}^{\oplus b_{1}}\oplus\cdots\oplus V_{r}^{b_{r}}$
for unique $a_{i},b_{i}\geq 1$, and setting $D_{k}\coloneqq\operatorname{End}_{L\rtimes G}(V_{k})$ for k=1,,rk=1,...,r,
$\langle\chi_{V},\chi_{W}\rangle=(a_{1}b_{1}\dim_{K}D_{1}+\cdots+a_{r}b_{r}\dim_{K}D_{r})\cdot 1_{L},$
and each
$\dim_{K}D_{k}\cdot 1_{L}=\langle\psi_{k},\psi_{k}\rangle=m_{K}^{L}(\chi_{k})^{2}\cdot|{\mathcal{C}}_{k}|\cdot\langle\chi_{k},\chi_{k}\rangle.$
Proof.
This follows directly from Theorem 8.20, Corollary 12.2, and the fact that $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ is semisimple. The only thing to note is that $\langle\chi,\chi\rangle=\langle\chi^{\prime},\chi^{\prime}\rangle$ for any $\chi,\chi^{\prime}$ in the same $\Gamma$-orbit on $\operatorname{Irr}(H,L)$, which follows from the fact that
$\operatorname{End}_{L[H]}(W)=\operatorname{End}_{L[H]}(g*W)$
for any $W\in\operatorname{\mathbf{Rep}}_{L}(H)$ and $g\in G$, which is clear from the definition of gWg*W. ∎
From Corollary 9.4, we can go further is when $L$ is finite.
Theorem 12.7.
Suppose that $L$ is finite. Then $m_{K}^{L}(\chi_{k})=1$ for all k=1,,rk=1,...,r.
In particular, the $\Gamma$-set $\operatorname{Irr}(H,L)$ completely describes the category $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ when $L$ is finite.
Corollary 12.8.
Suppose that $W\in\operatorname{Irr}_{L}^{\rtimes}(G)$. Then $\chi_{W}\neq 0$ if and only if $m(W)\cdot 1_{L}\neq 0$, and those characters $\chi_{W}$ where $\chi_{W}\neq 0$ are distinct and $L$-linearly independent. In particular, the association
$\operatorname{Irr}_{L}^{\rtimes}(G)\rightarrow\operatorname{Fun}(H/\!/H,L),\qquad W\mapsto\chi_{W},$
is injective if $L$ has characteristic 0 or if $L$ is finite.
Proof.
This follows directly from Theorem 12.6, Theorem 12.7, and the fact that $\chi_{V}$ for $V\in\operatorname{Irr}_{L}(H)$ are all non-zero, distinct, and $L$-linearly independent [ISA, Cor. 9.22]. ∎
13. Example: $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ for $L/K$ of Degree $2$ Revisited
In this section we give an example of the theory of the previous sections, to describe $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ for any field $L$ where $S_{3}$ acts on $L$ with kernel $H\coloneqq\langle(123)\rangle$ through an order two automorphism, so that $L/K$ has degree two. We will see that this recovers the classification of Section 5 in characteristic 0 deduced by more hands-on methods.
Let $x\in S_{3}\setminus H$ be any element, and let $\omega$ be a primitive third root of unity in $\overline{L}$.
13.1. Characteristic $\neq 3$
Let us first suppose that $\text{char}(K)\neq 3$, so $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ is semisimple by Corollary 8.21 and we may use the character theory of Section 12. From Theorem 8.20 we need to understand the character table of $H$ over $L$, and its action of $\Gamma=\operatorname{Gal}(L/K)$. There are three possible cases, which exactly match and explain the three cases encountered in Section 5.
Case 1: $\omega\in K$. Then the character table of $H$ is
[…]
and, as all entries are contained inside $K$, the action of $\operatorname{Gal}(L/K)$ sends $\chi\mapsto\chi^{x}$. Therefore, there are two orbits, $\{\chi_{1}\}$ and $\{\chi_{\omega},\chi_{\omega^{2}}\}$, corresponding to some $V_{1}$ and $V_{2}$ in $\operatorname{Irr}_{L}^{\rtimes}(S_{3})$ respectively. For each we can compute $m(V_{i})$.
For $V_{1}$, $\chi_{1}$ is the restriction of the trivial semilinear representation $L_{\text{triv}}$. In particular, $\chi_{1}$ is extendable to a semilinear representation of $S_{3}$ over $L$, and so $m(V_{1})=1$. Furthermore, $V_{1}=L_{\text{triv}}$.
For $V_{2}$, note that $|\Gamma_{\chi_{\omega}}|=1$, and so $m(V_{2})=1$ as $m(V_{2})\mid|\Gamma_{\chi_{\omega}}|$ by Theorem A. What this means concretely is that $\chi_{\omega}+\chi_{\omega^{2}}$ uniquely extends to a semilinear representation, $V_{2}$, of $S_{3}$ over $L$.
Because $L$ splits $H$, we can also compute the endomorphism rings of $V_{1}$ and $V_{2}$. We have
$\operatorname{End}_{L\rtimes S_{3}}(V_{1})\cong L^{\operatorname{Gal}(L/K)}=K$
by Proposition 8.29, and similarly
$\operatorname{End}_{L\rtimes S_{3}}(V_{2})\cong L^{\operatorname{Gal}(L/L)}=L.$
Note that this second isomorphism $L\xrightarrow{\sim}\operatorname{End}_{L\rtimes S_{3}}(V_{2})$ is not the natural action of $L$ on $V_{2}$.
Case 2: $\omega\in L\setminus K$. In this case, the character table of $H$ is the same as in Case 1, but now the natural Galois action of $\operatorname{Gal}(L/K)$ on $L$ swaps $\omega$ and $\omega^{2}$. In particular, the induced action of $\operatorname{Gal}(L/K)$ on the character table of $H$ is trivial. Therefore, all orbits of the action of $\operatorname{Gal}(L/K)$ are singletons, and there are three irreducible representations $V_{1}$, $V_{\omega}$ and $V_{\omega^{2}}$, which corresponds each to the characters $\chi_{1},\chi_{\omega}$ and $\chi_{\omega^{2}}$ respectively.
As $m(V_{\omega^{i}})\mid|\Gamma_{\chi_{\omega^{i}}}|$ by Theorem A, we have that $m(V_{\omega^{i}})\in\{1,2\}$. Similarly to Case 1 above, $\chi_{1}$ is the restriction of the trivial representation $L_{\text{triv}}$, and so $m(V_{1})=1$. To compute $m(V_{\omega})$ and $m(V_{\omega^{2}})$ however, this is as far as Theorem A alone gets us.
To determine $m(V_{\omega})$ and $m(V_{\omega^{2}})$ we use the induction functor
$L\otimes_{K}-\colon\operatorname{\mathbf{Rep}}_{K}(S_{3})\rightarrow\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$
of Section 4.1. The character table of $S_{3}$ over $K$ is
[…]
Let $U$ denote the irreducible $K$-linear representation of $S_{3}$ with character $\psi_{3}$, and suppose that
$L\otimes_{K}U=V_{1}^{\oplus a_{1}}\oplus V_{\omega}^{\oplus a_{\omega}}\oplus V_{\omega^{2}}^{\oplus a_{\omega^{2}}}$
for integers $a_{1},a_{\omega},a_{\omega^{2}}\geq 0$. Considering the character of the restriction to $H$, we therefore have that
$\chi_{\omega}+\chi_{\omega^{2}}=\psi_{3}|_{H}=a_{1}m(V_{1})\chi_{1}+a_{\omega}m(V_{\omega})\chi_{\omega}+a_{\omega^{2}}m(V_{\omega^{2}})\chi_{\omega^{2}}.$
Comparing coefficients, we have that $a_{\omega}m(V_{\omega})=1=a_{\omega^{2}}m(V_{\omega^{2}})$, and therefore $m(V_{\omega})=1=m(V_{\omega^{2}})$.
Because $L$ splits $H$, we can use Proposition 8.29 to see that the endomorphism rings of $V_{1}$, $V_{\omega}$ and $V_{\omega^{2}}$ are all $K$.
Case 3: $\omega\not\in L$. Now, for $\chi\coloneqq\chi_{\omega}+\chi_{\omega^{2}}$ the character table of $H$ over $L$ is
[…]
which has trivial action of $\operatorname{Gal}(L/K)$. Therefore, both irreducible characters $\chi_{1},\chi$ correspond to a unique irreducible semilinear representations $V_{1},V$ respectively of $S_{3}$ over $L$. As in Case 1, $V_{1}$ is the trivial semilinear representation $L_{\text{triv}}$ of $S_{3}$, with endomorphism ring $K$ and $m(V_{1})=1$, and similarly to above we either have that $m(V)=1$ or $m(V)=2$.
To determine $m(V)$, we again use the functor $L\otimes_{K}$-. Observing that $\chi=\psi_{3}|_{H}$, we see that $L\otimes_{K}U$ extends $\chi$, and therefore $m(V)=1$ and $V=L\otimes_{K}U$ is irreducible.
13.2. Characteristic 3
When $L$ has characteristic $3$, $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$ is no longer semisimple, and we need to understand $\operatorname{Ind}_{L}^{\rtimes}(S_{3})$. From Theorem 8.20, this is identified with $\operatorname{Ind}_{L}(H)/\operatorname{Gal}(L/K)$. The set $\operatorname{Ind}_{L}(H)$ consists of three indecomposble representations $V_{1},V_{2},V_{3}$ of dimensions $1,2,3$ respectively, and thus the action of $\operatorname{Gal}(L/K)$ on $\operatorname{Ind}_{L}(H)$ must be trivial as these dimensions are distinct. It remains to determine the semilinear Schur indices $m(V_{i})$, which we now show are all equal to $1$.
Each $V_{i}$ is defined over $K$, being given by the matrix representations where $(123)$ acts via
$(1),\qquad\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\qquad\begin{pmatrix}1&1&0\\ 0&1&1\\ 0&0&1\end{pmatrix}$
respectively. Each extends to a linear representation of $S_{3}$ over $K$, where $(12)$ acts via the matrices
$(1),\qquad\begin{pmatrix}1&0\\ 0&2\end{pmatrix},\qquad\begin{pmatrix}2&1&0\\ 0&1&1\\ 0&0&2\end{pmatrix}$
respectively, and therefore the same matrices define extensions $\tilde{V_{i}}$ of each $V_{i}$ to $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$.
In particular, $\tilde{V_{1}},\tilde{V_{2}},\tilde{V_{3}}$ is a complete list of all indecomposable objects of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(S_{3})$.
13.3. General Algorithm
It is natural to ask if one can always follow the same process used in this section to determine all $m_{K}^{L}(W)$ when $G$ is finite and $L$ has characteristic 0. In general, for an orbit $\mathcal{O}$} of the action of $\Gamma$ on $\operatorname{Irr}(H,L)$ we can consider
$\chi_{{\mathcal{O}}}=\sum_{\chi\in\mathcal{O}}\chi.$
Then for any $U\in\operatorname{Irr}_{K}(G)$, with character $\psi_{U}$, $(L\otimes_{K}U)|_{H}$ has character $\psi_{U}|_{H}$, and
$\psi_{U}|_{H}=\sum_{{\mathcal{O}}}a_{U,{\mathcal{O}}}\cdot\chi_{\mathcal{O}}$
for unique $a_{U,{\mathcal{O}}}\geq 0$. From Proposition 9.3(4) we have that $m_{K}^{L}(\mathcal{O})\mid a_{U,{\mathcal{O}}}$, and so
$m_{K}^{L}(\mathcal{O})\mid\operatorname{hcf}\{a_{U,{\mathcal{O}}}\mid U\in\operatorname{Irr}_{K}(G)\}$
where $m_{K}^{L}(\mathcal{O})$ is the semilinear Schur index of an element of $\mathcal{O}$. We also know that for a fixed orbit $\mathcal{O}$}, not all $a_{U,{\mathcal{O}}}$ can be zero by Corollary 4.2.
In the examples of this section, this divisibility relation always turned out to always be an equality. However, this is not true in general, as the following example shows.
Example 13.1.
Let $L/K$ be ${\mathbb{Q}}(\sqrt{2})/{\mathbb{Q}}$, $G=C_{4}=\langle y\rangle$ with natural surjection to $\Gamma=\operatorname{Gal}({\mathbb{Q}}(\sqrt{2})/{\mathbb{Q}})$ with kernel $H=\langle x\rangle$, $y^{2}=x$. The three irreducible characters of $C_{4}$ over $\mathbb{Q}$} are
[…]
The two orbits of $\Gamma$ on $\operatorname{Irr}_{L}(C_{2})$ are $\chi_{1}$ and $\chi_{-1}$, and we can compute that
$\psi_{1}|_{H}$ $=\chi_{1},$
$\psi_{2}|_{H}$ $=\chi_{1},$
$\psi_{3}|_{H}$ $=2\cdot\chi_{-1}.$
However $m_{K}^{L}(\chi_{-1})=1$, by Example 9.5, as $1=x^{2}-2y^{2}$ has a solution over $\mathbb{Q}$}, with $x=1,y=1$.
In particular, this algorithm, whilst still giving restrictions on the $m_{K}^{L}(W)$, won’t determine all semilinear Schur indices in general. A more precise process for determining the semilinear Schur indices $m_{K}^{L}(W)$ from linear representation-theoretic data when $L/K$ of degree two is the subject of the forthcoming work [RUMTAY4].
14. Conjugacy Classes and the Number of Semilinear Representations
In this section we suppose that
  1. (1)
    $G$ is finite, $L$ is a field, with $|H|\in L^{\times}$, and
  2. (2)
    $L=K(\mu_{n})$ for some $n\geq 1$ with $\text{exp}(H)\mid n$.
In this situation, there is a canonical embedding
$\epsilon\colon\Gamma\hookrightarrow({\mathbb{Z}}/n{\mathbb{Z}})^{\times}.$
which is determined by the property that $\gamma(\zeta)=\zeta^{\epsilon(\gamma)}$ for any $n$th root of unity $\zeta$ and all $\gamma\in\Gamma$. Using this embedding, there is a left action of $G\times G$ on $H$ by the formula:
$(g_{1},g_{2})*h=g_{1}h^{\epsilon(\sigma_{g_{2}}^{-1})}g_{1}^{-1},$
This induces an action of $G/H\times G/H$ and hence $\Gamma\times\Gamma$ on $\operatorname{Cl}(H)$. Similarly, there is a left action of $G\times G$ on $\operatorname{Irr}_{L}(H)$, where $(g_{1},g_{2})*V$ is $V$ as an abelian group but with $L\rtimes H$-module structure defined by
$\lambda h*v\coloneqq\sigma_{g_{2}^{-1}}(\lambda)\cdot(g_{1}^{-1}hg_{1})(v),$
which induces an action of $G/H\times G/H$ and hence $\Gamma\times\Gamma$ on $\operatorname{Irr}_{L}(H)$. The diagonal action is the action of $\Gamma$ on $\operatorname{Irr}_{L}(H)$ considered before. If $V\in\operatorname{Irr}_{L}(H)$ has character $\chi$, then we set
$(g_{1},g_{2})*\chi\coloneqq\chi_{(g_{1},g_{2})*V}=\sigma_{g_{2}}(\chi(g_{1}^{-1}-g_{1})).$
These actions are compatible in the following sense.
Lemma 14.1.
For any $\chi\in\operatorname{Irr}(H,L)$, $h\in H$ and $\underline{\gamma}\in\Gamma\times\Gamma$,
$(\underline{\gamma}*\chi)(h)=\chi(\underline{\gamma}^{-1}*h).$
Proof.
Expressing $\underline{\gamma}$ as the reduction of $(g_{1},g_{2})\in G\times G$, choose a basis of $V$ (the representation associated to $\chi$) where $g_{1}^{-1}hg_{1}$ acts diagonally. Then $\chi(g_{1}^{-1}hg_{1})=\zeta_{1}+\cdots+\zeta_{r}$ is the sum of $n$th roots of unity, using the assumption that $n$ divides the exponent of $H$. We may then compute that
$(\underline{\gamma}*\chi)(h)$ $=\sigma_{g_{2}}(\chi(g_{1}^{-1}hg_{1})),$
$=\sigma_{g_{2}}(\zeta_{1}+\cdots+\zeta_{r}),$
$=\zeta_{1}^{\epsilon(\sigma_{g_{2}})}+\cdots+\zeta_{r}^{\epsilon(\sigma_{g_{2}})},$
$=\chi((g_{1}^{-1}hg_{1})^{\epsilon(\sigma_{g_{2}})}),$
$=\chi(g_{1}^{-1}h^{\epsilon(\sigma_{g_{2}})}g_{1}),$
$=\chi(\underline{\gamma}^{-1}*h),$
as required. ∎
Consider the $L$-vector space $\operatorname{Fun}(H/\!/H,L)$ of functions from $H$ to $L$ which are conjugation invariant. This has a natural action of $G\times G$ where
$(g_{1},g_{2})*f\coloneqq\sigma_{g_{2}}(f(g_{1}^{-1}-g_{1})).$
This induces a natural action of $G/H\times G/H$ and hence of $\Gamma\times\Gamma$ on $\operatorname{Fun}(H/\!/H,L)$, the subspace of functions invariant by conjugation by $H$. Because $L$ splits $H$, this has a basis given by irreducible characters $\chi\in\operatorname{Irr}(H,L)$ [LOR, Cor. 3.21], and with respect to this basis $\operatorname{Fun}(H/\!/H,L)$ is nothing but the permutation representation $L[\operatorname{Irr}_{L}(H)]$ of $\Gamma\times\Gamma$. Similarly, the centre $Z(L[H])$ of $L[H]$ has a basis given by conjugacy class sums $c_{h}$, which induces a permutation action of $\Gamma\times\Gamma$ on $Z(L[H])$, which is explicitly given by
$\underline{\gamma}\cdot\left(\sum_{h}\lambda_{h}h\right)=\sum_{h}\lambda_{\underline{\gamma}^{-1}*h}h.$
Lemma 14.2.
The perfect pairing
$\operatorname{Fun}(H/\!/H,L)\times Z(L[H])\rightarrow L,\qquad\left(f,\sum_{h}\lambda_{h}h\right)=\sum_{h}\lambda_{h}f(h)$
is $\Gamma\times\Gamma$-equivariant. In particular, this induces a $\Gamma\times\Gamma$-equivariant isomorphism
$\operatorname{Fun}(H/\!/H,L)\xrightarrow{\sim}\operatorname{Hom}_{L}(Z(L[H]),L),\qquad f\mapsto\Phi_{f}.$
Proof.
Expressing any $f\in\operatorname{Fun}(H/\!/H,L)$ as a sum of $\chi\in\operatorname{Irr}_{L}(H)$, this follows from Lemma 14.1. ∎
For any group which acts on a finite set $X$, there is a equivariant perfect pairing
$L[X]\times L[X]\rightarrow L,\qquad\left(\sum_{x}\lambda_{x}x,\sum_{y}\mu_{y}y\right)=\sum_{x}\lambda_{x}\mu_{x},$
which induces a canonical isomorphism of representations
$L[X]\xrightarrow{\sim}\operatorname{Hom}_{L}(L[X],L),$
where $x\in X$ is sent to its indicator function. In particular:
Corollary 14.3.
The permutation representations of $\Gamma\times\Gamma$ associated to $\operatorname{Irr}_{L}(H)$ and $\operatorname{Cl}(H)$ are canonically isomorphic.
Remark 14.4.
Even though the permutation representations of the $\Gamma\times\Gamma$-sets $\operatorname{Irr}_{L}(H)$ and $\operatorname{Cl}(H)$ are the same, it is not true in general that $\operatorname{Irr}_{L}(H)$ and $\operatorname{Cl}(H)$ are isomorphic as $\Gamma\times\Gamma$-sets. For example, for the field extension ${\mathbb{Q}}(i)/{\mathbb{Q}}$, an example of a pair $(G,H)$ of the smallest order where this fails is $\text{SmallGroup}(32,9)$ with $H$ the unique subgroup of $G$ isomorphic to $\text{SmallGroup}(16,4)$. The $\Gamma\times\Gamma=K_{4}$-sets $\operatorname{Irr}_{L}(H)$ and $\operatorname{Cl}(H)$ are non-isomorphic as one can compute that $|\operatorname{Irr}_{L}(H)^{K_{4}}|=4$ whereas $|\operatorname{Cl}(H)^{K_{4}}|=6$. This condition is natural to check, it being true that for any degree two extension $L/K$, if $\operatorname{Irr}_{L}(H)$ and $\operatorname{Cl}(H)$ are non-isomorphic then one must necessarily have that $|\operatorname{Irr}_{L}(H)^{K_{4}}|\neq|\operatorname{Cl}(H)^{K_{4}}|$ by [SERRE, Ex. 13.5] as $K_{4}$ is the only non-cyclic subgroup of $K_{4}$.
However, when $L/K$ is $\mathbb{C}$$\mathbb{R}$}, the $\Gamma\times\Gamma$-sets $\operatorname{Irr}_{L}(H)$ and $\operatorname{Cl}(H)$ are isomorphic [RUMTAY3, Cor. 1.5].
Now we consider the diagonal action of $\Gamma$ on $\operatorname{Irr}_{L}(H)$ and $\operatorname{Cl}(H)$, which allows us to count the number of irreducible semilinear representations of $G$ from just the group theory of $G$. This generalises [RUMTAY, Thm. 5.6] for $\mathbb{C}$$\mathbb{R}$}, and [SERRE, §12.4, Cor. 2] in the split case (cf. Section 10).
Corollary 14.5.
The number of irreducible semilinear representations of $G$ over $L$ is the same as the number of $G$-orbits on $\operatorname{Cl}(G)$: $|\operatorname{Irr}_{L}^{\rtimes}(G)|=|\operatorname{Cl}(H)/\hskip 1.00006pt\Gamma|$.
Proof.
This follows from [SERRE, Ex. 13.5], Corollary 14.3, and the bijection between $\operatorname{Irr}_{L}(H)/\Gamma$ and $\operatorname{Irr}_{L}^{\rtimes}(G)$ of Theorem 8.20. ∎
15. Cohomological Interpretation
We can interpret the main results of Section 8 in terms of Galois cohomology.
From the matrix form of semilinear representations described in Section 2.1, for any $n\geq 1$ we obtain an identification between $\mathrm{H}^{1}(G,\operatorname{GL}_{n}(L))$ and the set of isomorphism classes of $n$-dimensional semilinear representations of $G$ over $L$. From the exact sequence
$1\rightarrow H\rightarrow G\rightarrow\Gamma\rightarrow 1$
we have the restriction-inflation exact sequence of pointed sets [SERGC, §5.8(a)]
$1\rightarrow\mathrm{H}^{1}(\Gamma,\operatorname{GL}_{n}(L)^{H})\rightarrow\mathrm{H}^{1}(G,\operatorname{GL}_{n}(L))\rightarrow\mathrm{H}^{1}(H,\operatorname{GL}_{n}(L))^{\Gamma},$
which has trivial first term
$\mathrm{H}^{1}(\Gamma,\operatorname{GL}_{n}(L)^{H})=\mathrm{H}^{1}(\Gamma,\operatorname{GL}_{n}(L))=1$
by Hilbert’s Theorem 90. Furthermore, $\mathrm{H}^{1}(H,\operatorname{GL}_{n}(L))$ is the set of isomorphism classes of $n$-dimensional linear representations of $H$ over $L$, and the action of $\Gamma$ on $H^{1}(H,\operatorname{GL}_{n}(L))$ corresponds to the action of $\Gamma$ on isomorphism classes of objects of $\operatorname{\mathbf{Rep}}_{L}(H)$ as described in Definition 8.16.
The fact that the image of $\mathrm{H}^{1}(G,\operatorname{GL}_{n}(L))$ is contained in $\mathrm{H}^{1}(H,\operatorname{GL}_{n}(L))^{\Gamma}$ amounts to the fact that any representation of $H$ obtained by restricting a semilinear representation will be fixed by the action of $\Gamma$. The fact that
$1\rightarrow\mathrm{H}^{1}(G,\operatorname{GL}_{n}(L))\rightarrow\mathrm{H}^{1}(H,\operatorname{GL}_{n}(L))^{\Gamma}$
is an exact sequence of pointed sets amounts to the fact that the only $n$-dimensional semilinear representation which restricts to the trivial $n$-dimensional linear representation of $H$ is the trivial semilinear representation. This can be considered as a weaker form of Corollary 8.6.
15.1. One-Dimensional Representations
We can say more when $n=1$. In this case $\operatorname{GL}_{1}(L)=L^{\times}$ is abelian, and the pointed sets considered above are abelian groups. Furthermore, from the Lyndon-Hochschild-Serre spectral sequence there is an exact sequence
$0\rightarrow\mathrm{H}^{1}(\Gamma,(L^{\times})^{H})\rightarrow\mathrm{H}^{1}(G,L^{\times})\rightarrow\mathrm{H}^{1}(H,L^{\times})^{\Gamma}\xrightarrow{{\mathfrak{T}}}\mathrm{H}^{2}(\Gamma,L^{\times})\rightarrow\mathrm{H}^{2}(G,L^{\times})$
of low degree terms which extends the inflation-restriction sequence above [NSW, Prop. 1.6.7, Thm. 2.4.1]. Again the first term
$\mathrm{H}^{1}(\Gamma,(L^{\times})^{H})=\mathrm{H}^{1}(\Gamma,L^{\times})=0$
is trivial by Hilbert’s Theorem 90, and so we have an exact sequence of abelian groups
$0\rightarrow\mathrm{H}^{1}(G,L^{\times})\rightarrow\mathrm{H}^{1}(H,L^{\times})^{\Gamma}\xrightarrow{{\mathfrak{T}}}\mathrm{H}^{2}(\Gamma,L^{\times}).$
The group $\mathrm{H}^{1}(G,L^{\times})$ classifies $1$-dimensional semilinear representations of $G$ over $L$, and, as $H$ acts trivially on $L^{\times}$, the group $\mathrm{H}^{1}(H,L^{\times})^{\Gamma}$ is canonically identified with the set $\operatorname{Hom}(H,L^{\times})^{\Gamma}$ of linear characters of $H$ over $L$ which are fixed by the action of $\Gamma$, where
$(\gamma*\chi)(h)\coloneqq\gamma(\chi(g(\gamma)^{-1}hg(\gamma)))$
for any $\gamma\in\Gamma$, $h\in H$ and $\chi\in\operatorname{Hom}(H,L^{\times})$, and $g(\gamma)\in G$ is any lift of $\gamma$. The injectivity of
$\mathrm{H}^{1}(G,L^{\times})\rightarrow\operatorname{Hom}(H,L^{\times})^{\Gamma}$
amounts to the statement that any linear character of $H$ over $L$ which is invariant under $\Gamma$ has at most one extension of a semilinear character of $G$ over $L$, which is Corollary 8.6 for $1$-dimensional representations.
15.2. The Transgression Map
More interesting is the transgression map, $\mathfrak{T}$}. By [NSW, Thm. 2.4.3] this agrees with the map constructed in [NSW, Prop. 1.6.6], which, following this construction in our situation, is explicitly defined as follows.
Let $(g_{\gamma})_{\gamma\in\Gamma}$ be a set of lifts in $G$ of $\gamma\in\Gamma$, or in other words a set of coset representatives of $H$ in $G$, which we choose with $g_{1}=1\in G$. For any $\gamma_{1},\gamma_{2}\in\Gamma$, we have that
$g_{\gamma_{1}}g_{\gamma_{2}}=g_{\gamma_{1}\gamma_{2}}h_{\gamma_{1},\gamma_{2}},$
for a unique $h_{\gamma_{1},\gamma_{2}}\in H$. Then the transgression map $\mathfrak{T}$} is defined by
${\mathfrak{T}}\colon\operatorname{Hom}(H,L^{\times})^{\Gamma}\rightarrow\mathrm{H}^{2}(\Gamma,L^{\times}),\qquad{\mathfrak{T}}(\chi)\coloneqq[f_{\chi}],$
where $f_{\chi}$ is the $2$-cocycle
$f_{\chi}\colon\Gamma\times\Gamma\rightarrow L^{\times},\qquad f_{\chi}(\gamma_{1},\gamma_{2})=(\gamma_{1}\gamma_{2})(\chi(h_{\gamma_{1},\gamma_{2}}))^{-1}.$
There is also another natural map from $\operatorname{Hom}(H,L^{\times})^{\Gamma}$ to $\mathrm{H}^{2}(\Gamma,L^{\times})$, which arises from Theorem B. Writing $\operatorname{Br}(L/K)$ for the subgroup of classes of $\operatorname{Br}(K)$ which are split over the Galois extension $L$, there is an explicit group isomorphism
$\Psi\colon\operatorname{Br}(L/K)\xrightarrow{\sim}\mathrm{H}^{2}(\Gamma,L^{\times})$
as described in [milneCFT, §IV.3]. We therefore describe our map as a map
$\Phi\colon\operatorname{Hom}(H,L^{\times})^{\Gamma}\rightarrow\operatorname{Br}(L/K),$
which we define as follows. For any $\chi\in\operatorname{Hom}(H,L^{\times})^{\Gamma}$, let $L_{\chi}\in\operatorname{Irr}_{L}(H)$ denote the associated $1$-dimensional representation of $H$, and let $V_{\chi}\in\operatorname{Irr}_{L}^{\rtimes}(G)$ denote the corresponding semilinear representation of $G$ over $L$ (cf. Theorem B). Because $L_{\chi}$ is $1$-dimensional, we have that
$\operatorname{End}_{L[H]}(L_{\chi})=L,$
and therefore the division algebra
$D_{\chi}\coloneqq\operatorname{End}_{L\rtimes G}(V_{\chi})$
has
$Z(D_{\chi})\cong L^{\Gamma}=K$
by Corollary 8.27, using the fact that $\Gamma$ fixes $\chi$. In particular, $D_{\chi}$ is a central division algebra over $K$, and we set
$\Phi(\chi)\coloneqq[D_{\chi}]\in\operatorname{Br}(K).$
From Corollary 8.7 restriction defines an isomorphism
$L\otimes_{K}D_{\chi}\xrightarrow{\sim}\operatorname{End}_{L[H]}(V|_{H}),$
and by Theorem B we have that
$V_{\chi}|_{H}\cong L_{\chi}^{m(V_{\chi})},$
so
$\operatorname{End}_{L[H]}(V|_{H})\cong\operatorname{End}_{L[H]}(L_{\chi}^{m(V_{\chi})})\cong M_{m(V_{\chi})}(L).$
This shows that $\Phi(\chi)\in\operatorname{Br}(L/K)$, and therefore that $\Phi$ is well-defined.
The following shows that the two maps $\mathfrak{T}$} and $\Phi$ agree.
Proposition 15.1.
The transgression homomorphism factors as the composition ${\mathfrak{T}}=\Psi\circ\Phi$.
Proof.
Let $\chi\in\operatorname{Hom}(H,L^{\times})^{\Gamma}$. To see that ${\mathfrak{T}}(\chi)=(\Psi\circ\Phi)(\chi)$, it is equivalent to show that
${\mathfrak{T}}(\chi)^{-1}=\Psi(\Phi(\chi))^{-1}.$
The map $\Psi$ is a homomorphism, so $\Psi(\Phi(\chi))^{-1}=\Psi(\Phi(\chi)^{-1})$, and $\Phi(\chi)=[D_{\chi}]\in\operatorname{Br}(L/K)$, so $\Phi(\chi)^{-1}=[D_{\chi}^{\text{op}}]$. We therefore want to show that
$\Psi([D_{\chi}^{\text{op}}])={\mathfrak{T}}(\chi)^{-1}=[f_{\chi}]^{-1}\in\mathrm{H}^{2}(\Gamma,L^{\times}).$
To see this, let us recall how to compute $\Psi([D_{\chi}^{\text{op}}])$, following [milneCFT, Thm. IV.3.11]. First, one takes a central simple $K$-algebra AA which has dimension $|\Gamma|^{2}$ over $K$, and $D_{\chi}^{\text{op}}]=[A]$. This can be found, following [milneCFT, Cor. IV.3.6], by using the isomorphism
$L\otimes_{K}D_{\chi}\xrightarrow{\sim}\operatorname{End}_{L}(V_{\chi}),$
and setting $A\coloneqq C_{\operatorname{End}_{K}(V_{\chi})}(D_{\chi})$ to be the centraliser of $D_{\chi}$ in $\operatorname{End}_{K}(V_{\chi})$. Then, as described in [milneCFT, Thm. IV.3.11], one takes elements $(e_{\gamma})_{\gamma\in\Gamma}$ in $A^{\times}$ which satisfy
(8) $e_{\gamma}\cdot a=\gamma(a)\cdot e_{\gamma}$
for any $a\in L$, and sets $\Psi([D_{\chi}^{\text{op}}])\coloneqq\Psi([A])\coloneqq[f]$, where
$f\colon\Gamma\times\Gamma\rightarrow L^{\times},\qquad f(\gamma_{1},\gamma_{2})\coloneqq e_{\gamma_{1}}\cdot e_{\gamma_{2}}\cdot e_{\gamma_{1}\gamma_{2}}^{-1}\in L^{\times}.$
For our $A=C_{\operatorname{End}_{K}(V_{\chi})}(D_{\chi})$ we can find these elements explicitly. Writing $\rho\colon G\rightarrow\operatorname{End}_{K}(V_{\chi})$ for the action map of the semilinear $G$-representation $V_{\chi}$, we may set
$e_{\gamma}\coloneqq\rho(g_{\gamma}),$
where the $(g_{\gamma})_{\gamma\in\Gamma}$ are the lifts of elements of $\Gamma$ we used to define $f_{\chi}$ and ${\mathfrak{T}}(\chi)$ above. These commute with all elements of $D_{\chi}$ directly from the definition of $D_{\chi}$, and so lie in $A$. They also satisfy the equation (8). We may therefore compute that
$f(\gamma_{1},\gamma_{2})$ $=\rho(g_{\gamma_{1}})\cdot\rho(g_{\gamma_{2}})\cdot\rho(g_{\gamma_{1}\gamma_{2}})^{-1},$
$=\rho(g_{\gamma_{1}}g_{\gamma_{2}})\cdot\rho(g_{\gamma_{1}\gamma_{2}})^{-1},$
$=\rho(g_{\gamma_{1}\gamma_{2}}h_{\gamma_{1},\gamma_{2}})\cdot\rho(g_{\gamma_{1}\gamma_{2}})^{-1},$
$=\rho(g_{\gamma_{1}\gamma_{2}})\cdot\chi(h_{\gamma_{1},\gamma_{2}})\cdot\rho(g_{\gamma_{1}\gamma_{2}})^{-1},$
$=(\gamma_{1}\gamma_{2})(\chi(h_{\gamma_{1},\gamma_{2}}))\cdot\rho(g_{\gamma_{1}\gamma_{2}})\cdot\rho(g_{\gamma_{1}\gamma_{2}})^{-1},$
$=(\gamma_{1}\gamma_{2})(\chi(h_{\gamma_{1},\gamma_{2}})).$
In particular, $f(\gamma_{1},\gamma_{2})=f_{\chi}(\gamma_{1},\gamma_{2})^{-1}$, and so $\Psi([D_{\chi}^{\text{op}}])={\mathfrak{T}}(\chi)^{-1}=[f_{\chi}]^{-1}$ as required. ∎
We get the following immediate consequence, which is not at all obvious from the definition of $\Phi$.
Corollary 15.2.
$\Phi\colon\operatorname{Hom}(H,L^{\times})^{\Gamma}\rightarrow\operatorname{Br}(L/K)$ is a group homomorphism.
Remark 15.3.
We note that when $H\hookrightarrow G$ is split, the map $\Phi$ is zero, as explained in the comment following the proof of [NSW, Thm. 2.4.4]. One can also see this more directly as a consequence of Proposition 10.6.
Corollary 15.2 has consequences for the numbers $m_{K}^{L}(\chi)$, for $\chi\in\operatorname{Hom}(H,L^{\times})^{\Gamma}$.
Corollary 15.4.
For $\chi\in\operatorname{Hom}(H,L^{\times})^{\Gamma}$, the prime factors of $m_{K}^{L}(\chi)$ are a subset of the prime factors of $\operatorname{Ord}(\chi)$. In particular, if $\operatorname{Ord}(\chi)$ is coprime to $[L:K]$, then $m_{K}^{L}(\chi)=1$.
Proof.
Because $\Phi$ is a group homomorphism, $\operatorname{Ord}(\Phi(\chi))\mid\operatorname{Ord}(\chi)$. By [GS, Prop. 4.5.13], we also have that $\operatorname{Ord}(\Phi(\chi))\mid\operatorname{Deg}(\Phi(\chi))$ and $\operatorname{Ord}(\Phi(\chi))$ and $\operatorname{Deg}(\Phi(\chi))$ share the same prime factors. The result follows as $\operatorname{Deg}(\Phi(\chi))=m_{K}^{L}(\chi)$ by Proposition 8.29 and $m_{K}^{L}(\chi)\mid[L:K]$ by Proposition 9.3. ∎
Remark 15.5.
Whenever $K$ is such that the period coincides with the index of elements of $\operatorname{Br}(K)$, one further has that $m_{K}^{L}(\chi)\mid\operatorname{Ord}(\chi)$. For example, this is true for local fields [milneCFT, Rem. IV.4.4(b)] and global fields [milneCFT, Thm. VIII.2.6].
16. Realisation of Division Algebras
The Schur subgroup of a field $K$ is the subgroup $S(K)$ of $\operatorname{Br}(K)$ consisting of those classes $[A]$, where $A$ is a central simple algebra over $k$ for which there exists a surjection $K[H]\twoheadrightarrow A$ for some finite group $H$. In general $S(K)\neq\operatorname{Br}(K)$, and is typically much smaller, being explicitly described when $K$ has characteristic 0 as the subgroup $C(K)$ of $\operatorname{Br}(K)$ consisting of classes represented by cyclotomic algebras [YAM].
In this section we show that the semilinear representations of the type we have considered in this paper are sufficiently more general than ordinary group algebras to realise all elements of $\operatorname{Br}(K)$. We also direct the reader towards [MEIR], where similar ideas are used to show that any central simple algebra over $K$ is Brauer equivalent to a quotient of a finite-dimensional Hopf algebra over $k$.
Theorem 16.1.
Suppose that $K$ be a field, and $A$ is a central simple algebra over $K$. Then there is a finite Galois extension $L/K$, a group $G$ and a surjective group homomorphism $G\rightarrow\operatorname{Gal}(L/K)$ such that $A$ is Brauer equivalent to a quotient of $L\rtimes G$. If the order of $[A]$ in $\operatorname{Br}(K)$ is coprime to $\operatorname{char}(K)$, then $G$ can be taken to be finite.
Proof.
Let $L$ be a finite Galois extension that splits $A$, and set $\Gamma\coloneqq\operatorname{Gal}(L/K)$. Then from the isomorphism, $\mathrm{H}^{2}(\operatorname{Gal}(L/K),L^{\times})\cong\operatorname{Br}(L/K)$ [milneCFT, Thm. IV.3.14], $A]=[A(L/K,\alpha)]$ in $\operatorname{Br}(K)$ for some normalised cocycle $\alpha\colon\Gamma\times\Gamma\rightarrow L^{\times}$, where $A(L/K,\alpha)$ is the central simple $K$-algebra
$A(L/K,\alpha)=\bigoplus_{\gamma\in\Gamma}L\cdot u_{\gamma}$
for formal symbols $(u_{\gamma})_{\gamma\in\Gamma}$, with multiplication
$\lambda u_{\gamma}\cdot\mu u_{\delta}=\lambda\gamma(\mu)\alpha(\gamma,\delta)u_{\gamma\delta}.$
We take
$G\coloneqq\{L^{\times}\cdot u_{\gamma}\mid\gamma\in\Gamma\}\subset A(L/K,\alpha)^{\times},$
for which there is a natural surjective group homomorphism
$G\rightarrow\Gamma,\qquad\lambda\cdot u_{\gamma}\mapsto\gamma.$
Then we can define a $K$-algebra homomorphism
$\phi\colon L\rtimes G\rightarrow A(L/K,\alpha),\qquad\lambda\cdot(\mu\cdot u_{\gamma})\mapsto\lambda\mu\cdot u_{\gamma}$
which is clearly surjective, and multiplicative from the definition of the action of $G$ on $L$ and the fact that $\alpha$ is normalised. This gives the first statement.
For the second statement, write $G_{K}\coloneqq\operatorname{Gal}(K^{\text{sep}}/K)$. If $[A]$ has order $m$ in $\operatorname{Br}(K)$, and $m$ is coprime to the characteristic of $K$, then the natural inclusion $\mathrm{H}^{2}(G_{K},\mu_{m}(K^{\text{sep}}))\hookrightarrow\mathrm{H}^{2}(G_{K},(K^{\text{sep}})^{\times})$ has image $\mathrm{H}^{2}(G_{K},(K^{\text{sep}})^{\times})[m]$ [GS, Cor. 4.4.9]. In particular, under the isomorphism $\operatorname{Br}(K)[m]\cong\mathrm{H}^{2}(G_{K},(K^{\text{sep}})^{\times})[m]$ [milneCFT, Cor. IV.3.16], $A]=[A(L/K,\alpha)]$ for some finite Galois extension $L/K$ and normalised cocycle $\alpha\colon\Gamma\times\Gamma\rightarrow\mu_{m}(L)$, where again $\Gamma\coloneqq\operatorname{Gal}(L/K)$. Then we may proceed as above, taking $G$ to be the, now finite, group
$G\coloneqq\{\mu_{m}(L)\cdot u_{\gamma}\mid\gamma\in\Gamma\}\subset A(L/K,\alpha)^{\times},$
which is a subgroup because $\alpha$ is valued in $\mu_{m}(L)$. ∎
This allows us to show that all central division algebras over $K$ arise as endomorphism rings of semilinear representations, and that all possible values for the semilinear Schur indices $m(V)$ occur.
Corollary 16.2.
Suppose that $K$ be a field, and $D$ is a central division algebra over $K$. Then there is a finite Galois extension $L/K$, a group $G$, a surjective group homomorphism $G\rightarrow\operatorname{Gal}(L/K)$ and $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$ such that $D\cong\operatorname{End}_{L\rtimes G}(V)$ and $m(V)=\operatorname{Deg}(D)$. If the order of $[D]$ in $\operatorname{Br}(K)$ is coprime to $\operatorname{char}(K)$, then $G$ can be taken to be finite.
Proof.
Taking $A\coloneqq D^{\text{op}}$ in Theorem 16.1, we have a surjection $L\rtimes G\twoheadrightarrow M_{n}(D^{\text{op}})$ for some $n\geq 1$, and we can take $V\coloneqq(D^{\text{op}})^{n}$, which is irreducible as a $M_{n}(D^{\text{op}})$-module with endomorphism ring $D$. To see that $m(V)=\operatorname{Deg}(D)$, note that $V|_{H}$ is semisimple by Corollary 8.14, and that
$L\otimes_{K}D=L\otimes_{K}\operatorname{End}_{L\rtimes G}(V)\xrightarrow{\sim}\operatorname{End}_{L[H]}(V|_{H})$
by Corollary 8.7. From how $L$ is defined in the proof of Theorem 16.1, $D^{\text{op}}$ is split over $L$, and therefore the same is true of $D$. In particular, $\operatorname{End}_{L[H]}(V|_{H})$ is a matrix ring over $L$, and so the semisimple $V|_{H}$ has only one isotypic factor: $V|_{H}=W^{m(V)}$ for some $W\in\operatorname{Irr}_{L[H]}(H)$ with $\operatorname{End}_{L[H]}(W)=L$. Then the fact that $m(V)=\operatorname{Deg}(D)$ follows by comparing dimensions in the above isomorphism. ∎
Remark 16.3.
We point out that in Theorem 16.1 (and therefore in Corollary 16.2), the field $L$ can be taken to be any field that splits $A$. However, when one takes $G$ to be finite, one loses control the field $L$, and the field $L$ constructed by the proof in the case that $G$ is finite might have degree greater than that of a splitting field for $A$.
For example, the group $\operatorname{Br}({\mathbb{Q}})[2]$ parametrises quaternion algebras over $\mathbb{Q}$}. For any such quaternion algebra $D$, the proof realises $D$ as the endomorphism ring of some irreducible $V\in\operatorname{Irr}_{L}^{\rtimes}(G)$, for some Galois extension $L/{\mathbb{Q}}$ and takes $H\coloneqq\mu_{2}(L)\cong C_{2}$.
On the other hand, we can compute the quaternion algebras $D$ which arise from semilinear representations of degree two extensions $L/{\mathbb{Q}}$ and group extensions $H\leq G$ of index two with $H\cong C_{2}$.
There are two options for $H\leq G$: $C_{2}\leq C_{2}\times C_{2}$ or $C_{2}\leq C_{4}$. The first has $m_{{\mathbb{Q}}}^{L}(W)=1$ for all $W\in\operatorname{Irr}_{L}(C_{2})$ and degree two $L/{\mathbb{Q}}$, and so does not realise any quaternion algebras over $\mathbb{Q}$}. For the second, $L={\mathbb{Q}}(\sqrt{d})$ for some $d\in{\mathbb{Q}}^{\times}\setminus{\mathbb{Q}}^{\times 2}$, and from Example 11.6 $D\cong(-1,d)_{{\mathbb{Q}}}$ for some $a\in{\mathbb{Q}}^{\times}$. However such $D$ are exactly those quaternion algebras which define elements of the proper subgroup $\operatorname{Br}({\mathbb{Q}}(i)/{\mathbb{Q}})[2]$ of $\operatorname{Br}({\mathbb{Q}})[2]$.
Question 16.4.
The above remark shows that not all elements of $\operatorname{Br}({\mathbb{Q}})[2]$ are realised when $L/{\mathbb{Q}}$ has degree two and $H\cong C_{2}$, but all elements are realised when $L$ is allowed to be any extension of $\mathbb{Q}$} (actually, computing explicitly with cocycles, one can always take $L$ to be at most degree $4$ over $\mathbb{Q}$}).
It is natural to ask if the same is true when instead the degree of $L$ is fixed and $H$ is allowed to grow: are all elements of $\operatorname{Br}({\mathbb{Q}})[2]$ realised for degree two extensions $L/{\mathbb{Q}}$, for a degree two extension $H\leq G$ of finite groups of arbitrary order?
17. Extension to Infinite Galois Extensions
We now give a brief indication of how the results of this paper can be extended to understand semilinear representations of infinite Galois extensions.
Let $L/K$ be a potentially infinite Galois extension, and suppose that $\sigma\colon G\twoheadrightarrow\operatorname{Gal}(L/K)$ is a continuous surjection of topological groups which is open.
Example 17.1.
As an example to illustrate that such situations arise naturally, let $F$ be a finite extension of ${\mathbb{Q}}_{p}$. From local class field theory that there is an isomorphism of topological groups
$\mathcal{O}_{F}^{\times}\xrightarrow{\sim}\operatorname{Gal}(F^{\text{ab}}/F^{\text{nr}})$
where $F^{\text{nr}}$ is the maximal unramified extension of $F$, and $F^{\text{ab}}$ is the maximal abelian extension of $F$.
The determinant
$\det\colon\operatorname{GL}_{n}(\mathcal{O}_{F})\rightarrow\mathcal{O}_{F}^{\times}\xrightarrow{\sim}\operatorname{Gal}(F^{\text{ab}}/F^{\text{nr}})$
satisfies the above assumptions, as does the reduced norm
$\text{Nrd}\colon\mathcal{O}_{D}^{\times}\rightarrow\mathcal{O}_{F}^{\times}\xrightarrow{\sim}\operatorname{Gal}(F^{\text{ab}}/F^{\text{nr}})$
for any central division algebra $D$ over $F$.
Definition 17.2.
We denote by
$\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$
the full subcategory of $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$ of objects $V$ for which
$V=\bigcup_{H\leq_{o}G}V^{H},$
as $H$ ranges over the open subgroups of $G$.
Example 17.3.
When $L/K$ is finite and $G$ is any group with a surjective group homomorphism $G\twoheadrightarrow\operatorname{Gal}(L/K)$, one can take $G$ to have the discrete topology. This satisfies the above assumptions, and $\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$ is nothing but $\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$.
We can characterise this in several equivalent ways.
Lemma 17.4.
Suppose that $V\in\operatorname{\mathbf{Rep}}_{L}^{\rtimes}(G)$. Then the following are equivalent:
  1. (1)
    $V\in\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$,
  2. (2)
    Every $v\in V$ is fixed by some open subgroup of $G$,
  3. (3)
    $\operatorname{Stab}_{G}(v)$ is open in $G$,
  4. (4)
    There is an open subgroup $H$ of $G$ such that $L\cdot V^{H}=V$.
When $G$ has a basis of open normal subgroups, these are equivalent to:
  1. (5)
    There is an open normal subgroup $N$ of $G$ such that $L\cdot V^{N}=V$.
For example, the assumption that $G$ has a basis of open normal subgroups is satisfied in the examples of Example 17.1 above.
Proof.
The equivalence of the first three conditions is straightforward. To see that (4) implies (2), suppose that $L\cdot V^{H}=V$ for some open subgroup $H$ of $G$. Then any $v\in V$ is of the form
$\sum_{i}\lambda_{i}v_{i}\in L\cdot V^{H}.$
The $\lambda_{i}$ lie in some finite Galois extension $F$ of $K$, and so are fixed by some open subgroup $\operatorname{Gal}(L/F)$. In particular, $v$ is fixed by the open subgroup $H\cap\sigma^{-1}(\operatorname{Gal}(L/F))$. To see that (2) implies (4), let $e_{1},...,e_{n}$ be a basis of $V$ over $L$. Taking $H$ to be the intersection of the open subgroups $\operatorname{Stab}_{G}(e_{i})$, we have that $L\cdot V^{H}=V$. The equivalence of point (5) with (4) is immediate. ∎
The link between the categories we have considered and $\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$ is the following. If $N$ is an open normal subgroup of $G$, then $L^{N}/K$ is a finite Galois extension of $K$, because $\sigma(N)$ is open and normal in $\operatorname{Gal}(L/K)$. In particular, there is a surjective map
$G/N\twoheadrightarrow\operatorname{Gal}(L^{N}/K),$
and we may consider the category $\operatorname{\mathbf{Rep}}_{L^{N}}^{\rtimes}(G/N)$, which satisfies the assumptions of Section 3, and therefore is described by the results of this paper. The following is a restatement of [TAY4, Cor. 3.9].
Proposition 17.5.
Suppose that $N$ is an open normal subgroup of $G$. Then the functor
$L\otimes_{L^{N}}-\colon\operatorname{\mathbf{Rep}}_{L^{N}}^{\rtimes}(G/N)\rightarrow\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$
is exact, monoidal, fully faithful, and the essential image is closed under sub-quotients. The essential image is explicitly described as those $V\in\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$ for which $L\cdot V^{N}=V$, and on this full subcategory
$(-)^{N}\colon\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)\rightarrow\operatorname{\mathbf{Rep}}_{L^{N}}^{\rtimes}(G/N)$
defines a quasi-inverse to $L\otimes_{L^{N}}$-.
In particular, the indecomposable (resp. irreducible) objects $V$ of $\operatorname{\mathbf{Rep}}_{L,\operatorname{sm}}^{\rtimes}(G)$ which satisfy $L\cdot V^{N}=V$ are canonically identified with $\operatorname{Ind}_{L^{N}}^{\rtimes}(G/N)$ (resp. $\operatorname{Irr}_{L^{N}}^{\rtimes}(G/N)$), which is described by Theorem B.
Example 17.6.
Continuing Example 17.1 above, then taking $N\coloneqq 1+\pi^{r}M_{n}(\mathcal{O}_{F})$ for $r\geq 1$ and a uniformiser $\pi$ of $\mathcal{O}_{F}$,
$G/N=\operatorname{GL}_{n}(\mathcal{O}_{F}/1+\pi^{r}\mathcal{O}_{F}),\ \text{ and }\ L^{N}=F^{\text{nr}}_{r},$
where $F^{\text{nr}}_{r}$ denotes the $r$th Lubin-Tate extension of $F^{\text{nr}}$ from explicit local class field theory.
For example, when $F={\mathbb{Q}}_{p}$ and $r=1$, the surjection $G/N\rightarrow\operatorname{Gal}(L^{N}/K)$ becomes
$\det\colon\operatorname{GL}_{n}({\mathbb{F}}_{p})\twoheadrightarrow{\mathbb{F}}_{p}^{\times}\xrightarrow{\sim}\operatorname{Gal}({\mathbb{Q}}_{p}^{\text{nr}}(\zeta_{p})/{\mathbb{Q}}_{p}^{\text{nr}}).$